### v1

# Time Series Foundation Models: Benchmarking Challenges and Requirements

Abstract
Time Series Foundation Models (TSFMs) represent a new paradigm for time series forecasting, offering zero-shot forecasting capabilities without the need for domain-specific pre-training or fine-tuning. However, as with Large Language Models (LLMs), evaluating TSFMs is tricky, as with ever more extensive training sets, it becomes more and more challenging to ensure the integrity of benchmarking data. Our investigation of existing TSFM evaluation highlights multiple challenges, ranging from the representativeness of the benchmark datasets, over the lack of spatiotemporal evaluation, to risks of information leakage due to overlapping and obscure datasets, and the memorization of global patterns caused by external shocks like economic crises or pandemics. Our findings reveal widespread confusion regarding data partitions, risking inflated performance estimates and incorrect transfer of global knowledge to local time series. We argue for the development of robust evaluation methodologies to prevent pitfalls already observed in LLM and classical time series benchmarking, and call upon the research community to design new, principled approaches, such as evaluations on truly out-of-sample future data, to safeguard the integrity of TSFM assessment.
1 Introduction
Time Series Foundation Models (TSFMs) are an emerging class of time series forecasting models inspired by the architecture and training procedures of foundation models in natural language processing, especially Large Language Models (LLMs). Prominent examples of TSFMs include Chronos (Ansari et al., 2024), TimesFM (Das et al., 2024), Moirai/Moirai-MoE (Woo et al., 2024; Liu et al., 2024b), MOMENT (Goswami et al., 2024) and Time-MoE (Shi et al., 2024). In contrast to traditional time series forecasting models, which need to be trained from scratch on the time series that should be forecasted, TSFMs use transfer learning. That is, they are pre-trained on massive amounts of generic and/or domain-specific time series and are, thus, able to produce zero-shot forecasts.
However, pre-training TSFMs on massive amounts of time series data, typically scraped from publicly accessible online repositories, introduces challenges similar to those observed in LLM evaluation. The practice of training LLMs on much of the Internet has led to an LLM “evaluation crisis" (Liao & Xiao, 2023), describing a situation in which most benchmark data has already been exposed to an LLM during pre-training. Combining such test set contamination (Mirzadeh et al., 2024; Ravaut et al., 2024; Li et al., 2024a) with memorization effects (Chang et al., 2024) can result in overly-optimistic estimates of a model’s predictive performance and complicate the fair comparison of models through common benchmarks. Furthermore, as with LLMs (Chang et al., 2024), cross-validation in TSFMs is omitted due to its exhaustive resource requirements with large foundation models (Das et al., 2024; Gruver et al., 2024), and the diversity and generalizability of benchmarking datasets can be questioned (Aksu et al., 2024).
As the awareness about evaluation challenges in TSFMs is growing, researchers have begun to propose mitigation strategies, such as held-out-driven benchmarks (Aksu et al., 2024) and clean train-test splits within benchmarks (Qiu et al., 2024) that include more diverse and generalizing test sets Aksu et al. (2024). Especially when comparing TSFMs against each other, current approaches face limitations in defining an information-leakage-free test set for all investigated models (Aksu et al., 2024). This challenge is even more pronounced for larger models developed based on scaling laws, as their increased capacity may amplify the memorization effect, leading to a greater impact of test set contamination (see Section 2).
In this article, we discuss the evaluation challenges arising from current practices of training and evaluating TSFMs and propose requirements that a TSFM benchmark should fulfill. Empirically, our position is supported by an analysis of the lineage of the training and test sets of 15 prominent TSFMs and a further investigation of existing TSFM benchmarks [^1] TSFM benchmarks that existed in the beginning of 2025.. In addition, we provide a theoretical derivation of the phenomenon of global pattern memorization, which is specific to TSFMs and, to the best of our knowledge, has not been formally introduced before.
First, while we did not identify cases of information leakage within any of the underlying studies, our analysis highlights the increasing confusion with regards to a clear separation between training and test sets, one of the fundamental principles of time series forecasting and machine learning, and outlines possible adverse effects. Concretely, we found that one model’s training set is another model’s test set. This confusion, which is further exacerbated through the practice of remixing and renaming of datasets, stems from an ever-growing size of datasets, combined with non-transparent documentation about which exact data points have been used for training and testing, and complicates the fair benchmarking between TSFMs. We will show that this is not just a danger, but is actually already happening, using two examples where information leakage occurs in benchmarks 3.1.
Second, we highlight a more subtle instance of information leakage, particularly in time series forecasting, namely the memorization of global patterns. Global external shocks like the COVID-19 pandemic or global economic crises can profoundly impact numerous aspects of our economies and societies and, hence, are reflected in many diverse time series. A TSFM trained and evaluated on time series that are disjunct but stem from the same period and are both impacted by such an external shock can still profit from memorizing these patterns.
In addition to the issue of information leakage, we point out that current TSFM evaluation methods are not fully consistent with previous approaches in time series forecasting. In Section 3.3 and 3.4 we show that TSFMs typically do not implement best practices such as time series cross-validation or choosing representative time series for a domain.
This article argues that in the age of TSFMs a new benchmarking methodology is required to prevent a repetition of the LLM evaluation crisis and calls for new proposals that fulfill the requirements developed in this paper.
2 Foundation Models in Time Series
A foundation model is “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks" (Bommasani et al., 2021). The term is primarily associated with Transformer-based LLMs, especially decoder-only models like the ChatGPT or Llama families of models (OpenAI, 2023; Grattafiori et al., 2024). The success of foundation models can be attributed to scaling laws (Kaplan et al., 2020), that is, increasing training data, model parameters, and computational resources significantly improves the model’s performance.
LLMs have demonstrated remarkable zero-shot learning capabilities across diverse benchmarks (Liang et al., 2022). However, concerns have been raised that these models might just memorize publicly available questions and answers of well-known benchmarks (Chang et al., 2024). Memorization (Carlini et al., 2022) is closely related to test set contamination, describing the situation where exact or highly similar samples from a test set are leaked to the LLM’s pre-training data (Li et al., 2024a). This contamination is a side-effect of the vast size and incomprehensible nature of today’s pre-training datasets, such as Common Crawl [^2] https://commoncrawl.org/ or The Pile [^3] https://pile.eleuther.ai/ (Li et al., 2024a; Singh et al., 2024). Despite efforts to improve dataset curation in recent LLMs (Grattafiori et al., 2024), evidence shows that test set contamination remains prevalent in various forms inflating generalization performance (Mirzadeh et al., 2024; Ravaut et al., 2024; Li et al., 2024a).
Like generative LLMs, time series forecasting models are trained to autoregressively predict the next value(s) given a sequential series of past data points. Hence, it is not surprising, that researchers started to adapt decoder-only Transformers for time series forecasting. Some of these models encoded numerical time series values as natural language tokens (Gruver et al., 2024; Xue & Salim, 2024), others have developed novel tokenization techniques for numerical data, such as TimeGPT-1 (Garza et al., 2024) and Chronos (Ansari et al., 2024). Like LLMs, decoder-only Transformers for time series forecasting can be pre-trained on generic time series and adapted or fine-tuned to domain-specific time series optionally. At this, they seem to leverage the same scaling laws as LLMs (Edwards et al., 2024), leading to Time Series Foundation Models able to zero or few-shot forecast time series of different frequencies and domains adapting to diverse lengths of input context and forecasting horizons (Das et al., 2024; Liang et al., 2024).
Traditional statistical or machine-learning-based time series forecasting models have to be trained on the time series that should be forecasted. These local models exhibit only limited generalization capabilities, that is, they can only predict values that are independent and identically distributed (i.i.d.) to their training data. Combined with the common practice to use time-based train/test splits when evaluating time series models (Hyndman & Athanasopoulos, 2021), this property of local models protects them from information leakage.
Table 1: Summary of models, citations, and release years.
Model Name Author Year
Time-MoE Shi et al. (2024) 2025
Chronos Ansari et al. (2024) 2024
Lag-LLama Rasul et al. (2024) 2024
Moirai Woo et al. (2024) 2024
Moirai-MoE Liu et al. (2024b) 2024
MOMENT Goswami et al. (2024) 2024
Timer Liu et al. (2024c) 2024
TimesFM Das et al. (2024) 2024
Time-LLM Jin et al. (2024) 2024
TTM Ekambaram et al. (2024) 2024
UniTime Liu et al. (2024a) 2024
VisionTS Chen et al. (2024) 2024
ForecastPFN Dooley et al. (2023) 2023
GPT4TS Zhou et al. (2023) 2023
LLMTime Gruver et al. (2024) 2023
TimeGPT-1 Garza et al. (2024) 2023
TSFMs, in contrast, are often evaluated on zero-shot or few-shot forecasting tasks using held-out datasets, similar to LLM evaluation (Ansari et al., 2024; Liu et al., 2024b; Woo et al., 2024). Notably, the terms “zero-shot” and analogously “few-shot” need to be distinguished between LLMs and TSFMs: while LLMs adapt to different tasks without any example in zero-shot learning (Brown et al., 2020), TSFMs usually forecast zero-shot by adapting to different input lengths, frequencies, horizons and domains (Ansari et al., 2024; Das et al., 2024; Woo et al., 2024).
The specialized practices of pre-training and zero-shot forecasting in TSFMs introduce novel challenges while simultaneously amplifying existing difficulties in the evaluation of time series models, as discussed in the following sections.
3 Challenges of Evaluating Time Series Foundation Models
3.1 Challenge: Test Set Contamination Through Multi-Purpose Use of Datasets
In order to empirically study whether test set contamination is a problem in existing TSFM evaluations, we tried to track the lineage of the various datasets used for training and evaluating the TSFM listed in Table 1. Specifically, we distinguish between three types of data usages: pre-training datasets, train/test datasets, and zero-shot datasets.
Pre-training datasets are utilized exclusively for foundational training to teach typical time series patterns (e.g., trends, seasonality) to a model. Here, the entire time series is usually used for training, and no evaluation is conducted on these datasets. Train/test datasets, also called in-domain or in-distribution datasets, fit a model to time series from a specific domain (e.g., finance, energy, transportation). Following the traditional time series forecasting methodology, these datasets are typically split at a particular time point into training and test sets (Hyndman & Athanasopoulos, 2021; Kapoor & Narayanan, 2023). The training data is included in the model’s training corpus during pre-training or subsequent task-specific adaptation/fine-tuning. The test data is used to evaluate the model’s performance within the given domain. When benchmarking multiple models in a specific domain, all models should use the same datasets with the same train-test splits in order to ensure a fair comparison. Zero-shot datasets (held-out datasets) are entirely excluded from the pre-training and task-specific fine-tuning process and used for inference only. Unlike LLM zero-shot datasets with diverse tasks, TSFM datasets often share frequencies and domains with pre-training data, primarily evaluating frequency and domain generalization. However, in zero-shot evaluation, the TSFMs make forecasts, typically in a rolling manner, taking a limited context of historical values as inputs and producing output for future time steps (Aksu et al., 2024). Ideally, all TSFMs would be benchmarked on the same zero-shot datasets.
Figure 1: Lineage of dataset collections (left) used for training and evaluating recent Time Series Foundation Models (right). Typically, a collection contains multiple different datasets. Lines indicate cases where at least one dataset of a collection was used for pre-training, train/test, or zero-shot evaluation.
Our lineage analysis, visualized in Figure 1, reveals that each TSFM uses a different combination of pre-training, train/test, and zero-shot datasets.
While we did not find a TSFM that used the same set for different purposes, we still argue that this multi-purpose use of datasets is a problem. For example, suppose Model A uses a dataset for pre-training, and Model B uses the same dataset for evaluation. In that case, it becomes impossible to evaluate these models against each other on that dataset fairly.
A notable case for the multi-purpose use of datasets is the subset Australian Electricity Demand from the Monash dataset collection (Godahewa et al., 2021), which is used for pre-training (e.g., Lag-Llama, Timer), train/test evaluation (e.g., Moirai), and zero-shot forecasting (e.g., Chronos, TimesFM). This multi-purpose use makes the dataset practically useless for benchmarking between models (e.g., it cannot be used to compare Chronos to Lag-Llama).
Another interesting case is the Informer collection (Zhou et al., 2021), which is primarily used for zero-shot evaluations (indicated by green connections in Figure 1). However, some individual time series of the Informer collection, such as ETTh1, ETTh2, and ETTm1, have also been used for pre-training (e.g., in LagLlama) or train/test evaluation (e.g., in UniTime). Again, this means that LagLlama or UniTime cannot be benchmarked against any other model using the Informer collection. Note that it is not easy to exactly determine which datasets a model has been trained on. It requires careful reading of the original paper and sometimes even analyzing published source code. Doing this for a dozen or more models in a benchmarking study requires substantial effort.
No dataset has been universally used across all TSFMs for train/test or zero-shot evaluation (see Table 2 and Table 3). Considering all investigated datasets, only 7% have never been used for pre-training or train/test fine-tuning, making them the only suitable candidates for genuine zero-shot model benchmarking.
Renaming datasets is another issue we encountered during our analysis. For example, the Monash dataset "Elecdemand" is a scaled subset ($1/1000$) of the "Australian Electricity Demand" dataset. Normalization practices could introduce unintended information leakage if these datasets are treated as separate. Similarly, the dataset "ElectricityLoadDiagrams20112014" appears under various names, such as "Electricity" in the Autoformer and Monash collections (Godahewa et al., 2021; Wu et al., 2022) and "ECL" in the Informer (Zhou et al., 2021) collection. Dataset collections sometimes also contain the same variables but different observations. For instance, the car ride shares dataset is included in both the Monash and GluonTS collections. Although both repositories cite the same GitHub source [^4] https://github.com/fivethirtyeight/uber-tlc-foil-response, a close examination reveals that they represent different time series of distinct For-Hire Vehicles.
The empirical analysis using the Moirai TSFM by Aksu et al. (2024) demonstrates the effects of information leakage between training and test sets on the predictive performance of TSFMs. The authors prepared a new pre-training and held-out evaluation dataset without any overlap and trained the Moirai architecture on it. A different Moirai model from a previous publication was taken which pre-training dataset contained 0.1% data of the newly defined held-out evaluation dataset leading to a deliberate small portion of information-leakage. They then compared the performance of the models with different numbers of parameters and different forecasting horizons. The models trained on the training set with information leakage achieved an average MAPE between 7.6 and 59 percentage points lower than the same architecture trained on leakage-free training data (see Appendix B.2 for detailed results). These results vividly demonstrate the significant impact that test set contamination has on the performance of TSFMs. Larger model sizes especially seem prone to this, suggesting that bigger models tend to memorize rather than generalize. Another example of the information leakage effect is observable in a recently proposed TSFM benchmark. The benchmark creators, due to the TSFMs’ incomprehensible datasets, unintentionally included three evaluation datasets as test sets that had already been used in the pretraining of TimesFM, UniTS, and TTM (Li et al., 2024b). Based on an analysis of the benchmark results, this leads to an advantage of 47% - 184% lower mean squared error (MSE) rate compared to best models not pre-trained on the leaking datasets. The advantage of the best TSFM on non-leaked datasets is only between 0.3% and 14%. Of course, the effect should be evaluated using comparisons on the same datasets, but the performance benefit through information leakage appears to align with findings from the Moirai leakage example. This highlights the difficulties for benchmarking between TSFMs.
3.2 Challenge: Memorization of Global Patterns
Direct overlap between observations of the training and test sets is not the only cause for information leakage. A more subtle case of information leakage can happen due to indirect dependencies between data points in the training and test sets (Kapoor & Narayanan, 2023). Typically, these indirect dependencies stem from confounding bias, that is, an external factor independently associated with observations in both datasets. In the context of time series forecasting, global events or crises like COVID-19 or the European energy crisis, caused by Russia’s invasion of Ukraine, can influence a large number of time series from different domains.
A key strength of TSFMs is learning patterns and transferring a suitable pattern to an unseen time series. But when a TSFM is trained on data including such global events or crises, information may inadvertently carry over to forecasts on test set time series from the same period and affected by the same external factors. This introduces a problematic dependency between the training and test data, violating the assumption of independence that is crucial for robust evaluation. As a result, the model gains access to information during training that is usually inaccessible in a real-world scenario at the time of prediction, leading to inflated performance metrics. The time component and the actual information inside a time series plays a new role for globally trained models such as TSFMs.
Figure 2: Global pattern memorization of TSFM A when trained on time series 1 and zero-shot forecasting time series 2.
3.3 Challenge: Evaluation Dimensions
The typical evaluation in time series forecasting is performed across the time dimension to assess a model’s capabilities over time (Hyndman & Athanasopoulos, 2021), particularly its performance across different seasonal patterns (Figure 3, left side). Using time-series cross-validation, one can verify whether a model provides reliable forecasts, for example during winter and summer. It should be emphasized that a rolling evaluation does not necessarily require retraining the model for each rolling interval. One can just as well use a fixed training dataset and then roll inference over the various test sets (Hewamalage et al., 2023), making it a suitable evaluation strategy for TSFMs.
Figure 3: Evaluation strategies along the time or the domain dimension.
Interestingly, current TSFMs seem to evaluate only using a single time split per series (Figure 3, right side) without performing time series cross-validation, usually to reduce resource utilization or costs (Das et al., 2024; Gruver et al., 2024). Results are then often reported based on a single prediction per time series (Ansari et al., 2024; Das et al., 2024; Shi et al., 2024). They demonstrate foundation model capabilities as the models are evaluated against multiple time series from different domains (see Section 2). But we have not found any TSFM for which a combined time and domain cross-validation is explicitly conducted (Figure 4), even though this approach would easily increase the number of predictions and ultimately lead to a more meaningful evaluation.
Figure 4: Evaluation along time and domain dimension.
3.4 Challenge: Domain Representative Time Series
Foundation models in particular aim to be generalizable across different domains. Still, there is no systematic derivation of why domains such as energy, transportation, weather, or web trends in particular should be predicted. A particularly illustrative example is the ETT dataset (Zhou et al., 2021), which is used in many TSFMs as a zero-shot benchmark (see Section 3.1). However, this dataset is based on two years of measurements from just two transformers in China. The exact location of these transformers within China is not specified, and only very general information about their origin is provided. Moreover, while the dataset indeed captures some noteworthy characteristics, such as both short- and long-term seasonal patterns, it still consists of only two specific time-series use cases among an uncountable variety of possible temporal dynamics observed in real-world applications. It is therefore unclear to what extent results obtained on ETT can be considered generalizable, whether to transformers deployed in other regions such as Africa, America, Europe, or even to a third transformer within China. Accordingly, it remains questionable whether the dataset allows for any broader conclusions about the overall capability of models in a special domain (Hewamalage et al., 2023).
4 Current Benchmarking Approaches
One could argue that existing best practices for model training and testing are sufficient to mitigate risks such as test set contamination and global pattern memorization. But are these strategies still adequate in the age of TSFMs, or must they be adapted, combined, or restructured?
Existing strategies can be broadly categorized into benchmarking strategies and research best practices. As the dominant logic of time series benchmarking (Hyndman & Athanasopoulos, 2021), time-based train/test splits, including those in the Monash dataset collection (Godahewa et al., 2021) and TFB (Qiu et al., 2024), prevent leakage within a single TSFM. However, it is still unsuitable for fair benchmarking, as teams may choose different training and testing sets (see Section 2). Moreover, such predefined train/test splits typically do not use a global split point across all series in a collection, which would protect against information leakage through memorizing global patterns. For example, the Monash collection series uses varying start and end points, preventing the possibility of a global time-based split across all included time series (see Appendix 7).
In terms of foundation models, held-out datasets like GIFT-Eval (Aksu et al., 2024) focus on creating curated pre-training and test datasets without any overlap to prevent test set contamination. This approach enables a robust zero-shot forecasting evaluation, which is critical to TSFMs. However, the held-out dataset approach only applies to TSFMs developed in the future, as past models cannot be benchmarked reliably due to potential data contamination caused by overlap between their pre-training data and the held-out dataset. Furthermore, in case researchers want to include more data into pre-training than provided in a collection like GIFT-Eval, it requires careful checks for potential test set contamination in zero-shot forecasting. Lastly, global patterns like events or crises can still affect both the training and held-out test sets of such collection as long as no global time-based split is enforced.
Combining held-out datasets, global time-based train/test splits and retraining and benchmarking all TSFMs on a standard pre-training and test set would prevent information leakage and enable fair benchmarking. Yet, retraining all model architectures used in a benchmark is obviously very resource-intensive and prone to implementation errors. In addition, forcing all models to use the same data for pre-training mitigates the advantages of the scaling laws (Edwards et al., 2024).
Using anonymized real-world benchmarking test datasets (e.g., masking or scaling them) can also help to address information leakage concerns. Anonymizing real-world time series would help prevent the simple recall of original data points in case they have been included in a model’s pre-training set (Der et al., 2023). However, pre-processing techniques like scaling, feature engineering, and tokenization, which are usually implemented in a TSFM architecture, could potentially reverse such anonymization efforts, leading to the original or a very similar time series. Hence, the effectiveness of this approach is questionable and would require sophisticated anonymization techniques that preserve statistical properties while preventing direct recalling of data points.
Competition-based splits, like in the M-competitions (Makridakis et al., 2022), require forecasts on unpublished time-splits of known series. Such competitions start by publishing the training set while the test set remains private. Forecasts are registered during the competition, usually organized via Kaggle [^5] https://www.kaggle.com/competitions/m5-forecasting-accuracy/overview and metrics are calculated using the private test set. Although these approaches prevent test set contamination by registering forecasts based on time-based splits, they do not ensure a rigorous held-out dataset split for zero-shot forecasting. Furthermore, the hidden test set usually contains static, past data for evaluation throughout the competition. Besides enabling TSFMs to benefit from global pattern memorization, organizing such competitions requires substantial resources for administration, leading to long waiting periods between competitions. In a more innovative way, the recent M6 competition employed a strategy of using real-world live data from 100 financial assets and registering predictions into the real future (Makridakis et al., 2024). The main task was not forecasting but ranking the assets according to their future values. Still, it was a closed competition in a predefined time frame where long waiting period persists for results as with previous competitions which can be a hurdle in the fast developments of TSFMs.
In contrast, synthetic data for testing promises an unlimited amount of truly novel and so far unknown data. In fact, for pre-training, synthetic data has already been used successfully (Dooley et al., 2023). Yet, the external validity of solely using synthetic data for testing remains questionable, as synthetic data generation itself needs to be evaluated (Yuezhang et al., 2021).
Research best practices such as pre-registration for predictive modeling provide indirect mechanisms to mitigate risks of test set contamination due to multi-purpose use of datasets or global pattern memorization. Pre-registration for predictive modeling has been proposed as a rigorous research method originating from the experimental sciences (e.g., medicine, psychology) (Hofman et al., 2023). This method requires an ex-ante documentation and publication of the whole predictive modeling process before model training, including all planned data collection and preparation as well as modeling and evaluation activities. Model cards can complement this ex-post, by providing all information on data, training, and benchmarking to other researchers in one document (Mitchell et al., 2019). While these best research practices have the potential to prevent cherry picking of training and test sets and evaluation metrics and provide deeper insights into the models, they cannot mitigate the risks stemming from multi-purpose use of datasets and memorization of global patterns.
Lastly, the academic review process ensures proper data splits and detects any information leakage. While the peer review process has proven its ability to spot methodological issues, relying solely on this approach may not comprehensively prevent unintentional benchmark contamination (Li et al., 2024b). Furthermore, reviewers indeed cannot assess the validity of all models included in a benchmarking study. Hence, as witnessed in LLM evaluation research, academic review processes may oversee test set contaminations (see Section 2).
In sum, while pre-defined distinctions between training, test, and zero-shot sets improve evaluation integrity, they hinder comparability with prior models and mitigate scaling laws. Adhering to best practices may help identify test set contamination, but cannot entirely prevent it, especially in the presence of global pattern memorization. Given the rapid evolution of TSFMs, waiting long periods between competitions is increasingly impractical and risks stagnating meaningful progress, while anonymization of data cannot fully mitigate TSFM benchmarking challenges.
5 TSFM Benchmark Requirements
The essence of time series forecasting is to predict the future. Paradoxically, using data from the past to evaluate their ability to do so is standard practice as we have seen in previous Section 4. Before the advent of TSFMs, this paradox was not a serious issue, as traditional time series models have to be trained on past values of the same time series they are expected to continue. The only pitfall to avoid is adhering to the correct ordering of training and test observations. Yet, as we have outlined in Section 2, the possibility of pre-training TSFMs on vast collections of time series from different domains and using them without training on the target time series introduces new risks of information leakage, possibly resulting in test set contamination and memorization of global patterns. Now there are a variety of ways (Section 4) to test and compare TSFM, but in our opinion none of them fully solves the problems identified (Section 2).
Instead of only pointing out which benchmarking methodologies are not suitable for TSFMs, we want to be proactive and develop guidelines how a TSFM benchmark should be designed. Based on the advantages and disadvantages of the current benchmarking approaches (Section 4) combined with our analysis (Section 2), we propose the following requirements for a time series foundation model benchmark methodology:
  1. 1.
    Prevention of Information Leakage. The benchmark must guarantee that no portion of test data is present within any model’s training corpus.
  2. 2.
    Unambiguous Data Splits and Transparency. It should employ clear and standardized protocols for dividing data into training and test sets.
  3. 3.
    Time- and Domain Dimension Evaluation. It should apply cross-validation across both time and domain dimensions, verifying model consistency across seasonal periods (e.g., summer/winter) and across diverse datasets or domains to ensure generalization and robustness.
  4. 4.
    Retrospective and Prospective Evaluation. The methodology should not be limited to models developed after the introduction of the benchmark, but must also be applicable to existing models and future models as the field evolves.
  5. 5.
    Flexibility in Pre-Training Data. Researchers should not face artificial restrictions on the amount, provenance, or specificity of pre-training data.
  6. 6.
    Mitigation of Global Pattern Memorization. Mechanisms must be in place such as global temporal cut-off points to reduce the risk of models exploiting global patterns.
  7. 7.
    Scalability and Sustainability. The framework should remain practical for long-term community use, avoiding excessive computational, infrastructural, or administrative requirements.
  8. 8.
    Real-World Representativeness. Evaluation data should reflect the diversity of time series in relevant real-world domains and frequencies.
  9. 9.
    Support for Zero-Shot and Fine-Tuned Evaluation. The protocol should enable assessment of models in both zero-shot and fine-tuned settings.
  10. 10.
    Inclusivity and Accessibility. Participation should be open to both academic and industry teams, including those using open-source or closed-source models.
  11. 11.
    Robustness Against Gaming and Overfitting. Safeguards must be implemented to prevent exploitation of the test data, such as rotating test series, or anonymizing series during evaluation.
Now, one could argue that these are wishful requirements for a benchmark, but that they are unrealistic to fulfill all and compromises must be made. We would like to challenge precisely this and encourage the community to look for solutions that cover as wide a range of requirements as possible.
For example, a benchmark based entirely on constantly generated synthetic data could meet more than half of the 11 requirements. While there has been impressive progress in mimicking real-world data (Narasimhan et al., 2024), a question arises: as the generation process becomes more similar to actual historical data, does the risk of information leakage or memorizing global patterns increase? On the other hand, if the generation process is more independent of past data, we once again encounter the issue of lacking real-world relevance. Additionally, such synthetic data may be vulnerable to gaming if someone can reverse-engineer the data generation process. Nevertheless, it is worth considering this approach further.
To more comprehensively address the spectrum of requirements, a conceptual shift in how TSFM benchmarks are designed might be necessary. Imagine a framework built around a continuously advancing global temporal split. In such a system, TSFMs would be tasked with forecasting actual future events, with evaluations performed retrospectively as real-world data becomes available. This dynamic approach, where the “test set" is always the yet-unseen future, inherently addresses critical issues like information leakage from test set contamination or rote memorization. This evaluation seems especially applicable for time-series forecasting, which is uniquely concerned with future prediction unlike typical NLP or computer vision tasks.
Furthermore, since the test data would not exist during model training, restrictions on training data could be completely lifted, facilitating the evaluation of a broader range of models, including closed-source TSFMs. Still, this paradigm introduces new complexities. For instance, distinguishing purely zero-shot forecasts from fine-tuned forecasts could be difficult and leading to nonfulfillment of requirement 9. Here it might require complementary ideas. To further promote transparency regarding training data, the adoption of model cards could be encouraged as a best practice, though community acceptance, as seen with LLMs, might present a hurdle.
Undoubtedly, operationalizing such a benchmark would present significant implementation challenges, likely necessitating a dedicated platform. However, we posit that pursuing this line of thinking offers a promising avenue to resolve many of the current impasses in TSFM evaluation. We therefore call upon the research community to collaboratively explore and develop these concepts further, shaping a more robust and reliable future for time series forecasting benchmarks.
6 Conclusion
Remarkably, TSFMs appear to follow a development trajectory similar to that of LLMs in terms of both data and model size. Scaling laws seem to apply to TSFMs, and there are notable similarities in the evaluation challenges they face, particularly concerning potential information leakage. We highlight key evaluation challenges that arise from the emergence of TSFMs. By investigating the data lineage of 15 prominent TSFMs and reflecting the common TSFM training and evaluation strategy, we identified two significant sources of information leakage: test set contamination through multi-purpose use of datasets and memorization of global patterns induced by external events or crises. Furthermore, we question the current evaluation strategy that considers only the domain dimension while disregarding the temporal dimension, as well as the selection of time series used to represent a domain.
Current benchmark approaches from time-series or machine learning do not mitigate all information leakage risks. Based on our investigations and existing methodologies, we propose requirements for a benchmark in the age of TSFMs. We would like to encourage the community to develop new ideas which can fulfill the requirements like benchmarking on real future data. Besides the development of a benchmark methodology, we see several directions for future research that could also reinterpret our current results. The effect size of test set contamination (when a test dataset was present in the pretraining corpus) has only been proven to be significant in isolated cases so far. The magnitude of global pattern memorization is also largely unknown and is likely to vary depending on the domain, type of global event, and other factors. Carefully designed training regimes for TSFMs will be required to properly quantify these effects.
Ultimately, we believe that the criteria we have established lay the foundation for robust, fair, and information-leakage-free benchmarking. This is essential for preventing an evaluation crisis for TSFMs akin to what has been observed with LLMs.
Appendix A Time Series Foundation Model Analysis
A.1 Overlap in test data usage
Table 2: Comparison of train/test evaluations of selected different models ordered by the total number of overlaps. For full results see Section C.
Dataset Chronos Moirai MOMENT Time-LMM
Electricity (H) T/T - T/T T/T
Traffic Hourly (H) - T/T T/T T/T
ETT1 (H) - - T/T T/T
Weather (Australia) (D) - T/T - T/T
M4 Monthly (M) T/T T/T - -
M4 Hourly (H) T/T T/T - -
M4 Daily (D) T/T T/T - -
M4 Weekly (W-SUN) T/T T/T - -
ETT1 (15T) - - T/T T/T
ETT2 (15T) - - T/T T/T
ETT2 (H) - - T/T T/T
CDC Fluview ILINet (W) - - T/T T/T
Table 3: Comparison of zero-shot evaluations of selected different models ordered by the total number of overlaps. Only datasets with 5 or more overlaps are shown. Models which have not included any of the shown datasets into their train/test evaluation are not displayed. For full results see Section C.
Dataset Chronos Moirai Moirai-MoE Lag-Llama LLMTime
Jena Weather (10T) - ZS ZS - ZS
ETT2 (15T) ZS ZS - ZS -
Electricity (H) - ZS ZS - ZS
ETT2 (H) ZS ZS - - -
ETT1 (H) ZS ZS - - -
ETT1 (15T) ZS ZS - - -
Traffic Hourly (H) ZS - - - ZS
A.2 Data Analysis
Additional visualizations based on the analysis of Time Series Foundation Models training and evaluation data.
Figure 5: Most common frequencies in investigated datasets.
Appendix B Leakage Investigations
B.1 Potential global pattern leakage in dataset collections
Figure 6: Test set contaminations when comparing TSFM A with TSFM B with overlapping pre-training and test sets on the test set of TSFM A.
It may be theoretically possible that existing model evaluation results have already been influenced by global pattern leakage. We analyse the overlapping time frames in the Monash Dataset  [^6] https://huggingface.co/datasets/Monash-University/monash_tsf as can be seen in Figure 7. Test datasets shouldn’t lie in the same time period as other training datasets. As an example, the proposed Australian Electricity Demand test time frame (April 2015) lays in the same period as the training period of the "Weather (Australia)" training set (1900-2021). The Moirai and Moirai-MoE models use both datasets for train/test forecasts, meaning if their are any global patterns in the weather data which could influence the Australian Electricity Demand data, this would be a kind of information leakage. Similarly, the Oikolab Weather dataset ranges from 2010 January to 2021 May, which is weather data located in Melbourne. The test set of the dataset Pedestrian Count (also Melbourne) is located at the end of April 2020. This test period lays within the training data of the Oikolab Weather, which could be a greater potential global pattern information leakage than in previous example. Both datasets are included as well in the train/test datasets of Moirai and Moirai-MoE.
Moreover, LagLlama uses the Beijing Multi-Site Air Quality dataset from 12 districts in Beijing from 2013 to 2017 as a pre-training dataset and simultaneously uses the Beijing PM2.5 air quality dataset from the Beijing airport between 2010 and 2013.
Figure 7: Date range of datasets within Monash dataset collection. Note that the “cif_2016" dataset was originally delivered without a timestamp and the start year was defined as 1900. No data cleaning was done.
B.2 Leakage in Moirai Model
Table 4: Information leakage in TSFM comparing Moirai with (MoiLeak) and without (Moi) information leakage during training phase for different model sizes. Metric is MAPE. Based on experiments by Aksu et al. (2024).
Horizon Short Medium Long
Moirai Leakage Moirai Moirai Leakage Moirai Moirai Leakage Moirai
Dataset Frequency Model Size
loop_seattle 5T S 0.84 0.87 0.75 0.77 0.70 0.75
B 0.67 0.84 0.42 0.83 0.50 0.78
L 0.66 0.83 0.33 0.85 0.46 0.81
loop_seattle H S 1.22 1.19 0.73 0.70 0.70 0.71
B 0.96 1.08 0.54 0.65 0.49 0.59
L 0.84 0.89 0.53 0.71 0.47 1.18
m_dense H S 0.73 0.70 0.70 0.71 0.71 0.91
B 0.54 0.65 0.49 0.59 0.53 0.61
L 0.53 0.71 0.47 1.18 0.49 1.69
sz_taxi 15T S 0.95 1.11 0.65 0.60 2.12 2.30
B 0.90 0.84 0.71 0.64 2.16 2.42
L 0.78 0.82 0.71 0.60 2.14 2.24
The investigated data includes 9 datasets on short horizons, four datasets each on medium (10x short horizon) and long horizons (15x short horizon). We filtered the results to 4 datasets where data is available for all horizons to reasonably compare the metrics and model sizes, as shown in Figure 8. On short horizons, both models perform better with increasing model size while Moirai Leakage has on average 7.6% points lower MAPE score. At medium horizons, the leakage-free MoiraiL generally underperforms compared to other model sizes, exhibiting an average performance deficit of 32% points compared to the Moirai LeakageL model. On long horizons, all models perform significantly worse. However, Moirai Leakage achieves an average MAPE score that is 29% points lower than the leakage-free Moirai. Across all horizons, the biggest difference is between MoiraiL and Moirai LeakageL, similar to typical LLM memorization behavior, when large models better memorize pre-train data than smaller models.
Figure 8: Comparing the model Moirai once with information leakage (Moirai Leakage), where the model was exposed to the test data during pre-training and the same model trained without information leakage (Moirai). Short, Medium and Long forecasting horizons were investigated. Model sizes are S (small), B (base) and L (large) and compared by the MAPE metric (the lower the better). Based on experiments by Aksu et al. (2024).
Appendix C TSFM Data Lineage
Table 5: Analysis of datasets used by investigated TSFMs with Pre-training (P), train-test splits (T/T) and test (T).
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
Jena Weather Autoformer (Wu et al., 2021) Nature 10T ZS ZS T/T T/T ZS ZS ZS ZS T/T
Jena Weather Autoformer (Wu et al., 2021) Nature H
Jena Weather Autoformer (Wu et al., 2021) Nature D ZS
BizITObs - Application AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S ZS
BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps 5T ZS
BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps H
BizITObs - Service AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S
BDG-2 Bear BuildingsBench (Emami et al., 2023) Energy H P P P
BDG-2 Fox BuildingsBench (Emami et al., 2023) Energy H P P P
BDG-2 Panther BuildingsBench (Emami et al., 2023) Energy H P P P
BDG-2 Rat BuildingsBench (Emami et al., 2023) Energy H P P P
Borealis BuildingsBench (Emami et al., 2023) Energy H P P P
Buildings900K BuildingsBench (Emami et al., 2023) Energy H P P P
IDEAL BuildingsBench (Emami et al., 2023) Energy H P P P
Low Carbon London BuildingsBench (Emami et al., 2023) Energy H P P
Sceaux BuildingsBench (Emami et al., 2023) Energy H P P P
SMART BuildingsBench (Emami et al., 2023) Energy H P P P
CMIP6 ClimateLearn (Nguyen et al., 2023) Climate 6H P P P
ERA5 ClimateLearn (Nguyen et al., 2023) Climate H P P P P
Alibaba Cluster Trace 2018 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P
Azure vM Traces 2017 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P
Borg Cluster Data 2011 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P
Air Passengers Darts Dataset (Herzen et al., 2021) Econ/Fin M ZS ZS
Aus Beer Darts Dataset (Herzen et al., 2021) Econ/Fin Q ZS ZS
Gas Rate CO2 Darts Dataset (Herzen et al., 2021) Econ/Fin Y ZS ZS
Monthly Milk Darts Dataset (Herzen et al., 2021) Econ/Fin M ZS ZS
Wine Darts Dataset (Herzen et al., 2021) Sales M ZS ZS
Wooly Darts Dataset (Herzen et al., 2021) Econ/Fin Q ZS ZS
Heart Rate Darts Dataset (Herzen et al., 2021) Nature 0.5S ZS ZS
Sunspots Darts Dataset (Herzen et al., 2021) Nature M ZS
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H
FRED FRED (Oreshkin et al., 2021) Econ/Fin D P
FRED FRED (Oreshkin et al., 2021) Econ/Fin Y P
FRED FRED (Oreshkin et al., 2021) Econ/Fin W P
FRED FRED (Oreshkin et al., 2021) Econ/Fin M P
FRED FRED (Oreshkin et al., 2021) Econ/Fin Q P
M5 GluonTS (Alexandrov et al., 2020a) Sales D ZS P P P
Taxi GluonTS (Alexandrov et al., 2020a) Transport 30T T/T P P P
Taxi GluonTS (Alexandrov et al., 2020a) Transport H P P
Uber TLC Daily GluonTS (Alexandrov et al., 2020a) Transport D T/T P P P
Uber TLC Hourly GluonTS (Alexandrov et al., 2020a) Transport H T/T P P P P
Wiki-Rolling GluonTS (Alexandrov et al., 2020a) Web D P
Trends hourly Google Trends Web H P
Trends weekly Google Trends Web W P
Trends monthly Google Trends Web D P
Requests Huawei Cloud (Joosen et al., 2023) CloudOps T ZS
Function Delay Huawei Cloud (Joosen et al., 2023) CloudOps T P
Platform Delay Huawei Cloud (Joosen et al., 2023) CloudOps T ZS
CPU Usage Huawei Cloud (Joosen et al., 2023) CloudOps T P
Memory Usage Huawei Cloud (Joosen et al., 2023) CloudOps T P
CPU Limit Huawei Cloud (Joosen et al., 2023) CloudOps T P
Memory Limit Huawei Cloud (Joosen et al., 2023) CloudOps T P
Instances Huawei Cloud (Joosen et al., 2023) CloudOps T P
ETT1 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T
ETT1 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T
ETT1 Informer (Zhou et al., 2020) Energy D ZS ZS
ETT1 Informer (Zhou et al., 2020) Energy W-THU
ETT2 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T ZS T/T ZS ZS ZS T/T T/T T/T
ETT2 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T
ETT2 Informer (Zhou et al., 2020) Energy D ZS ZS
ETT2 Informer (Zhou et al., 2020) Energy W-THU
Weather (US) Informer (Zhou et al., 2020) Weather H ZS
Brazilian Cities Temperature Kaggle Competition Weather M P
Walmart Store Sales Kaggle Competition Sales W ZS ZS ZS
Istanbul Traffic Kaggle Competition Transport H ZS ZS ZS
Turkey Power Kaggle Competition Energy H ZS ZS ZS ZS
TSMC Stock Kaggle Competition Econ/Fin D ZS
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
LargeST LargeST (Liu et al., 2023a) Transport 5T P P P
Beijing Subway LibCity (Wang et al., 2023a) Transport 30T P P P
HZMetro LibCity (Wang et al., 2023a) Transport 15T P P P
Loop Seattle LibCity (Wang et al., 2023a) Transport 5T P P P
Loop Seattle LibCity (Wang et al., 2023a) Transport H
Loop Seattle LibCity (Wang et al., 2023a) Transport D
Los-Loop LibCity (Wang et al., 2023a) Transport 5T P P P P
M_DENSE LibCity (Wang et al., 2023a) Transport H
M_DENSE LibCity (Wang et al., 2023a) Transport D ZS
PEMS Bay LibCity (Wang et al., 2023a) Transport 5T P P P P
PEMS03 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P
PEMS04 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P
PEMS07 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P
PEMS08 LibCity (Wang et al., 2023a) Transport 5T P P T/T ZS P P
Q-Traffic LibCity (Wang et al., 2023a) Transport 15T P P P P P
SHMetro LibCity (Wang et al., 2023a) Transport 15T P P P P
SZ-Taxi LibCity (Wang et al., 2023a) Transport 15T P P P
SZ-Taxi LibCity (Wang et al., 2023a) Transport H P
Rotterdam LibCity (Wang et al., 2023a) Transport 2T P P
Beijing Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P P
CDC Fluview ILINet LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T ZS T/T T/T P
CDC Fluview WHO NREVSS LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T P
China Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P
Favorita Sales LOTSA_Others (Woo et al., 2024) Sales D P P P P
Favorita Transactions LOTSA_Others (Woo et al., 2024) Sales D P P P
GoDaddy LOTSA_Others (Woo et al., 2024) Econ/Fin M P P
KDD Cup 2022 LOTSA_Others (Woo et al., 2024) Energy 10T P P P
Project Tycho LOTSA_Others (Woo et al., 2024) Healthcare W P P P
Residential Load Power LOTSA_Others (Woo et al., 2024) Energy T P P P
Residential PV Power LOTSA_Others (Woo et al., 2024) Energy T P P P
Restaurant LOTSA_Others (Woo et al., 2024) Sales D P
Solar LSTNet (Lai et al., 2017) Energy 10T
Solar LSTNet (Lai et al., 2017) Energy 5T P
Solar LSTNet (Lai et al., 2017) Energy H P ZS ZS ZS
Solar LSTNet (Lai et al., 2017) Energy D
Solar LSTNet (Lai et al., 2017) Energy W-FRI
Exchange Rate LSTNet (Lai et al., 2017) Econ/Fin D ZS ZS T/T ZS ZS T/T P
Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales D P
Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales W-WED
Australian Electricity Demand Monash (Godahewa et al., 2021) Energy 30T ZS ZS T/T T/T P P P ZS P ZS P
Bitcoin Monash (Godahewa et al., 2021) Econ/Fin D ZS T/T T/T P P ZS P ZS P
Car Parts Monash (Godahewa et al., 2021) Sales M T/T T/T P ZS P
CIF 2016 Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P
COVID Deaths Monash (Godahewa et al., 2021) Healthcare D ZS ZS T/T T/T P ZS ZS P
Covid Mobility Monash (Godahewa et al., 2021) Transport D T/T T/T P
Elecdemand Monash (Godahewa et al., 2021) Energy 30T T/T T/T P
Extended Web Traffic Monash (Godahewa et al., 2021) Web D T/T T/T P P P
FRED MD Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P
Hospital Monash (Godahewa et al., 2021) Healthcare M ZS ZS T/T T/T P ZS ZS P
Kaggle Web Traffic Weekly Monash (Godahewa et al., 2021) Web W P P P P
KDD Cup 2018 Monash (Godahewa et al., 2021) Nature H T/T T/T T/T P P P ZS P P
KDD Cup 2018 Monash (Godahewa et al., 2021) Nature D
London Smart Meters Monash (Godahewa et al., 2021) Energy 30T T/T P P P P P P P
M1 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T P ZS P
M1 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P P P
M1 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P P P
M3 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T ZS T/T ZS P
M3 Other Monash (Godahewa et al., 2021) Econ/Fin Q T/T T/T T/T ZS
M3 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P ZS T/T P
M3 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P ZS T/T P
M4 Daily Monash (Godahewa et al., 2021) Econ/Fin D T/T P T/T T/T P T/T ZS P
M4 Hourly Monash (Godahewa et al., 2021) Econ/Fin H T/T P T/T T/T P T/T ZS P
M4 Monthly Monash (Godahewa et al., 2021) Econ/Fin M T/T P T/T T/T ZS T/T ZS P
M4 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q-DEC ZS P P P ZS T/T P
M4 Weekly Monash (Godahewa et al., 2021) Econ/Fin W-SUN T/T P T/T T/T T/T ZS P
M4 Yearly Monash (Godahewa et al., 2021) Econ/Fin A-DEC ZS P P P ZS T/T P
NN5 Daily Monash (Godahewa et al., 2021) Econ/Fin D ZS ZS T/T T/T P ZS P ZS P
NN5 Weekly Monash (Godahewa et al., 2021) Econ/Fin W ZS ZS T/T T/T P ZS ZS
Oikolab Weather Monash (Godahewa et al., 2021) Climate H T/T T/T P
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
Pedestrian Counts Monash (Godahewa et al., 2021) Transport H T/T ZS T/T T/T P ZS P ZS ZS P
Rideshare Monash (Godahewa et al., 2021) Transport H T/T T/T T/T P ZS P
Saugeen Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P ZS P ZS P
Saugeen Monash (Godahewa et al., 2021) Nature W-THU
Saugeen Monash (Godahewa et al., 2021) Nature M
Solar Power (Australia) Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P
Solar Monash (Godahewa et al., 2021) Energy W ZS P P ZS
Solar Monash (Godahewa et al., 2021) Energy 10M P P P
Sunspot Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P P ZS P ZS P
Temperature Rain Monash (Godahewa et al., 2021) Nature D T/T T/T T/T P P ZS P
Tourism Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS P T/T ZS ZS P
Tourism Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS ZS P T/T ZS ZS P
Tourism Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS ZS P P P T/T ZS P
Traffic Hourly Monash (Godahewa et al., 2021) Transport H ZS P T/T T/T T/T P T/T ZS ZS ZS ZS T/T P
Traffic Weekly Monash (Godahewa et al., 2021) Transport W ZS T/T T/T ZS P ZS ZS P
US Births Monash (Godahewa et al., 2021) Healthcare D ZS T/T T/T P P ZS P ZS P
US Births Monash (Godahewa et al., 2021) Healthcare W-TUE
US Births Monash (Godahewa et al., 2021) Healthcare M
Vehicle Trips Monash (Godahewa et al., 2021) Transport D T/T T/T P ZS P
Weather (Australia) Monash (Godahewa et al., 2021) Climate D ZS ZS T/T T/T P ZS ZS P T/T T/T P
Wind Farms Monash (Godahewa et al., 2021) Energy T T/T T/T P P
Wind Farms Monash (Godahewa et al., 2021) Energy H P P
Wind Farms Monash (Godahewa et al., 2021) Energy D P
Wind Farms Monash (Godahewa et al., 2021) Energy 5T P
Wind Power Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P
Kaggle Web Traffic Daily Monash (Godahewa et al., 2021) Web D P
Restaurant Recruit Other Sales D
Subseasonal Other Climate D P P P
Subseasonal Precipitation Other Climate D P P P
USHCN Other Weather D P
Mexico City Bikes Other Transport H P P
Dominick Other Sales D ZS P P
ERCOT Load Other Energy H ZS P
Weather (US) Other Weather 10T P ZS
Bike Sharing Other Transport H ZS ZS
CC Carbon Other Energy 2T ZS
Air Quality Other Nature H P
Beijing PM 2.5 Other Nature H ZS
TDBrain Other Healthcare 0.002S P
SensorData Other IoT 0.02S P
BDG-2 Bull ProEnFo (Wang et al., 2023b) Energy H P P P
BDG-2 Cockatoo ProEnFo (Wang et al., 2023b) Energy H P P P
BDG-2 Hog ProEnFo (Wang et al., 2023b) Energy H P P P
Covid19 Energy ProEnFo (Wang et al., 2023b) Energy H P P P
ELF ProEnFo (Wang et al., 2023b) Energy H P P
GEF12 ProEnFo (Wang et al., 2023b) Energy H P P P
GEF14 ProEnFo (Wang et al., 2023b) Energy H P P P
GEF17 ProEnFo (Wang et al., 2023b) Energy H P P
PDB ProEnFo (Wang et al., 2023b) Energy H P P P
Spanish ProEnFo (Wang et al., 2023b) Energy H P P P P
Australia Rainfall Tan et al., 2021 Nature H P
Beijing PM 2.5 Tan et al., 2021 Nature H P
BenzeneConcentration Tan et al., 2021 Nature H P
IEEEPPG Tan et al., 2021 Healthcare 0.008S P
BIDMC32HR Tan et al., 2021 Healthcare - P
Dodger TSB-UAD (Paparrizos et al., 2022) Traffic - P
ECG TSB-UAD (Paparrizos et al., 2022) Healthcare - P
IOPS TSB-UAD (Paparrizos et al., 2022) Web - P
KDD21 TSB-UAD (Paparrizos et al., 2022) Multiple - P
MGAB TSB-UAD (Paparrizos et al., 2022) Synthetic - P
NAB TSB-UAD (Paparrizos et al., 2022) Web - P
SensorScope TSB-UAD (Paparrizos et al., 2022) Nature - P
YAHOO TSB-UAD (Paparrizos et al., 2022) Web - P
NASA-MSL TSB-UAD (Paparrizos et al., 2022) Sensor - P
NASA-SMAP TSB-UAD (Paparrizos et al., 2022) Sensor - P
Daphnet TSB-UAD (Paparrizos et al., 2022) Gait - P
GHL TSB-UAD (Paparrizos et al., 2022) Sensor - P
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
Genesis TSB-UAD (Paparrizos et al., 2022) Sensor - P
MITDB TSB-UAD (Paparrizos et al., 2022) Healthcare - P
OPP TSB-UAD (Paparrizos et al., 2022) Sensor - P
Occupancy TSB-UAD (Paparrizos et al., 2022) Sensor - P
SMD TSB-UAD (Paparrizos et al., 2022) Web - P
SZSDB TSB-UAD (Paparrizos et al., 2022) Healthcare - P
Electricity UCI ML Archive (Trindade, 2015) Energy 15T T/T P T/T P
Electricity UCI ML Archive (Trindade, 2015) Energy H T/T P ZS ZS T/T P T/T ZS ZS ZS ZS T/T T/T
Electricity UCI ML Archive (Trindade, 2015) Energy D ZS
Electricity UCI ML Archive (Trindade, 2015) Energy W-FRI T/T
Electricity UCI ML Archive (Trindade, 2015) Energy M T/T
Electricity UCI ML Archive (Trindade, 2015) Energy W P P
SemgHandGenderCh2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
GestureMidAirD2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
UWaveGestureLibraryAll UCR Archive (Ismail Fawaz et al., 2019) Human - T/T
SelfRegulationSCP1 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
UWaveGestureLibraryX UCR Archive (Ismail Fawaz et al., 2019) Human - P
GesturePebbleZ2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
Healthcare5000 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - T/T
OSULeaf UCR Archive (Ismail Fawaz et al., 2019) Nature - T/T
MedicalImages UCR Archive (Ismail Fawaz et al., 2019) Healthcare - T/T
Haptics UCR Archive (Ismail Fawaz et al., 2019) Human - P
LargeKitchenAppliances UCR Archive (Ismail Fawaz et al., 2019) Power - P
Japanesevowels UCR Archive (Ismail Fawaz et al., 2019) Audio - P
Worms UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Ham UCR Archive (Ismail Fawaz et al., 2019) Facilities - T/T
DistalPhalanxTW UCR Archive (Ismail Fawaz et al., 2019) Nature - P
ProximalPhalanxOutlineCorrect UCR Archive (Ismail Fawaz et al., 2019) Nature - P
SemgHandMovementCh2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
RefrigerationDevices UCR Archive (Ismail Fawaz et al., 2019) Power - P
FreezerRegularTrain UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
PigAirwayPressure UCR Archive (Ismail Fawaz et al., 2019) Nature - P
TwoLeadECG UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
GunPointMaleversusFemale UCR Archive (Ismail Fawaz et al., 2019) Human - P
Trace UCR Archive (Ismail Fawaz et al., 2019) Power - P
SmoothSubspace UCR Archive (Ismail Fawaz et al., 2019) Generated - P
MiddlePhalanxTW UCR Archive (Ismail Fawaz et al., 2019) Nature - P
AtrialFibrillation UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
SyntheticControl UCR Archive (Ismail Fawaz et al., 2019) Generated - P
ShapesAll UCR Archive (Ismail Fawaz et al., 2019) Generated - P
Human UCR Archive (Ismail Fawaz et al., 2019) BodyZSerticalSignal - P
PLAID UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
AllGestureWiimoteX UCR Archive (Ismail Fawaz et al., 2019) Human - P
Heartbeat UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
Wafer UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
FaceFour UCR Archive (Ismail Fawaz et al., 2019) Generated - P
Phoneme UCR Archive (Ismail Fawaz et al., 2019) Audio - P P
InlineSkate UCR Archive (Ismail Fawaz et al., 2019) Human - P
CricketX UCR Archive (Ismail Fawaz et al., 2019) Human Body - P
SelfRegulationSCP2 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
DistalPhalanxOutlineCorrect UCR Archive (Ismail Fawaz et al., 2019) Nature - P
ChlorineConcentration UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Chinatown UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
GestureMidAirD1 UCR Archive (Ismail Fawaz et al., 2019) Human - P
MiddlePhalanxOutlineAgeGroup UCR Archive (Ismail Fawaz et al., 2019) Nature - P
UMD UCR Archive (Ismail Fawaz et al., 2019) Generated - P
Crop UCR Archive (Ismail Fawaz et al., 2019) Nature - P
PenDigits UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
GesturePebbleZ1 UCR Archive (Ismail Fawaz et al., 2019) Human - P
Handwriting UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
Mallat UCR Archive (Ismail Fawaz et al., 2019) Generated - P
ERing UCR Archive (Ismail Fawaz et al., 2019) Human - P
StarLightCurves UCR Archive (Ismail Fawaz et al., 2019) Nature - P P
WordSynonyms UCR Archive (Ismail Fawaz et al., 2019) Audio - P
PEMS-SF UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
FaceDetection UCR Archive (Ismail Fawaz et al., 2019) Human - P
Computers UCR Archive (Ismail Fawaz et al., 2019) Power - P
ArrowHead UCR Archive (Ismail Fawaz et al., 2019) Generated - P
Wine UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Coffee UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Earthquakes UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Herring UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
Lightning2 UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Beef UCR Archive (Ismail Fawaz et al., 2019) Nature - P
MiddlePhalanxOutlineCorrect UCR Archive (Ismail Fawaz et al., 2019) Nature - P
HealthcareFiveDays UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
Yoga UCR Archive (Ismail Fawaz et al., 2019) Human - P
Adiac UCR Archive (Ismail Fawaz et al., 2019) Nature - P
HandMovementDirection UCR Archive (Ismail Fawaz et al., 2019) Human - P
MoteStrain UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
Rock UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Strawberry UCR Archive (Ismail Fawaz et al., 2019) Nature - P
InsectWingbeatSound UCR Archive (Ismail Fawaz et al., 2019) Nature - P
DodgerLoopWeekend UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
MixedShapesSmallTrain UCR Archive (Ismail Fawaz et al., 2019) Generated - P
EthanolConcentration UCR Archive (Ismail Fawaz et al., 2019) Power - P
OliveOil UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Meat UCR Archive (Ismail Fawaz et al., 2019) Nature - P
MelbournePedestrian UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
Car UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
FaceAll UCR Archive (Ismail Fawaz et al., 2019) Human - P
FacesUCR UCR Archive (Ismail Fawaz et al., 2019) Human - P
AllGestureWiimoteY UCR Archive (Ismail Fawaz et al., 2019) Human - P
NATOPS UCR Archive (Ismail Fawaz et al., 2019) Human - P
SemgHandSubjectCh2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
ShakeGestureWiimoteZ UCR Archive (Ismail Fawaz et al., 2019) Human - P
Cricket UCR Archive (Ismail Fawaz et al., 2019) Human - P
BME UCR Archive (Ismail Fawaz et al., 2019) Generated - P
EigenWorms UCR Archive (Ismail Fawaz et al., 2019) Nature - P P
FordB UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
NonInvasiveFetalHealthcareThorax1 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
UWaveGestureLibrary UCR Archive (Ismail Fawaz et al., 2019) Human - P
CinCHealthcareTorso UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
PigArtPressure UCR Archive (Ismail Fawaz et al., 2019) Nature - P P
Fish UCR Archive (Ismail Fawaz et al., 2019) Nature - P
SonyAIBORobotSurface2 UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
FiftyWords UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
MotorImagery UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
ToeSegmentation1 UCR Archive (Ismail Fawaz et al., 2019) Human - P
PhonemeSpectra UCR Archive (Ismail Fawaz et al., 2019) Audio - P
FreezerSmallTrain UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
TwoPatterns UCR Archive (Ismail Fawaz et al., 2019) Generated - P
ShapeletSim UCR Archive (Ismail Fawaz et al., 2019) Generated - P
Plane UCR Archive (Ismail Fawaz et al., 2019) Generated - P
GestureMidAirD3 UCR Archive (Ismail Fawaz et al., 2019) Human - P
DiatomSizeReduction UCR Archive (Ismail Fawaz et al., 2019) Generated - P
Human UCR Archive (Ismail Fawaz et al., 2019) BodyHorizontalSignal - P
CricketZ UCR Archive (Ismail Fawaz et al., 2019) Human - P
StandWalkJump UCR Archive (Ismail Fawaz et al., 2019) Human - P
WormsTwoClass UCR Archive (Ismail Fawaz et al., 2019) Human - P
Lightning7 UCR Archive (Ismail Fawaz et al., 2019) Nature - P
UWaveGestureLibraryY UCR Archive (Ismail Fawaz et al., 2019) Human - P
GunPointAgeSpan UCR Archive (Ismail Fawaz et al., 2019) Human - P
DistalPhalanxOutlineAgeGroup UCR Archive (Ismail Fawaz et al., 2019) Nature - P
SwedishLeaf UCR Archive (Ismail Fawaz et al., 2019) Nature - P
LSST UCR Archive (Ismail Fawaz et al., 2019) Nature - P
CBF UCR Archive (Ismail Fawaz et al., 2019) Generated - P
BeetleFly UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Libras UCR Archive (Ismail Fawaz et al., 2019) Human - P
HouseTwenty UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
ScreenType UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
InsectEPGSmallTrain UCR Archive (Ismail Fawaz et al., 2019) Nature - P
AllGestureWiimoteZ UCR Archive (Ismail Fawaz et al., 2019) Sensor - P
DodgerLoopDay UCR Archive (Ismail Fawaz et al., 2019) Sensor - P
NonInZSasiZSeFetalHealthcareThorax2 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
BasicHuman UCR Archive (Ismail Fawaz et al., 2019) Bodys - P
GunPointOldversusYoung UCR Archive (Ismail Fawaz et al., 2019) Human - P
FordA UCR Archive (Ismail Fawaz et al., 2019) Sensor - P
InsectWingbeat UCR Archive (Ismail Fawaz et al., 2019) Nature - P
ItalyPowerDemand UCR Archive (Ismail Fawaz et al., 2019) Power - P
ProximalPhalanxOutlineAgeGroup UCR Archive (Ismail Fawaz et al., 2019) Nature - P
ACSF1 UCR Archive (Ismail Fawaz et al., 2019) Power - P
GunPoint UCR Archive (Ismail Fawaz et al., 2019) Human - P
RacketSports UCR Archive (Ismail Fawaz et al., 2019) Human - P
SmallKitchenAppliances UCR Archive (Ismail Fawaz et al., 2019) Power - P
ProximalPhalanxTW UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
DuckDuckGeese UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
PickupGestureWiimoteZ UCR Archive (Ismail Fawaz et al., 2019) Human - P
EthanolLevel UCR Archive (Ismail Fawaz et al., 2019) Power - P
SpokenArabicDigits UCR Archive (Ismail Fawaz et al., 2019) Audio - P
SonyAIBORobotSurface1 UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
HandOutlines UCR Archive (Ismail Fawaz et al., 2019) Human - P
PowerCons UCR Archive (Ismail Fawaz et al., 2019) Power - P
PhalangesOutlinesCorrect UCR Archive (Ismail Fawaz et al., 2019) Nature - P
BirdChicken UCR Archive (Ismail Fawaz et al., 2019) Nature - P
ToeSegmentation2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
PigCVP UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
CricketY UCR Archive (Ismail Fawaz et al., 2019) Human - P
FingerMovements UCR Archive (Ismail Fawaz et al., 2019) Human - P
ElectricDevices UCR Archive (Ismail Fawaz et al., 2019) Power - P
InsectEPGRegularTrain UCR Archive (Ismail Fawaz et al., 2019) Nature - P
DodgerLoopGame UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
Fungi UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Symbols UCR Archive (Ismail Fawaz et al., 2019) Generated - P
MixedShapesRegularTrain UCR Archive (Ismail Fawaz et al., 2019) Generated - P
ArticularyWordRecognition UCR Archive (Ismail Fawaz et al., 2019) Human - P
UWaveGestureLibraryZ UCR Archive (Ismail Fawaz et al., 2019) Human - P
Epilepsy UCR Archive (Ismail Fawaz et al., 2019) Human - P
Healthcare200 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
Weatherbench (Daily) Weather Bench Weather D P P
Weatherbench (Hourly) Weather Bench Weather H P P
Weatherbench (Weekly) Weather Bench Weather W P P
Wiki hourly Wikipedia Metrics Web H P
Wiki daily Wikipedia Metrics Web D P P P
Wiki weekly Wikipedia Metrics Web W P
Wiki monthly Wikipedia Metrics Web M P
Global Temp Corrformer (Wu et al., 2023) Weather 10T T/T
TSMixup 10M Chronos (Ansari et al., 2024) Synthetic - P
KernelSynth 1M Chronos (Ansari et al., 2024) Synthetic - P
TaxiBJ Zhang et al., 2017 Traffic and weather 30T ZS

### v2

@@
-Time Series Foundation Models: Benchmarking Challenges and Requirements
+Challenges and Requirements for Benchmarking Time Series Foundation Models
 Abstract
-Time Series Foundation Models (TSFMs) represent a new paradigm for time series forecasting, offering zero-shot forecasting capabilities without the need for domain-specific pre-training or fine-tuning.
-However, as with Large Language Models (LLMs), evaluating TSFMs is tricky, as with ever more extensive training sets, it becomes more and more challenging to ensure the integrity of benchmarking data.
-Our investigation of existing TSFM evaluation highlights multiple challenges, ranging from the representativeness of the benchmark datasets, over the lack of spatiotemporal evaluation, to risks of information leakage due to overlapping and obscure datasets, and the memorization of global patterns caused by external shocks like economic crises or pandemics.
-Our findings reveal widespread confusion regarding data partitions, risking inflated performance estimates and incorrect transfer of global knowledge to local time series.
-We argue for the development of robust evaluation methodologies to prevent pitfalls already observed in LLM and classical time series benchmarking, and call upon the research community to design new, principled approaches, such as evaluations on truly out-of-sample future data, to safeguard the integrity of TSFM assessment.
+Time Series Foundation Models (TSFMs) represent a new paradigm for time-series forecasting, promising zero-shot predictions without the need for task-specific training or fine-tuning.
+However, similar to Large Language Models (LLMs), the evaluation of TSFMs is challenging: as training corpora grow increasingly large, it becomes difficult to ensure the integrity of the test sets used for benchmarking.
+Our investigation of existing TSFM evaluation studies identifies two kinds of information leakage: (1) train-test sample overlaps arising from the multi-purpose reuse of datasets and (2) temporal overlap of correlated train and test series.
+Ignoring these forms of information leakage when benchmarking TSFMs risks producing overly optimistic performance estimates that fail to generalize to real-world settings.
+We therefore argue for the development of novel evaluation methodologies that avoid pitfalls already observed in both LLM and classical time-series benchmarking, and we call on the research community to adopt principled approaches to safeguard the integrity of TSFM evaluation.
 1 Introduction
-Time Series Foundation Models (TSFMs) are an emerging class of time series forecasting models inspired by the architecture and training procedures of foundation models in natural language processing, especially Large Language Models (LLMs).
-Prominent examples of TSFMs include Chronos (Ansari et al., 2024), TimesFM (Das et al., 2024), Moirai/Moirai-MoE (Woo et al., 2024; Liu et al., 2024b), MOMENT (Goswami et al., 2024) and Time-MoE (Shi et al., 2024).
-In contrast to traditional time series forecasting models, which need to be trained from scratch on the time series that should be forecasted, TSFMs use transfer learning.
-That is, they are pre-trained on massive amounts of generic and/or domain-specific time series and are, thus, able to produce zero-shot forecasts.
-However, pre-training TSFMs on massive amounts of time series data, typically scraped from publicly accessible online repositories, introduces challenges similar to those observed in LLM evaluation.
-The practice of training LLMs on much of the Internet has led to an LLM “evaluation crisis" (Liao & Xiao, 2023), describing a situation in which most benchmark data has already been exposed to an LLM during pre-training.
-Combining such test set contamination (Mirzadeh et al., 2024; Ravaut et al., 2024; Li et al., 2024a) with memorization effects (Chang et al., 2024) can result in overly-optimistic estimates of a model’s predictive performance and complicate the fair comparison of models through common benchmarks.
-Furthermore, as with LLMs (Chang et al., 2024), cross-validation in TSFMs is omitted due to its exhaustive resource requirements with large foundation models (Das et al., 2024; Gruver et al., 2024), and the diversity and generalizability of benchmarking datasets can be questioned (Aksu et al., 2024).
-As the awareness about evaluation challenges in TSFMs is growing, researchers have begun to propose mitigation strategies, such as held-out-driven benchmarks (Aksu et al., 2024) and clean train-test splits within benchmarks (Qiu et al., 2024) that include more diverse and generalizing test sets Aksu et al.
+Time Series Foundation Models (TSFMs) represent an emerging paradigm in forecasting, drawing inspiration from the architecture and training methodologies of foundation models in natural language processing (NLP).
+Despite being a relatively new research area, the field has evolved into a highly dynamic landscape featuring a wide variety of architectures and approaches.
+This rapid proliferation is exemplified by diverse models such as Chronos (Ansari et al., 2024a), Chronos-bolt (Ansari et al., 2024b), Chronos-2 (Ansari et al., 2025), Moirai/Moirai-MoE (Woo et al., 2024; Liu et al., 2024b), MOMENT (Goswami et al., 2024), Time-MoE (Shi et al., 2024), TimesFM (Das et al., 2024), TimesFM-2.0 (Das et al., 2024), TimesFM-2.5 (Das et al., 2024), or TiRex (Auer et al., 2025).
+In contrast to traditional time series models that are trained from scratch for specific tasks and time series, TSFMs leverage transfer learning: they are pre-trained on large collections of generic time series and can perform zero-shot forecasting without domain adaptation or fine-tuning (Liang et al., 2024).
+The reliance on massive amounts of time series training data, typically scraped from public repositories, raises the question of whether the proper evaluation of TSFMs faces similar challenges to those observed in the evaluation of Large Language Models (LLMs).
+In the NLP domain, training on vast portions of the internet has given rise to an “evaluation crisis" (Liao and Xiao, 2023), in which test set contamination (Mirzadeh et al., 2024; Ravaut et al., 2024; Li et al., 2024a) and memorization effects (Chang et al., 2024) have led to overly optimistic performance estimates.
+Whether current TSFM evaluations are vulnerable to analogous issues remains an open question, with important implications for fair and reliable model comparison.
+While benchmark strategies such as benchmarking on held-out data (Aksu et al., 2024) and clean train-test splits (Qiu et al., 2024) have been proposed recently, we doubt that these approaches effectively prevent information leakage in the context of TSFMs.
+Beyond the risk of direct information leakage, the increased memorization capacity of scaled foundation models poses additional challenges (see Section 3).
+This article examines the validity and reliability of current TSFM evaluation practices and discusses the risk of an evaluation crisis in the TSFM domain.
+We address this issue through two contributions: first, we conduct a comprehensive train-test set lineage analysis of 22 leading TSFMs and review all major evaluation benchmarks available as of early 2026.
+Second, we empirically demonstrate the phenomenon of temporal overlaps of correlated train and test series, a challenge specific to TSFMs, through a dedicated experiment.
+Our investigation yields two primary findings that suggest that TSFM evaluation is indeed compromised:
+  • First, while we find no evidence of unintended direct information leakage within the training and evaluation of individual TSFMs, our analysis reveals substantial ambiguity in data separation between models.
+In several cases, one model’s training set serves as another’s test set.
+This issue is further exacerbated by dataset remixing, renaming, and inadequate documentation, which already pose significant evaluation challenges for newer models (see Section 3.1).
+  • Second, we identify the temporal overlap of correlated train and test series as an indirect form of information leakage specific to TSFMs.
+External shocks, such as the COVID-19 pandemic to stock prices, as well as exogenous drivers like local weather conditions to local solar production, act as factors that directly shape or distort the causal dynamics between multiple time series (see Section 3.2).
+As a result, a TSFM trained and evaluated on disjoint time series covering the same time period can effectively memorize patterns induced by such causal relations (see Figure 3).
+Based on these findings, we argue that the rise of TSFMs necessitates new benchmarking methodologies to prevent an evaluation crisis akin to that observed in NLP.
+2 Background
+2.1 Time Series Foundation Models
+A foundation model is defined as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted or fine-tuned to a wide range of downstream tasks” (Bommasani et al., 2021).
+The term is primarily associated with Transformer-based Large Language Models (LLMs) like the ChatGPT or Llama families (OpenAI, 2023; Grattafiori et al., 2024).
+Their success is largely attributed to scaling laws (Kaplan et al., 2020), describing the practice of increasing training data, parameters, and compute to improve a model’s predictive performance.
+Although LLMs demonstrate impressive zero-shot capabilities across benchmarks (Liang et al., 2022), concerns persist that they partly rely on memorizing benchmark data (Chang et al., 2024; Carlini et al., 2022).
+Such test set contamination, where test samples appear in pre-training data like Common Crawl [^1] https://commoncrawl.org/ or The Pile [^2] https://pile.eleuther.ai/, remains common despite improved dataset curation, leading to inflated performance estimates (Li et al., 2024a; Singh et al., 2024; Grattafiori et al., 2024; Mirzadeh et al., 2024; Ravaut et al., 2024).
+Recognizing the structural similarity between language modeling and time series analysis, recent research has adapted the Transformer architecture to forecast future values of time series, ranging from the classical encoder-decoder Transformer (Ansari et al., 2024a) over Decoder-only models (Das et al., 2024) to Encoder-only models often applied in multivariate time series forecasting (Goswami et al., 2024).
+Beyond this, TSFMs are based on a variety of other architectural backbones such as LLMs Gruver et al.
+(2024); Xue and Salim (2024), Vision-based Autoencoders Chen et al.
+(2024), and mixed architectures that leverage LLM reprogramming Jin et al.
 (2024).
-Especially when comparing TSFMs against each other, current approaches face limitations in defining an information-leakage-free test set for all investigated models (Aksu et al., 2024).
-This challenge is even more pronounced for larger models developed based on scaling laws, as their increased capacity may amplify the memorization effect, leading to a greater impact of test set contamination (see Section 2).
-In this article, we discuss the evaluation challenges arising from current practices of training and evaluating TSFMs and propose requirements that a TSFM benchmark should fulfill.
-Empirically, our position is supported by an analysis of the lineage of the training and test sets of 15 prominent TSFMs and a further investigation of existing TSFM benchmarks [^1] TSFM benchmarks that existed in the beginning of 2025..
-In addition, we provide a theoretical derivation of the phenomenon of global pattern memorization, which is specific to TSFMs and, to the best of our knowledge, has not been formally introduced before.
-First, while we did not identify cases of information leakage within any of the underlying studies, our analysis highlights the increasing confusion with regards to a clear separation between training and test sets, one of the fundamental principles of time series forecasting and machine learning, and outlines possible adverse effects.
-Concretely, we found that one model’s training set is another model’s test set.
-This confusion, which is further exacerbated through the practice of remixing and renaming of datasets, stems from an ever-growing size of datasets, combined with non-transparent documentation about which exact data points have been used for training and testing, and complicates the fair benchmarking between TSFMs.
-We will show that this is not just a danger, but is actually already happening, using two examples where information leakage occurs in benchmarks 3.1.
-Second, we highlight a more subtle instance of information leakage, particularly in time series forecasting, namely the memorization of global patterns.
-Global external shocks like the COVID-19 pandemic or global economic crises can profoundly impact numerous aspects of our economies and societies and, hence, are reflected in many diverse time series.
-A TSFM trained and evaluated on time series that are disjunct but stem from the same period and are both impacted by such an external shock can still profit from memorizing these patterns.
-In addition to the issue of information leakage, we point out that current TSFM evaluation methods are not fully consistent with previous approaches in time series forecasting.
-In Section 3.3 and 3.4 we show that TSFMs typically do not implement best practices such as time series cross-validation or choosing representative time series for a domain.
-This article argues that in the age of TSFMs a new benchmarking methodology is required to prevent a repetition of the LLM evaluation crisis and calls for new proposals that fulfill the requirements developed in this paper.
-2 Foundation Models in Time Series
-A foundation model is “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks" (Bommasani et al., 2021).
-The term is primarily associated with Transformer-based LLMs, especially decoder-only models like the ChatGPT or Llama families of models (OpenAI, 2023; Grattafiori et al., 2024).
-The success of foundation models can be attributed to scaling laws (Kaplan et al., 2020), that is, increasing training data, model parameters, and computational resources significantly improves the model’s performance.
-LLMs have demonstrated remarkable zero-shot learning capabilities across diverse benchmarks (Liang et al., 2022).
-However, concerns have been raised that these models might just memorize publicly available questions and answers of well-known benchmarks (Chang et al., 2024).
-Memorization (Carlini et al., 2022) is closely related to test set contamination, describing the situation where exact or highly similar samples from a test set are leaked to the LLM’s pre-training data (Li et al., 2024a).
-This contamination is a side-effect of the vast size and incomprehensible nature of today’s pre-training datasets, such as Common Crawl [^2] https://commoncrawl.org/ or The Pile [^3] https://pile.eleuther.ai/ (Li et al., 2024a; Singh et al., 2024).
-Despite efforts to improve dataset curation in recent LLMs (Grattafiori et al., 2024), evidence shows that test set contamination remains prevalent in various forms inflating generalization performance (Mirzadeh et al., 2024; Ravaut et al., 2024; Li et al., 2024a).
-Like generative LLMs, time series forecasting models are trained to autoregressively predict the next value(s) given a sequential series of past data points.
-Hence, it is not surprising, that researchers started to adapt decoder-only Transformers for time series forecasting.
-Some of these models encoded numerical time series values as natural language tokens (Gruver et al., 2024; Xue & Salim, 2024), others have developed novel tokenization techniques for numerical data, such as TimeGPT-1 (Garza et al., 2024) and Chronos (Ansari et al., 2024).
-Like LLMs, decoder-only Transformers for time series forecasting can be pre-trained on generic time series and adapted or fine-tuned to domain-specific time series optionally.
-At this, they seem to leverage the same scaling laws as LLMs (Edwards et al., 2024), leading to Time Series Foundation Models able to zero or few-shot forecast time series of different frequencies and domains adapting to diverse lengths of input context and forecasting horizons (Das et al., 2024; Liang et al., 2024).
-Traditional statistical or machine-learning-based time series forecasting models have to be trained on the time series that should be forecasted.
-These local models exhibit only limited generalization capabilities, that is, they can only predict values that are independent and identically distributed (i.i.d.) to their training data.
-Combined with the common practice to use time-based train/test splits when evaluating time series models (Hyndman & Athanasopoulos, 2021), this property of local models protects them from information leakage.
-Table 1: Summary of models, citations, and release years.
-Model Name Author Year
-Time-MoE Shi et al. (2024) 2025
-Chronos Ansari et al. (2024) 2024
-Lag-LLama Rasul et al. (2024) 2024
-Moirai Woo et al. (2024) 2024
-Moirai-MoE Liu et al. (2024b) 2024
-MOMENT Goswami et al. (2024) 2024
-Timer Liu et al. (2024c) 2024
-TimesFM Das et al. (2024) 2024
-Time-LLM Jin et al. (2024) 2024
-TTM Ekambaram et al. (2024) 2024
-UniTime Liu et al. (2024a) 2024
-VisionTS Chen et al. (2024) 2024
-ForecastPFN Dooley et al. (2023) 2023
-GPT4TS Zhou et al. (2023) 2023
-LLMTime Gruver et al. (2024) 2023
-TimeGPT-1 Garza et al. (2024) 2023
-TSFMs, in contrast, are often evaluated on zero-shot or few-shot forecasting tasks using held-out datasets, similar to LLM evaluation (Ansari et al., 2024; Liu et al., 2024b; Woo et al., 2024).
-Notably, the terms “zero-shot” and analogously “few-shot” need to be distinguished between LLMs and TSFMs: while LLMs adapt to different tasks without any example in zero-shot learning (Brown et al., 2020), TSFMs usually forecast zero-shot by adapting to different input lengths, frequencies, horizons and domains (Ansari et al., 2024; Das et al., 2024; Woo et al., 2024).
-The specialized practices of pre-training and zero-shot forecasting in TSFMs introduce novel challenges while simultaneously amplifying existing difficulties in the evaluation of time series models, as discussed in the following sections.
+Across these approaches, a key commonality lies in how data are represented: LLM-based models encode numerical values directly as language tokens like LLMTime (Gruver et al., 2024) and PromptCast (Xue and Salim, 2024), whereas some Transformer-based variants rely on specialized tokenization schemes to discretize continuous time series, as exemplified by TimeGPT-1 (Garza et al., 2024) and the T5-based Chronos (Ansari et al., 2024a).
+Furthermore, Vision-based approaches represent time series as gray-scaled images, while reprogramming-based approaches mix textual and continuous time series representations Jin et al.
+(2024).
+Similar to LLMs, all presented architectures are pre-trained on large collections of generic time series and appear to follow scaling laws (Edwards et al., 2024), enabling them to produce zero-shot forecasts and handle a wide range of frequencies and forecasting horizons.
+Only one exception—MOMENT—requires task-adaptive fine-tuning for high-quality forecasts Goswami et al.
+(2024).
+Based on an encoder-only architecture this model is capable of time series regression, classification, anomaly detection, and forecasting Goswami et al.
+(2024).
+However, due to its broad application area, the model needs task-adaptive fine-tuning of the last neural network layer that is specialized for each task.
+A key distinction between the foundation models in NLP and time series analysis concerns the definition of “zero-shot” capabilities.
+While LLMs typically adapt to entirely different tasks (Brown et al., 2020), TSFMs typically perform zero-shot forecasting by generalizing across domains, time series, input lengths, frequencies, and horizons without the need for model retraining (Ansari et al., 2024a; Das et al., 2024; Woo et al., 2024).
+2.2 Benchmarking Methodologies
+Traditional statistical or machine-learning-based models for time series forecasting generally require explicit training on the target time series.
+These local models exhibit only limited generalization capabilities; that is, they can only predict values that are identically distributed to their training data.
+Consequently, the dominant logic in time series benchmarking relies on time-based train/test splits (Hyndman and Athanasopoulos, 2021).
+This approach creates a temporal separation between training and evaluation data to prevent information leakage and is employed in major time series collections such as the Monash repository (Godahewa et al., 2021) and TFB (Qiu et al., 2024).
+To rigorously assess a model’s predictive capabilities under evolving temporal dynamics, Time-Series Cross-Validation (TSCV) is widely regarded as best practice.
+Instead of using a single time-based split, TSCV uses the idea of rolling windows (see left in Figure 1).
+This evaluation strategy does not necessarily require retraining; instead, the training set can be fixed while the inference window is rolled forward, reducing evaluation costs to inference-only computation (Hewamalage et al., 2023).
+Forecasting competitions, like the M-competitions or Kaggle, try to avoid the risk of test set contamination by soliciting forecasts for unpublished private test sets (Makridakis et al., 2022; Bojer and Meldgaard, 2021).
+These competitions are typically held at irregular intervals.
+While all of the above evaluation strategies rely on historical data for testing, the recent M6 competition introduced a live evaluation using future values of financial assets on selected dates as the target (Makridakis et al., 2024).
+Figure 1: Evaluation strategies along the time (left) or the domain dimension (right).
+While current TSFM benchmarking studies also rely on historical evaluation data, their logic diverges from the above described time-series cross-validation strategy.
+Instead of using TSCV on one series as illustrated in Figure 1 (a), they typically evaluate models using many series but with only a single time-based split per series (Das et al., 2024; Gruver et al., 2024).
+Figure 1 (b) shows the focus of most TSFM evaluations.
+Here, the focus is not on temporal depth, but on generalization across heterogeneous series, especially in zero-shot scenarios where complete datasets or domains are held out of the pre-training (Shi et al., 2024; Ansari et al., 2024a).
+This approach is focused on the generalization capabilities of TSFMs and prevents cherry-picking of test sets.
+Two prominent examples of this strategy are TSFM-Bench, which comprises 21 test sets with time series from diverse domains (e.g., energy, finance), statistical characteristics (e.g., stationary, trend), and frequencies (e.g., quarter-hourly, hourly) Li et al.
+(2025) and GIFT-Eval, which comprises 23 curated test sets with time series of seven domains, ten frequencies, and multivariate inputs (Aksu et al., 2024).
+While this series cross validation focuses on evaluating the generalization capabilities of TSFMs, it is not without drawbacks.
+Roque et al.
+(2025) showed that with just four cherry-picked test sets, out of the distribution of available test sets, 46% of models could appear as state-of-the-art purely due to selection bias.
+Notably, deep learning models were found to be more susceptible to this variance than classical baselines.
+Especially when evaluation general-purpose TSFMs, generalization is a core evaluation focus.
+Figure 1 (c) illustrates the current best practice in TSFM evaluation, that is, a combined time and series cross validation (Yu et al., 2016; Salinas et al., 2020).
+While this approach has first adopters, notably Goktas et al.
+(2025b); Shchur et al.
+(2025), it is not yet commonly used in TSFM benchmarking.
+The scaling of pre-training and evaluation datasets in TSFMs gives rise to another challenge that can also be observed in LLM evaluation.
+As LLMs are pre-trained on large datasets crawled from the internet like Common Crawl [^3] https://commoncrawl.org/ or The Pile [^4] https://pile.eleuther.ai/, they have already seen most benchmark tasks during pre-training.
+Ultimately, this led to an LLM “evaluation crisis” Liao and Xiao (2023).
+Likewise, with large and obscure pre-training datasets, defining test datasets that do not overlap with models’ pre-training datasets becomes increasingly difficult in TSFMs Aksu et al.
+(2024); Goktas et al.
+(2025a); Xu et al.
+(2025).
+As shown in the GIFT-Eval ablation studies, information leakage can strongly inflate evaluation performance of TSFMs Aksu et al.
+(2024).
+Thus, investigating information leakage in TSFMs is an important research direction.
+Addressing this, current benchmarks developed strategies to prevent information leakage.
+GIFT-Eval provides an explicit pretraining dataset containing approximately 230 billion data points and a dedicated test with 177 million data points Aksu et al.
+(2024) staying with a series cross validation (see Figure 1 (a)) but no time cross validation.
+Two other benchmarks TempusBench and Fidel-TS, use test datasets from unknown or restricted sources to prevent information leakage Goktas et al.
+(2025a); Xu et al.
+(2025).
+Particularly, TempusBench leverages a new test set with 48 uni- and multi-variate time series forecasting tasks that is currently not in any TSFM pre-training corpus Goktas et al.
+(2025a).
+While they apply a time and series cross-validation strategy, they further propose to leverage real-world live data for evaluations Goktas et al.
+(2025a).
+Fidel-TS apply a similar strategy by leveraging restricted live APIs but implement a series cross validation without time cross validation Xu et al.
+(2025).
+Overall, all current approaches to mitigate information leakage and the possibly resulting inflation of performance rely on the development of new datasets that can currently not be in the pre-training corpus of TSFMs.
 3 Challenges of Evaluating Time Series Foundation Models
-3.1 Challenge: Test Set Contamination Through Multi-Purpose Use of Datasets
-In order to empirically study whether test set contamination is a problem in existing TSFM evaluations, we tried to track the lineage of the various datasets used for training and evaluating the TSFM listed in Table 1.
-Specifically, we distinguish between three types of data usages: pre-training datasets, train/test datasets, and zero-shot datasets.
-Pre-training datasets are utilized exclusively for foundational training to teach typical time series patterns (e.g., trends, seasonality) to a model.
-Here, the entire time series is usually used for training, and no evaluation is conducted on these datasets.
-Train/test datasets, also called in-domain or in-distribution datasets, fit a model to time series from a specific domain (e.g., finance, energy, transportation).
-Following the traditional time series forecasting methodology, these datasets are typically split at a particular time point into training and test sets (Hyndman & Athanasopoulos, 2021; Kapoor & Narayanan, 2023).
-The training data is included in the model’s training corpus during pre-training or subsequent task-specific adaptation/fine-tuning.
-The test data is used to evaluate the model’s performance within the given domain.
-When benchmarking multiple models in a specific domain, all models should use the same datasets with the same train-test splits in order to ensure a fair comparison.
-Zero-shot datasets (held-out datasets) are entirely excluded from the pre-training and task-specific fine-tuning process and used for inference only.
-Unlike LLM zero-shot datasets with diverse tasks, TSFM datasets often share frequencies and domains with pre-training data, primarily evaluating frequency and domain generalization.
-However, in zero-shot evaluation, the TSFMs make forecasts, typically in a rolling manner, taking a limited context of historical values as inputs and producing output for future time steps (Aksu et al., 2024).
-Ideally, all TSFMs would be benchmarked on the same zero-shot datasets.
+Based on the need for an examination of information leakage in current TSFMs, we decided to first conduct a TSFM train-test dataset lineage analysis (Section 3.1).
+We then further analyze the outcomes of potential indirect overlaps in time series between pre-training and test sets of TSFMs.
+3.1 Direct Information Leakage: Train-Test Sample Overlap
+To examine potential test set contamination in existing TSFM evaluations, we searched for peer-reviewed and accepted papers including the terms "Time Series Foundation Model" at leading venues in the machine learning area (TMLR, ICML, NeurIPS, ICLR, WWW) and their workshops, published up to January 2026.
+This search resulted in 22 published TSFMs that met our inclusion criteria (Table 1).
+For each of these papers, we analyzed their dataset lineage, categorizing the used datasets into pre-training, train/test and zero-shot.
+Table 1: Time series foundation models published in recent years
+Model Name Author Year Conference/Journal
+ForecastPFN Dooley et al. (2023) 2023 NeurIPS 23
+GPT4TS Zhou et al. (2023) 2023 NeurIPS 23
+Lag-Llama Rasul et al. (2024) 2024 NeurIPS 23 Workshop R0-FoMo
+LLMTime Gruver et al. (2024) 2024 NeurIPS 23
+Time-LLM Jin et al. (2024) 2024 ICLR 24
+Chronos Ansari et al. (2024a) 2024 TMLR
+TimesFM Das et al. (2024) 2024 ICML 24
+Moirai Woo et al. (2024) 2024 ICML 24
+MOMENT Goswami et al. (2024) 2024 ICML 24
+Timer Liu et al. (2024c) 2024 ICML 24
+UniTime Liu et al. (2024a) 2024 WWW 24
+TinyTimeMixer Ekambaram et al. (2024) 2024 NeurIPS 24
+Time-MoE Shi et al. (2024) 2025 ICLR 25
+Moirai-MoE Liu et al. (2024b) 2024 ICML 25
+VisionTS Chen et al. (2024) 2024 ICML 25
+LightGTS Wang et al. (2025a) 2025 ICML 25
+Sundial Liu et al. (2025) 2025 ICML 25
+TimesFM-IC Faw et al. (2025) 2025 ICML 25
+ROSE Wang et al. (2025b) 2025 ICML 25
+FlowState Graf et al. (2025) 2025 NeurIPS 25 Workshop BERT2S
+TiRex Auer et al. (2025) 2025 NeurIPS 25
+Toto Cohen et al. (2025) 2025 NeurIPS 25
+Pre-training datasets are used to fit a model to general time series patterns, such as, trends and seasonality.
+They are employed for initial training only and, hence, should be excluded from any evaluation activities.
+Train/test datasets, also sometime called in-domain or in-distribution datasets in the context of TSFMs, are datasets used to fine-tune a model to domain-specific patterns.
+They are split at a defined point in time into training and test sets and the model is fitted on the training data and evaluated on the test data (Hyndman and Athanasopoulos, 2021; Kapoor and Narayanan, 2023).
+Zero-shot datasets are used for evaluation of TSFMs on unseen data and should therefore be excluded from any training and fine-tuning.
+In a zero-shot setting, a model is given a short historical context of an unseen time series and it is assessed how well it can predict future values of this series.
 Figure 2: Lineage of dataset collections (left) used for training and evaluating recent Time Series Foundation Models (right).
@@
 Lines indicate cases where at least one dataset of a collection was used for pre-training, train/test, or zero-shot evaluation.
-Our lineage analysis, visualized in Figure 1, reveals that each TSFM uses a different combination of pre-training, train/test, and zero-shot datasets.
-While we did not find a TSFM that used the same set for different purposes, we still argue that this multi-purpose use of datasets is a problem.
-For example, suppose Model A uses a dataset for pre-training, and Model B uses the same dataset for evaluation.
-In that case, it becomes impossible to evaluate these models against each other on that dataset fairly.
-A notable case for the multi-purpose use of datasets is the subset Australian Electricity Demand from the Monash dataset collection (Godahewa et al., 2021), which is used for pre-training (e.g., Lag-Llama, Timer), train/test evaluation (e.g., Moirai), and zero-shot forecasting (e.g., Chronos, TimesFM).
-This multi-purpose use makes the dataset practically useless for benchmarking between models (e.g., it cannot be used to compare Chronos to Lag-Llama).
-Another interesting case is the Informer collection (Zhou et al., 2021), which is primarily used for zero-shot evaluations (indicated by green connections in Figure 1).
-However, some individual time series of the Informer collection, such as ETTh1, ETTh2, and ETTm1, have also been used for pre-training (e.g., in LagLlama) or train/test evaluation (e.g., in UniTime).
-Again, this means that LagLlama or UniTime cannot be benchmarked against any other model using the Informer collection.
-Note that it is not easy to exactly determine which datasets a model has been trained on.
-It requires careful reading of the original paper and sometimes even analyzing published source code.
-Doing this for a dozen or more models in a benchmarking study requires substantial effort.
-No dataset has been universally used across all TSFMs for train/test or zero-shot evaluation (see Table 2 and Table 3).
-Considering all investigated datasets, only 7% have never been used for pre-training or train/test fine-tuning, making them the only suitable candidates for genuine zero-shot model benchmarking.
-Renaming datasets is another issue we encountered during our analysis.
-For example, the Monash dataset "Elecdemand" is a scaled subset ($1/1000$) of the "Australian Electricity Demand" dataset.
-Normalization practices could introduce unintended information leakage if these datasets are treated as separate.
+Smaller datasets (used n<10) are shown as "Other".
+Our lineage analysis, the details are described in Supplementary Information Section D and the results are visualized in Figure 2, reveals that there is no consensus between model developers which publicly available datasets should be used for what purpose.
+In fact, each TSFM team uses a different combination of pre-training, train/test, and zero-shot datasets.
+Whereas we find no indication that any individual TSFM suffers from information leakage within its own training and evaluation, the widespread practice of re-purposing datasets introduces substantial challenges for benchmarking across TSFMs.
+Once a dataset has been used for pre-training or fine-tuning by any model, it must be regarded as contaminated and is no longer admissible for TSFM benchmarking.
+Our lineage analysis shows that this is not merely a theoretical concern, but a concrete and pervasive practical issue.
+No dataset has been universally used across all TSFMs for train/test or zero-shot evaluation (see Supplementary Information Section D).
+And only 6% of the analyzed datasets have never been used for pre-training or fine-tuning, making them the only suitable candidates for genuine zero-shot model benchmarking.
+In Supplementary Information Section B, we summarize literature showing that leakage between training and test sets can easily lead to over 50% better test scores.
+Several documented leakage cases suggest that the often unintended overlaps between train and test data are not always identified during the peer-review process Saravanan et al.
+(2024); Montet et al.
+(2025); Li et al.
+(2024b).
+This is not surprising as exactly determining on which datasets a model has been trained requires careful reading of the original paper, its appendix, and sometimes even analyzing published source code.
+Doing this for a dozen or more models when performing a benchmarking study requires substantial effort from the authors and reviewers as well.
+As more TSFMs are published, authors increasingly face the challenge of navigating complicated dataset relationships.
+For example, the authors of the Tirex model discovered that a dataset called "Solar" appears in parts in the Gift-Eval test data Aksu et al.
+(2024) as well as in the Chronos pre-training data, which is used by several models including Tirex, Toto, and Sundial Ansari et al.
+(2024a); Auer et al.
+(2025); Liu et al.
+(2025); Cohen et al.
+(2025).
+As a precaution, Tirex and Toto excluded the dataset, though Tirex explicitly notes the difficulty of verifying whether the data is actually identical.
+Through our lineage analysis, we found that in the Chronos pre-training data, the "Solar" dataset includes solar power data from all US states at two frequencies: 5T (5 minutes) and 1H (hourly), where the 5T data was sourced directly from the original data provider [^5] https://www.nlr.gov/grid/solar-power-data.
+Gift-Eval, however, sourced only the Alabama subset as a test set, aggregated to 10T, 1H, 1D (daily), and 1W (weekly), through the LSTNet dataset repository [^6] https://github.com/laiguokun/multivariate-time-series-data/tree/master/solar-energy.
+LSTNet also sourced the data from the original provider but, as noted, included only Alabama rather than all states.
+In addition, the Alabama solar subset appears in the Monash Repository [^7] https://forecastingdata.org/ at frequencies 10T and 1W.
+This modified version of the source data was used as pretraining data by the models TinyTimeMixers and Lag-Llama Ekambaram et al.
+(2024); Rasul et al.
+(2024).
+While TinyTimeMixers and Lag-Llama were published before Gift-Eval, the authors of Sundial validate against Gift-Eval and state they excluded any overlaps without explicitly listing which datasets they excluded.
+This example illustrates how dataset provenance can become convoluted through multiple transformations and repositories, making it difficult for researchers to identify overlaps.
+While in the previous example the dataset name was constantly "Solar", relying on dataset names to avoid train-test overlaps is not sufficient.
+The identification of suitable datasets is complicated by dataset remixing and renaming.
+For example, the dataset "Elecdemand" from the Monash Repository is a scaled subset ($1/1000$) of the “Australian Electricity Demand” dataset.
 Similarly, the dataset "ElectricityLoadDiagrams20112014" appears under various names, such as "Electricity" in the Autoformer and Monash collections (Godahewa et al., 2021; Wu et al., 2022) and "ECL" in the Informer (Zhou et al., 2021) collection.
-Dataset collections sometimes also contain the same variables but different observations.
-For instance, the car ride shares dataset is included in both the Monash and GluonTS collections.
-Although both repositories cite the same GitHub source [^4] https://github.com/fivethirtyeight/uber-tlc-foil-response, a close examination reveals that they represent different time series of distinct For-Hire Vehicles.
-The empirical analysis using the Moirai TSFM by Aksu et al.
-(2024) demonstrates the effects of information leakage between training and test sets on the predictive performance of TSFMs.
-The authors prepared a new pre-training and held-out evaluation dataset without any overlap and trained the Moirai architecture on it.
-A different Moirai model from a previous publication was taken which pre-training dataset contained 0.1% data of the newly defined held-out evaluation dataset leading to a deliberate small portion of information-leakage.
-They then compared the performance of the models with different numbers of parameters and different forecasting horizons.
-The models trained on the training set with information leakage achieved an average MAPE between 7.6 and 59 percentage points lower than the same architecture trained on leakage-free training data (see Appendix B.2 for detailed results).
-These results vividly demonstrate the significant impact that test set contamination has on the performance of TSFMs.
-Larger model sizes especially seem prone to this, suggesting that bigger models tend to memorize rather than generalize.
-Another example of the information leakage effect is observable in a recently proposed TSFM benchmark.
-The benchmark creators, due to the TSFMs’ incomprehensible datasets, unintentionally included three evaluation datasets as test sets that had already been used in the pretraining of TimesFM, UniTS, and TTM (Li et al., 2024b).
-Based on an analysis of the benchmark results, this leads to an advantage of 47% - 184% lower mean squared error (MSE) rate compared to best models not pre-trained on the leaking datasets.
-The advantage of the best TSFM on non-leaked datasets is only between 0.3% and 14%.
-Of course, the effect should be evaluated using comparisons on the same datasets, but the performance benefit through information leakage appears to align with findings from the Moirai leakage example.
-This highlights the difficulties for benchmarking between TSFMs.
-3.2 Challenge: Memorization of Global Patterns
-Direct overlap between observations of the training and test sets is not the only cause for information leakage.
-A more subtle case of information leakage can happen due to indirect dependencies between data points in the training and test sets (Kapoor & Narayanan, 2023).
-Typically, these indirect dependencies stem from confounding bias, that is, an external factor independently associated with observations in both datasets.
-In the context of time series forecasting, global events or crises like COVID-19 or the European energy crisis, caused by Russia’s invasion of Ukraine, can influence a large number of time series from different domains.
-A key strength of TSFMs is learning patterns and transferring a suitable pattern to an unseen time series.
-But when a TSFM is trained on data including such global events or crises, information may inadvertently carry over to forecasts on test set time series from the same period and affected by the same external factors.
-This introduces a problematic dependency between the training and test data, violating the assumption of independence that is crucial for robust evaluation.
-As a result, the model gains access to information during training that is usually inaccessible in a real-world scenario at the time of prediction, leading to inflated performance metrics.
-The time component and the actual information inside a time series plays a new role for globally trained models such as TSFMs.
-Figure 2: Global pattern memorization of TSFM A when trained on time series 1 and zero-shot forecasting time series 2.
-3.3 Challenge: Evaluation Dimensions
-The typical evaluation in time series forecasting is performed across the time dimension to assess a model’s capabilities over time (Hyndman & Athanasopoulos, 2021), particularly its performance across different seasonal patterns (Figure 3, left side).
-Using time-series cross-validation, one can verify whether a model provides reliable forecasts, for example during winter and summer.
-It should be emphasized that a rolling evaluation does not necessarily require retraining the model for each rolling interval.
-One can just as well use a fixed training dataset and then roll inference over the various test sets (Hewamalage et al., 2023), making it a suitable evaluation strategy for TSFMs.
-Figure 3: Evaluation strategies along the time or the domain dimension.
-Interestingly, current TSFMs seem to evaluate only using a single time split per series (Figure 3, right side) without performing time series cross-validation, usually to reduce resource utilization or costs (Das et al., 2024; Gruver et al., 2024).
-Results are then often reported based on a single prediction per time series (Ansari et al., 2024; Das et al., 2024; Shi et al., 2024).
-They demonstrate foundation model capabilities as the models are evaluated against multiple time series from different domains (see Section 2).
-But we have not found any TSFM for which a combined time and domain cross-validation is explicitly conducted (Figure 4), even though this approach would easily increase the number of predictions and ultimately lead to a more meaningful evaluation.
-Figure 4: Evaluation along time and domain dimension.
-3.4 Challenge: Domain Representative Time Series
-Foundation models in particular aim to be generalizable across different domains.
-Still, there is no systematic derivation of why domains such as energy, transportation, weather, or web trends in particular should be predicted.
-A particularly illustrative example is the ETT dataset (Zhou et al., 2021), which is used in many TSFMs as a zero-shot benchmark (see Section 3.1).
-However, this dataset is based on two years of measurements from just two transformers in China.
-The exact location of these transformers within China is not specified, and only very general information about their origin is provided.
-Moreover, while the dataset indeed captures some noteworthy characteristics, such as both short- and long-term seasonal patterns, it still consists of only two specific time-series use cases among an uncountable variety of possible temporal dynamics observed in real-world applications.
-It is therefore unclear to what extent results obtained on ETT can be considered generalizable, whether to transformers deployed in other regions such as Africa, America, Europe, or even to a third transformer within China.
-Accordingly, it remains questionable whether the dataset allows for any broader conclusions about the overall capability of models in a special domain (Hewamalage et al., 2023).
-4 Current Benchmarking Approaches
-One could argue that existing best practices for model training and testing are sufficient to mitigate risks such as test set contamination and global pattern memorization.
-But are these strategies still adequate in the age of TSFMs, or must they be adapted, combined, or restructured?
-Existing strategies can be broadly categorized into benchmarking strategies and research best practices.
-As the dominant logic of time series benchmarking (Hyndman & Athanasopoulos, 2021), time-based train/test splits, including those in the Monash dataset collection (Godahewa et al., 2021) and TFB (Qiu et al., 2024), prevent leakage within a single TSFM.
-However, it is still unsuitable for fair benchmarking, as teams may choose different training and testing sets (see Section 2).
-Moreover, such predefined train/test splits typically do not use a global split point across all series in a collection, which would protect against information leakage through memorizing global patterns.
-For example, the Monash collection series uses varying start and end points, preventing the possibility of a global time-based split across all included time series (see Appendix 7).
-In terms of foundation models, held-out datasets like GIFT-Eval (Aksu et al., 2024) focus on creating curated pre-training and test datasets without any overlap to prevent test set contamination.
-This approach enables a robust zero-shot forecasting evaluation, which is critical to TSFMs.
-However, the held-out dataset approach only applies to TSFMs developed in the future, as past models cannot be benchmarked reliably due to potential data contamination caused by overlap between their pre-training data and the held-out dataset.
-Furthermore, in case researchers want to include more data into pre-training than provided in a collection like GIFT-Eval, it requires careful checks for potential test set contamination in zero-shot forecasting.
-Lastly, global patterns like events or crises can still affect both the training and held-out test sets of such collection as long as no global time-based split is enforced.
-Combining held-out datasets, global time-based train/test splits and retraining and benchmarking all TSFMs on a standard pre-training and test set would prevent information leakage and enable fair benchmarking.
-Yet, retraining all model architectures used in a benchmark is obviously very resource-intensive and prone to implementation errors.
-In addition, forcing all models to use the same data for pre-training mitigates the advantages of the scaling laws (Edwards et al., 2024).
-Using anonymized real-world benchmarking test datasets (e.g., masking or scaling them) can also help to address information leakage concerns.
-Anonymizing real-world time series would help prevent the simple recall of original data points in case they have been included in a model’s pre-training set (Der et al., 2023).
-However, pre-processing techniques like scaling, feature engineering, and tokenization, which are usually implemented in a TSFM architecture, could potentially reverse such anonymization efforts, leading to the original or a very similar time series.
-Hence, the effectiveness of this approach is questionable and would require sophisticated anonymization techniques that preserve statistical properties while preventing direct recalling of data points.
-Competition-based splits, like in the M-competitions (Makridakis et al., 2022), require forecasts on unpublished time-splits of known series.
-Such competitions start by publishing the training set while the test set remains private.
-Forecasts are registered during the competition, usually organized via Kaggle [^5] https://www.kaggle.com/competitions/m5-forecasting-accuracy/overview and metrics are calculated using the private test set.
-Although these approaches prevent test set contamination by registering forecasts based on time-based splits, they do not ensure a rigorous held-out dataset split for zero-shot forecasting.
-Furthermore, the hidden test set usually contains static, past data for evaluation throughout the competition.
-Besides enabling TSFMs to benefit from global pattern memorization, organizing such competitions requires substantial resources for administration, leading to long waiting periods between competitions.
-In a more innovative way, the recent M6 competition employed a strategy of using real-world live data from 100 financial assets and registering predictions into the real future (Makridakis et al., 2024).
-The main task was not forecasting but ranking the assets according to their future values.
-Still, it was a closed competition in a predefined time frame where long waiting period persists for results as with previous competitions which can be a hurdle in the fast developments of TSFMs.
-In contrast, synthetic data for testing promises an unlimited amount of truly novel and so far unknown data.
-In fact, for pre-training, synthetic data has already been used successfully (Dooley et al., 2023).
-Yet, the external validity of solely using synthetic data for testing remains questionable, as synthetic data generation itself needs to be evaluated (Yuezhang et al., 2021).
-Research best practices such as pre-registration for predictive modeling provide indirect mechanisms to mitigate risks of test set contamination due to multi-purpose use of datasets or global pattern memorization.
-Pre-registration for predictive modeling has been proposed as a rigorous research method originating from the experimental sciences (e.g., medicine, psychology) (Hofman et al., 2023).
-This method requires an ex-ante documentation and publication of the whole predictive modeling process before model training, including all planned data collection and preparation as well as modeling and evaluation activities.
-Model cards can complement this ex-post, by providing all information on data, training, and benchmarking to other researchers in one document (Mitchell et al., 2019).
-While these best research practices have the potential to prevent cherry picking of training and test sets and evaluation metrics and provide deeper insights into the models, they cannot mitigate the risks stemming from multi-purpose use of datasets and memorization of global patterns.
-Lastly, the academic review process ensures proper data splits and detects any information leakage.
-While the peer review process has proven its ability to spot methodological issues, relying solely on this approach may not comprehensively prevent unintentional benchmark contamination (Li et al., 2024b).
-Furthermore, reviewers indeed cannot assess the validity of all models included in a benchmarking study.
-Hence, as witnessed in LLM evaluation research, academic review processes may oversee test set contaminations (see Section 2).
-In sum, while pre-defined distinctions between training, test, and zero-shot sets improve evaluation integrity, they hinder comparability with prior models and mitigate scaling laws.
-Adhering to best practices may help identify test set contamination, but cannot entirely prevent it, especially in the presence of global pattern memorization.
-Given the rapid evolution of TSFMs, waiting long periods between competitions is increasingly impractical and risks stagnating meaningful progress, while anonymization of data cannot fully mitigate TSFM benchmarking challenges.
+Dataset collections sometimes also contain different observations of the the same variables.
+For instance, the car ride shares dataset is included with same frequency and from the same year 2015 in both the Monash and GluonTS collections.
+Although both repositories cite the same GitHub source [^8] https://github.com/fivethirtyeight/uber-tlc-foil-response, a close examination reveals that they contain different time series of distinct For-Hire Vehicles (Uber in GluonTS vs Lyft in Monash).
+Consequently it is necessary to always compare the underlying data itself, keeping in mind it could have been transformed in various ways (scaling, handling missing values, renaming, resampling, etc.).
+3.2 Indirect Information Leakage: Temporal Overlap of Correlated Train and Test Series
+Direct overlap in data points between training and test sets is not the only cause for information leakage.
+In the context of time series data, a more subtle case of information leakage can happen due to indirect dependencies between data points in the training and test sets (Kapoor and Narayanan, 2023).
+Often, these indirect dependencies stem from confounding bias, that is, an exogenous factor influencing observations from both datasets.
+For example, global events or crises like COVID-19 or the European energy crisis, caused by Russia’s invasion of Ukraine, can influence a large number of time series from different domains.
+However, not only global phenomena but also local factors can introduce such dependencies.
+Weather conditions at a specific location, for instance, directly influence solar energy production data, creating a causal relationship between meteorological measurements and power generation time series.
+Similarly, local traffic patterns may affect both air quality measurements and delivery time predictions within the same geographic area.
+A key strength of TSFMs lies in their ability to learn subtle patterns in a flexible, data-driven manner and to apply these patterns to forecast future values of previously unseen time series.
+At large scale, however, TSFMs may inadvertently memorize patterns observed in one or more training time series and transfer them to series in the test set from the same time period.
+As long as training and test series are correctly ordered in time and do not overlap, this behavior does not constitute a serious issue.
+However, if training and test series overlap temporally and have a casual relationship (confounding or direct causation), such pattern transfer can result in information leakage.
+In the example illustrated in Figure 3, a highly flexible TSFM may learn the characteristic shape of the stock market decline caused by the global COVID-19 pandemic from the time series of the DAX index and reproduce it when tasked with forecasting the S&P 500 index over the same period.
+Figure 3: Indirect Information Leakage - Temporal Overlap of Correlated Train Series A and Test Series B.
+This temporal overlap of correlated train-test samples introduces a problematic dependency between the training and test data, violating the assumption of independence that is crucial for robust evaluation.
+As a result, the model gains access to information during training that would be inaccessible in a real-world scenario at the time of prediction, leading to inflated performance metrics.
+Therefore the explicit time period of a pattern and its actually encoded information gain importance for globally trained models such as TSFMs.
+This issue is not merely theoretical, but has been demonstrated empirically.
+Rodrigo and Ortiz (2024) conducted an experiment isolating the influence of temporal overlap information leakage.
+They build a setup with highly correlated time series, such as metro, bus, road, and train usage in Madrid for the year 2024.
+They trained two models in a global manner with all series.
+One model was trained with a common train-test split date for all series.
+For the other model only the metro time-series was cutted at the train-test split date, keeping the other series fully in the train set.
+After training, they only predicted the test period of the metro series, which was never part of the training dataset.
+The second model with full access to the other correlated series during training achieved a Mean Absolute Error (MAE) of 248,24 compared to 439,14 to the first model with a common train-test split date across all series.
+This corresponds to a performance improvement of approximately 43.47% attributable solely to this indirect form of leakage, confirming that access to correlated and temporally overlapping external information significantly distorts evaluation results even without direct access to the target series.
+Figure 4: Zero-shot forecasting for S&P 500 stock index during the Covid shock in the year 2020.
+In blue is the ground truth, in green a pretrained model with data up to December 2019, in red a model pretrained until December 2020.
+In both cases S&P 500 is not part of the training data.
+While the study by Rodrigo and Ortiz (2024) used time series where the complete training data is highly correlated and temporally overlapping with the complete test data, we replicate an TSFM training process in a more realistic scenario.
+We incorporate diverse data where only a small percentage of observations is affected by indirect leakage.
+To demonstrate the impact of temporal train-test leakage on model performance under these conditions, we conducted an experiment using a trained-from-scratch Time-Series Transformer (TST) (Zerveas et al., 2020).
+As we have seen in Section 3.1, any other pre-trained TSFM could be already affected by this form of information leakage.
+The encoder-only model shares multi-variate characteristics with recent models such as MOMENT or Chronos-2 (Goswami et al., 2024; Ansari et al., 2025) but is significantly smaller, enabling faster training saturation.
+We trained the model using 10 different random seeds for two different training sets [^9] Details on the general setup, used datasets, scaling, hyperparameter-tuning and results can be found in Supplementary Information Section C: one with a temporal overlap to the test set during the COVID-19 pandemic and one without such overlap.
+The first training set comprised the NN5 Daily and Tourism Monthly (Monash collection), and the American (Dow Jones), Brazilian (Bovespa), British (FTSE 100), Canadian (TSX), French (CAC 40), German (DAX) and Japanese (Nikkei 225) stock indices from 2015 to the end of 2019.
+The only difference in the second training set was that the stock index data spanned from 2016 to the end of 2020, comprising observations affected by the COVID-19 outbreak.
+No data points were shared directly between the training and test sets.
+Table 2: Per-seed S&P 500 MAE breakdown for indirect leak from other countries stock indices vs no-leak experiments.
+The percentage difference shows how much better the model with leakage data performs than the non-leakage model.
+For each seed, the better MAE in displayed in bold.
+Seed MAE Before Split MAE After Split
+No Leak Indirect Leak % Diff No Leak Indirect Leak % Diff
+410 230.53 227.72 1.22% 157.85 199.33 -26.28%
+1425 219.68 178.09 18.93% 120.77 82.53 31.67%
+1680 193.52 187.19 3.27% 80.74 81.22 -0.60%
+1825 247.76 156.12 36.99% 110.47 70.36 36.31%
+2287 231.19 206.28 10.78% 102.42 100.54 1.84%
+3658 219.23 156.91 28.43% 117.25 84.62 27.84%
+4013 224.75 193.16 14.06% 114.19 97.98 14.20%
+4507 206.45 213.70 -3.51% 93.10 84.01 9.77%
+8936 189.40 199.38 -5.27% 103.59 153.67 -48.35%
+9675 206.30 182.99 11.30% 82.35 132.44 -60.82%
+Mean 216.88 190.15 11.62% 108.27 108.67 -1.44%
+Std Dev 17.15 21.88 21.10 38.83
+We selected stock prices as our target variable as they are specifically hard to model appropriately in a long-term valid manner Goyal et al.
+(2024), suggesting that any performance gaps must come from external leaked information.
+For both models, we used the S&P 500 index as a zero-shot test set, which was not part of both datasets.
+Additionally, we chose a time-frame where a common cause had influence on many different time series and chose therefore the COVID-19 outbreak.
+In Figure 3, time series 1 shows the DAX index and time series 2 represents the S&P 500 index.
+Note how the COVID-19 outbreak is visible in both series.
+Figure 4 displays the mean and standard deviation of the predictions for both models across the ten runs, and Table 2 shows the specific values for each seed.
+For almost all seeds, the model trained on data with temporal overlap to the test set predicted the initial market crash following the outbreak significantly better than the non-leakage model.
+In total, the temporal overlap model performed better eight out of ten times.
+Up to the first of May, the leakage model achieved an average MAE of about 190 compared to about 217 for the non-leakage model.
+For the remaining part, both models had a very similar average MAE, with the non-leakage model being slightly better (108.3 compared to 108.7 for the non-leakage model.
+When looking at the specific runs, shown in Table 2, the performance difference between the two models is up to about 37% in favor of the leakage model during the shock.
+Despite the temporal overlap representing only a small fraction of the overall training data, specifically seven time series out of 484 (or 1,600 out of 205,859 observations in total) (Table 6), this performance gap underscores the significant impact the small temporal overlap of correlated series can have on the predictive performance of a model.
 4 TSFM Benchmark Requirements
 The essence of time series forecasting is to predict the future.
-Paradoxically, using data from the past to evaluate their ability to do so is standard practice as we have seen in previous Section 4.
+Paradoxically, using data from the past to evaluate their ability to do so is standard practice (Section 2.2).
 Before the advent of TSFMs, this paradox was not a serious issue, as traditional time series models have to be trained on past values of the same time series they are expected to continue.
 The only pitfall to avoid is adhering to the correct ordering of training and test observations.
-Yet, as we have outlined in Section 2, the possibility of pre-training TSFMs on vast collections of time series from different domains and using them without training on the target time series introduces new risks of information leakage, possibly resulting in test set contamination and memorization of global patterns.
-Now there are a variety of ways (Section 4) to test and compare TSFM, but in our opinion none of them fully solves the problems identified (Section 2).
-Instead of only pointing out which benchmarking methodologies are not suitable for TSFMs, we want to be proactive and develop guidelines how a TSFM benchmark should be designed.
-Based on the advantages and disadvantages of the current benchmarking approaches (Section 4) combined with our analysis (Section 2), we propose the following requirements for a time series foundation model benchmark methodology:
-  1. 1.
-    Prevention of Information Leakage. The benchmark must guarantee that no portion of test data is present within any model’s training corpus.
-  2. 2.
-    Unambiguous Data Splits and Transparency. It should employ clear and standardized protocols for dividing data into training and test sets.
-  3. 3.
-    Time- and Domain Dimension Evaluation.
-It should apply cross-validation across both time and domain dimensions, verifying model consistency across seasonal periods (e.g., summer/winter) and across diverse datasets or domains to ensure generalization and robustness.
-  4. 4.
-    Retrospective and Prospective Evaluation.
-The methodology should not be limited to models developed after the introduction of the benchmark, but must also be applicable to existing models and future models as the field evolves.
-  5. 5.
-    Flexibility in Pre-Training Data. Researchers should not face artificial restrictions on the amount, provenance, or specificity of pre-training data.
-  6. 6.
-    Mitigation of Global Pattern Memorization. Mechanisms must be in place such as global temporal cut-off points to reduce the risk of models exploiting global patterns.
-  7. 7.
-    Scalability and Sustainability. The framework should remain practical for long-term community use, avoiding excessive computational, infrastructural, or administrative requirements.
-  8. 8.
-    Real-World Representativeness. Evaluation data should reflect the diversity of time series in relevant real-world domains and frequencies.
-  9. 9.
-    Support for Zero-Shot and Fine-Tuned Evaluation. The protocol should enable assessment of models in both zero-shot and fine-tuned settings.
-  10. 10.
-    Inclusivity and Accessibility. Participation should be open to both academic and industry teams, including those using open-source or closed-source models.
-  11. 11.
-    Robustness Against Gaming and Overfitting. Safeguards must be implemented to prevent exploitation of the test data, such as rotating test series, or anonymizing series during evaluation.
-Now, one could argue that these are wishful requirements for a benchmark, but that they are unrealistic to fulfill all and compromises must be made.
-We would like to challenge precisely this and encourage the community to look for solutions that cover as wide a range of requirements as possible.
-For example, a benchmark based entirely on constantly generated synthetic data could meet more than half of the 11 requirements.
-While there has been impressive progress in mimicking real-world data (Narasimhan et al., 2024), a question arises: as the generation process becomes more similar to actual historical data, does the risk of information leakage or memorizing global patterns increase?
-On the other hand, if the generation process is more independent of past data, we once again encounter the issue of lacking real-world relevance.
-Additionally, such synthetic data may be vulnerable to gaming if someone can reverse-engineer the data generation process.
-Nevertheless, it is worth considering this approach further.
-To more comprehensively address the spectrum of requirements, a conceptual shift in how TSFM benchmarks are designed might be necessary.
-Imagine a framework built around a continuously advancing global temporal split.
-In such a system, TSFMs would be tasked with forecasting actual future events, with evaluations performed retrospectively as real-world data becomes available.
-This dynamic approach, where the “test set" is always the yet-unseen future, inherently addresses critical issues like information leakage from test set contamination or rote memorization.
-This evaluation seems especially applicable for time-series forecasting, which is uniquely concerned with future prediction unlike typical NLP or computer vision tasks.
-Furthermore, since the test data would not exist during model training, restrictions on training data could be completely lifted, facilitating the evaluation of a broader range of models, including closed-source TSFMs.
-Still, this paradigm introduces new complexities.
-For instance, distinguishing purely zero-shot forecasts from fine-tuned forecasts could be difficult and leading to nonfulfillment of requirement 9.
-Here it might require complementary ideas.
-To further promote transparency regarding training data, the adoption of model cards could be encouraged as a best practice, though community acceptance, as seen with LLMs, might present a hurdle.
-Undoubtedly, operationalizing such a benchmark would present significant implementation challenges, likely necessitating a dedicated platform.
-However, we posit that pursuing this line of thinking offers a promising avenue to resolve many of the current impasses in TSFM evaluation.
-We therefore call upon the research community to collaboratively explore and develop these concepts further, shaping a more robust and reliable future for time series forecasting benchmarks.
+However, the paradigm shift behind TSFMs introduces new possibilities for information leakage.
+Unlike local models, TSFMs learn patterns globally.
+Consequently, strict temporal ordering within a single time series is no longer sufficient to guarantee a fair evaluation.
+To rigorously assess these models, we must address leakage across two distinct dimensions: the sample dimension and the temporal dimension (see Figure 5).
+Figure 5: Information leakage risks and challenges arising from the advent of TSFMs, and the requirements needed to address them.
+Flexibility in Pre-Training Data (Training Dimension) The current lack of transparency surrounding TSFM pre-training corpora creates a substantial risk of information leakage as shown in Section 3.1.
+This growing confusion is evident in newer benchmarks and models that report results contaminated by data leakage—whether unintentional or intentional Aksu et al.
+(2024); Auer et al.
+(2025); Li et al.
+(2024b).
+Attempting to mitigate this risk by restricting pre-training data is both counterproductive and ineffective: such restrictions artificially limit the potential of foundation models and, given the opacity of large-scale pre-training corpora, our lineage analysis shows that it is almost impossible to rule out train-test sample overlaps.
+We therefore argue for full flexibility in pre-training data.
+A robust evaluation framework should be agnostic to pre-training data choices, allowing the use of any existing data to fully exploit the predictive capabilities of TSFMs.
+This perspective shifts the responsibility for preventing leakage from the TSFM developer to the benchmark designer, thereby necessitating evaluation datasets that are genuinely novel and capable of assessing generalization rather than memorization.
+Strict Global Post-Training Test Periods (Inference Dimension) Indirect temporal dependencies between time series, originating from direct or indirect causal relationships, create another more subtle risk for information leakage.
+Preventing this risk by enforcing global cut-off dates for pre-training data to separate training from test data is not a viable solution to this challenge.
+First, it would require all researchers to tailor their training protocol to exactly this date and second, it prevents new time series patterns of current time periods after the cut-off from being trained into the models.
+We therefore argue for strict global post-training test periods.
+These test periods need to be flexibly determined based on the global last pre-training time point of all evaluated models, instead of forcing a pre-training cut-off.
+For a robust evaluation framework a strict temporal barrier must be enforced across the entire universe of data: if the last pre-training timestamp across all evaluated models is $t$, only observations from $t+1$ onwards can be present in the test set of a benchmark, ensuring that TSFMs rely solely on past dynamics.
+Addressing these two dimensions (unrestricted training data and strict temporal inference separation) is non-trivial and often overlooked in current evaluation standards.
+We ask the community to develop ideas for new benchmark methodologies which are grounded in these two requirements.
+Possible directions could be constantly newly sourced test data, synthetic test data, or fast-paced competitions on unreleased and unbiased test data, just to name a few.
 5 Conclusion
@@
 We highlight key evaluation challenges that arise from the emergence of TSFMs.
-By investigating the data lineage of 15 prominent TSFMs and reflecting the common TSFM training and evaluation strategy, we identified two significant sources of information leakage: test set contamination through multi-purpose use of datasets and memorization of global patterns induced by external events or crises.
-Furthermore, we question the current evaluation strategy that considers only the domain dimension while disregarding the temporal dimension, as well as the selection of time series used to represent a domain.
-Current benchmark approaches from time-series or machine learning do not mitigate all information leakage risks.
-Based on our investigations and existing methodologies, we propose requirements for a benchmark in the age of TSFMs.
-We would like to encourage the community to develop new ideas which can fulfill the requirements like benchmarking on real future data.
-Besides the development of a benchmark methodology, we see several directions for future research that could also reinterpret our current results.
-The effect size of test set contamination (when a test dataset was present in the pretraining corpus) has only been proven to be significant in isolated cases so far.
-The magnitude of global pattern memorization is also largely unknown and is likely to vary depending on the domain, type of global event, and other factors.
-Carefully designed training regimes for TSFMs will be required to properly quantify these effects.
-Ultimately, we believe that the criteria we have established lay the foundation for robust, fair, and information-leakage-free benchmarking.
+By investigating the data lineage of 22 prominent TSFMs and reflecting on common TSFM training and evaluation strategies, we identified two significant sources of information leakage: direct information leakage by train-test sample overlaps through multi-purpose use of datasets and indirect information leakage by temporal overlap of correlated train and test series.
+Current benchmark approaches from time series forecasting or machine learning do not mitigate all information leakage risks.
+Based on our investigations and existing methodologies, we propose two key requirements for information leakage free evaluations in the age of TSFMs: flexibility in pre-training data and strict global post-training test periods.
+Beyond the development of new benchmark methodologies, we identify several directions for future research .
+To date, the effect size of test-set contamination—when a test dataset is present in the pre-training corpus—has only been shown to be significant in isolated cases.
+Likewise, the magnitude of temporal leakage arising from correlated and overlapping time series remains largely unknown in real TSFM settings and is likely to vary across domains, types of global events, and other contextual factors.
+Carefully designed TSFM training regimes will be required to rigorously quantify these effects.
+Ultimately, we believe that the requirements we have established lay the foundation for robust, fair, and information-leakage-free benchmarking.
 This is essential for preventing an evaluation crisis for TSFMs akin to what has been observed with LLMs.
+6 Data Availability
+The resulting table of the TSFM data-lineage analysis is supplied in the Supplementary Information Section D.
+For the global pattern experiment in Section 3.2, the used data is available through the source code via zenodo: https://doi.org/10.5281/zenodo.18635478
+7 Code Availability
+The code to reproduce the experiment in Section 3.2 is provided via zenodo: https://doi.org/10.5281/zenodo.18635478
 Appendix A Time Series Foundation Model Analysis
-A.1 Overlap in test data usage
-Table 2: Comparison of train/test evaluations of selected different models ordered by the total number of overlaps. For full results see Section C.
+A.1 Multi-Purpose Use of Train and Test Data
+This section showcases the varying training and testing dataset strategies of selected TSFMs.
+Table 3 reveals that there is virtually no standardized approach among TSFMs, with each model employing different combinations of datasets for train/test (or in-domain) datasets.
+While some datasets like Electricity (H) and Traffic Hourly (H) appear across multiple models, the coverage remains fragmented.
+Table 4 further illustrates this inconsistency in zero-shot evaluation practices.
+Even among datasets with five or more overlaps, the specific models testing on each dataset vary considerably.
+For instance, ETT2 (15T) is evaluated by Chronos, Moirai, and Lag-Llama, while Electricity (H) appears in Moirai, Moirai-MoE, and LLMTime.
+Another notable case for the multi-purpose use of datasets is the subset Australian Electricity Demand from the Monash dataset collection (Godahewa et al., 2021), which is used for pre-training (e.g., Lag-Llama, Timer), train/test evaluation (e.g., Moirai), and zero-shot forecasting (e.g., Chronos, TimesFM).
+This multi-purpose use makes the dataset practically useless for benchmarking between models (e.g., it cannot be used to compare Chronos to Lag-Llama).
+To understand the problem of overlaps better, we provide Figure 6.
+Also an interesting case is the Informer collection (Zhou et al., 2021), which is primarily used for zero-shot evaluations.
+However, some individual time series of the Informer collection, such as ETTh1, ETTh2, and ETTm1, have also been used for pre-training (e.g., in LagLlama) or train/test evaluation (e.g., in UniTime).
+This means that LagLlama or UniTime cannot be benchmarked against any other model using the Informer collection.
+Taken together with investigations of Roque et al.
+(2025) that the selection of evaluation datasets can highly influence the final evaluation results, this strengthens the need for a more uniform benchmark approach.
+Figure 6: Test set contaminations when comparing TSFM A with TSFM B with overlapping pre-training and test sets on the test set of TSFM A.
+Table 3: Comparison of train/test evaluations of selected different models ordered by the total number of overlaps. For full results see Section D.
 Dataset Chronos Moirai MOMENT Time-LMM
@@
 Models which have not included any of the shown datasets into their train/test evaluation are not displayed.
-For full results see Section C.
+For full results see Section D.
 Dataset Chronos Moirai Moirai-MoE Lag-Llama LLMTime
@@
 Traffic Hourly (H) ZS - - - ZS
-A.2 Data Analysis
-Additional visualizations based on the analysis of Time Series Foundation Models training and evaluation data.
-Figure 5: Most common frequencies in investigated datasets.
 Appendix B Leakage Investigations
-B.1 Potential global pattern leakage in dataset collections
-Figure 6: Test set contaminations when comparing TSFM A with TSFM B with overlapping pre-training and test sets on the test set of TSFM A.
-It may be theoretically possible that existing model evaluation results have already been influenced by global pattern leakage.
-We analyse the overlapping time frames in the Monash Dataset  [^6] https://huggingface.co/datasets/Monash-University/monash_tsf as can be seen in Figure 7.
-Test datasets shouldn’t lie in the same time period as other training datasets.
-As an example, the proposed Australian Electricity Demand test time frame (April 2015) lays in the same period as the training period of the "Weather (Australia)" training set (1900-2021).
-The Moirai and Moirai-MoE models use both datasets for train/test forecasts, meaning if their are any global patterns in the weather data which could influence the Australian Electricity Demand data, this would be a kind of information leakage.
-Similarly, the Oikolab Weather dataset ranges from 2010 January to 2021 May, which is weather data located in Melbourne.
-The test set of the dataset Pedestrian Count (also Melbourne) is located at the end of April 2020.
-This test period lays within the training data of the Oikolab Weather, which could be a greater potential global pattern information leakage than in previous example.
-Both datasets are included as well in the train/test datasets of Moirai and Moirai-MoE.
-Moreover, LagLlama uses the Beijing Multi-Site Air Quality dataset from 12 districts in Beijing from 2013 to 2017 as a pre-training dataset and simultaneously uses the Beijing PM2.5 air quality dataset from the Beijing airport between 2010 and 2013.
-Figure 7: Date range of datasets within Monash dataset collection.
-Note that the “cif_2016" dataset was originally delivered without a timestamp and the start year was defined as 1900.
-No data cleaning was done.
-B.2 Leakage in Moirai Model
+B.1 Effects of Overlaps in Train-Test Samples
+This section collects different examples where the train-test sample overlaps unintentionally led to information leakage in recent benchmark papers.
+In one case, the benchmark creators, due to the TSFMs’ incomprehensible datasets, unintentionally included three evaluation datasets as test sets that had already been used in the pretraining of TimesFM, UniTS, and TTM (Li et al., 2024b).
+Based on our analysis of the benchmark results, this leads to an advantage of 47% - 184% lower mean squared error (MSE) rate compared to best models not pre-trained on the leaking datasets.
+The advantage of the best TSFM on non-leaked datasets is only between 0.3% and 14%.
+The performance benefit through information leakage appears to align with findings from the Moirai leakage example (see Section B.2).
+In another example a peer-reviewed paper benchmarked TSFM models on the Electricity dataset which is included in every tested TSFM pretraining data Saravanan et al.
+(2024).
+Also in the energy domain, a paper benchmarked TSFMs on the Spanish dataset (among others), which is also included in the pretraining data of the tested TSFMs.
+Chronos achieved a sMAPE of 4.854 where the best non-TSFM model TiDE achieved a sMAPE of 8.102 Montet et al.
+(2025).
+This highlights the increasing difficulties for benchmarking TSFMs.
+B.2 Intended Information-Leakage Experiment for Moirai Model
 Table 5: Information leakage in TSFM comparing Moirai with (MoiLeak) and without (Moi) information leakage during training phase for different model sizes.
@@
 L 0.78 0.82 0.71 0.60 2.14 2.24
-The investigated data includes 9 datasets on short horizons, four datasets each on medium (10x short horizon) and long horizons (15x short horizon).
-We filtered the results to 4 datasets where data is available for all horizons to reasonably compare the metrics and model sizes, as shown in Figure 8.
-On short horizons, both models perform better with increasing model size while Moirai Leakage has on average 7.6% points lower MAPE score.
+The empirical analysis using the Moirai TSFM by Aksu et al.
+(2024) demonstrates the effects of information leakage between training and test sets on the predictive performance of TSFMs.
+The authors prepared a new pre-training and held-out evaluation dataset without any overlap and trained the Moirai architecture on it.
+A different Moirai model from a previous publication was taken which pre-training dataset contained 0.1% data of the newly defined held-out evaluation dataset leading to a deliberate small portion of information-leakage.
+The investigated data includes nine datasets tested on three different forecast horizons (short, medium, long).
+Not for all horizons data was available, so we filtered the results on 4 datasets where the results are available for all horizons to reasonably compare the metrics and model sizes, as shown in Figure 7.
+On short horizons, both models, leakage-free and leakage model, perform better with increasing model size while Moirai Leakage has on average 7.6% points lower MAPE score.
 At medium horizons, the leakage-free MoiraiL generally underperforms compared to other model sizes, exhibiting an average performance deficit of 32% points compared to the Moirai LeakageL model.
@@
 However, Moirai Leakage achieves an average MAPE score that is 29% points lower than the leakage-free Moirai.
-Across all horizons, the biggest difference is between MoiraiL and Moirai LeakageL, similar to typical LLM memorization behavior, when large models better memorize pre-train data than smaller models.
+Across all horizons, the biggest difference is between MoiraiL and Moirai LeakageL.
+These results vividly demonstrate the significant impact that train-test sample overlap has on the performance of TSFMs. Larger model sizes especially seem prone to this.
 Figure 7: Comparing the model Moirai once with information leakage (Moirai Leakage), where the model was exposed to the test data during pre-training and the same model trained without information leakage (Moirai).
@@
 (2024).
-Appendix C TSFM Data Lineage
-Table 5: Analysis of datasets used by investigated TSFMs with Pre-training (P), train-test splits (T/T) and test (T).
-Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
-Jena Weather Autoformer (Wu et al., 2021) Nature 10T ZS ZS T/T T/T ZS ZS ZS ZS T/T
-Jena Weather Autoformer (Wu et al., 2021) Nature H
-Jena Weather Autoformer (Wu et al., 2021) Nature D ZS
-BizITObs - Application AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S ZS
-BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps 5T ZS
-BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps H
-BizITObs - Service AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S
-BDG-2 Bear BuildingsBench (Emami et al., 2023) Energy H P P P
-BDG-2 Fox BuildingsBench (Emami et al., 2023) Energy H P P P
-BDG-2 Panther BuildingsBench (Emami et al., 2023) Energy H P P P
-BDG-2 Rat BuildingsBench (Emami et al., 2023) Energy H P P P
-Borealis BuildingsBench (Emami et al., 2023) Energy H P P P
-Buildings900K BuildingsBench (Emami et al., 2023) Energy H P P P
-IDEAL BuildingsBench (Emami et al., 2023) Energy H P P P
-Low Carbon London BuildingsBench (Emami et al., 2023) Energy H P P
-Sceaux BuildingsBench (Emami et al., 2023) Energy H P P P
-SMART BuildingsBench (Emami et al., 2023) Energy H P P P
-CMIP6 ClimateLearn (Nguyen et al., 2023) Climate 6H P P P
-ERA5 ClimateLearn (Nguyen et al., 2023) Climate H P P P P
-Alibaba Cluster Trace 2018 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P
-Azure vM Traces 2017 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P
-Borg Cluster Data 2011 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P
+B.3 Potential Overlaps of Correlated Train and Test Series in Dataset Collections
+It may be theoretically possible that existing model evaluation results have already been influenced by overlaps of correlated train and test series.
+We analyse the overlapping time frames in the Monash Dataset  [^10] https://huggingface.co/datasets/Monash-University/monash_tsf as can be seen in Figure 8.
+Test datasets shouldn’t lie in the same time period as other training datasets.
+As an example, the proposed Australian Electricity Demand test time frame (April 2015) lays in the same period as the training period of the "Weather (Australia)" training set (1900-2021).
+The Moirai and Moirai-MoE models use both datasets for train/test forecasts, meaning if their are any global patterns in the weather data which could influence the Australian Electricity Demand data, this would be a kind of information leakage, at least on theoretical basis.
+Similarly, the Oikolab Weather dataset ranges from 2010 January to 2021 May, which is weather data located in Melbourne.
+The test set of the dataset Pedestrian Count (also Melbourne) is located at the end of April 2020.
+This test period lays within the training data of the Oikolab Weather, which could be a greater potential global pattern information leakage than in previous example.
+Both datasets are included in the datasets of Moirai, Time-Moe, Sundial and Toto.
+Moreover, LagLlama uses the Beijing Multi-Site Air Quality dataset from 12 districts in Beijing from 2013 to 2017 as a pre-training dataset and simultaneously uses the Beijing PM2.5 air quality dataset from the Beijing airport between 2010 and 2013, which could also in theory lead to leakage.
+Figure 8: Date range of datasets within Monash dataset collection.
+Note that the “cif_2016" dataset was originally delivered without a timestamp and the start year was defined as 1900.
+No data cleaning was done.
+Appendix C Temporal Overlap of Correlated Series Experiment
+This section provides implementation details of the experiment conducted to show the influence of the temporal overlap of correlated train and test series in a TSFM setting.
+C.1 Data visualizations and statistics
+Figure 9 shows two example time series from major stock indices, the DAX and S&P 500.
+Despite representing different markets, both series display remarkably similar patterns during key economic phases, including crisis periods and growth cycles.
+This visual comparison highlights how global financial markets often move in tandem, reflecting shared responses to worldwide economic events.
+Figure 10 illustrates the relationship between training data and test data that overlap temporally.
+It is important to note that while the training and test periods overlap in time, they do not share the same data points, the overlap is purely temporal, meaning both datasets cover in a small part the same time period but contain different observations or series as can be seen in Table 6.
+Figure 9: Comparison of DAX and S&P 500 points over time
+Figure 10: Comparison of time periods per dataset
+Category No Leakage Indirect Leakage
+Train Test Train Test
+# Time Series 484 1 484 1
+# Data Points 205,854 230 205,859 230
+# Indirect Leakage Data Points - - 1,600 -
+Table 6: Summary of time series and data points with and without leakage
+C.2 Scaling and Model Tuning
+Min-max scaling was applied separately to each dataset to account for differences in scale across datasets and ensure optimal model performance.
+Each scaler was fitted on the respective training data and subsequently used to transform the remaining data.
+For the S&P 500 dataset, the scaler was fitted explicitly on data from 2015 to the end of 2019 and then applied to the data from 2020.
+The Time-Series Transformer was implemented via the tsai library [^11] https://timeseriesai.github.io/tsai.
+The tuning was performed via Optuna [^12] https://github.com/optuna/optuna for 50 trials to minimize the MAE on the validation set.
+The training set consisted of the first 80% of NN5 Daily and Tourism Monthly datasets as well as January 1st, 2015 until June 30th, 2019 for no-leak and January 1st 2016 until June 30th, 2020 of the stock index data for leak.
+The validation set consisted of the remaining 20% of NN5 Daily and Tourism Monthly datasets, and the stock index data from July 1st, 2019 until December 31st, 2019 for no-leak and July 1st, 2020 until December 31st 2020 for leak.
+Parameter Lower Upper Stepsize Choice
+Learning rate log(1e-05) log(1e-01) - -
+Batch size - - - 16, 32, 64, 128
+Epochs 5 30 5 -
+Dropout 0.0 0.7 - -
+FC Dropout 0.0 0.7 - -
+Number of layers 2 8 1 -
+Model dimension - - - 64, 128, 256, 512, 1024
+Number of heads - - - 4, 8, 16, 32
+Dimension of feedforward network model - - - 32, 64, 128, 256, 512
+Table 7: Optuna search space for tuning
+Parameter No-leak Leak
+Learning rate 1.018e-03 1.431e-03
+Batch size 128 128
+Epochs 30 30
+Dropout 0.5 0.3
+FC Dropout 0.27 0.004
+Number of layers 2 4
+Model dimension 128 128
+Number of heads 32 4
+Dimension of feedforward network model 32 512
+Table 8: Final model parameters
+Appendix D TSFM Data Lineage
+Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer VisionTS TTM LLMTime Time-LMM UniTime Time-MoE Chronos-2 Flowstate(under review) LightGTS Sundial TabPFN-TS Time-MoE2 TinyTimeMixer TiRex Toto YingLong (under review) Gift-Eval TOTO2 Flowstate(T) TimesFM-ICL ROSE
+Jena Weather Autoformer (Wu et al., 2021) Nature 10T ZS ZS T/T T/T ZS ZS ZS ZS T/T ZS T/T ZS T/T ZS T/T
+Jena Weather Autoformer (Wu et al., 2021) Nature H ZS T/T ZS
+Jena Weather Autoformer (Wu et al., 2021) Nature D ZS ZS T/T ZS
+BizITObs - Application AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S ZS ZS T/T ZS
+BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps 5T ZS ZS T/T ZS
+BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps H ZS T/T ZS
+BizITObs - Service AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S ZS T/T ZS
+BDG-2 Bear BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
+BDG-2 Fox BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
+BDG-2 Panther BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
+BDG-2 Rat BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
+Borealis BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
+Buildings900K BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
+IDEAL BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
+Low Carbon London BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P
+Sceaux BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
+SMART BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
+CMIP6 ClimateLearn (Nguyen et al., 2023) Climate 6H P P P P P P P P
+ERA5 ClimateLearn (Nguyen et al., 2023) Climate H P P P P P P P P P
+Alibaba Cluster Trace 2018 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P P P P P
+Azure vM Traces 2017 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P P P P P P P P
+Borg Cluster Data 2011 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P P P P P P P P
 Air Passengers Darts Dataset (Herzen et al., 2021) Econ/Fin M ZS ZS
@@
 Sunspots Darts Dataset (Herzen et al., 2021) Nature M ZS
-Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T
-Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H
-Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T
-Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H
+Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T ZS T/T ZS
+Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H ZS T/T ZS
+Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T ZS T/T ZS
+Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H ZS T/T ZS
 FRED FRED (Oreshkin et al., 2021) Econ/Fin D P
@@
 FRED FRED (Oreshkin et al., 2021) Econ/Fin Q P
-M5 GluonTS (Alexandrov et al., 2020a) Sales D ZS P P P
-Taxi GluonTS (Alexandrov et al., 2020a) Transport 30T T/T P P P
-Taxi GluonTS (Alexandrov et al., 2020a) Transport H P P
-Uber TLC Daily GluonTS (Alexandrov et al., 2020a) Transport D T/T P P P
-Uber TLC Hourly GluonTS (Alexandrov et al., 2020a) Transport H T/T P P P P
-Wiki-Rolling GluonTS (Alexandrov et al., 2020a) Web D P
-Trends hourly Google Trends Web H P
-Trends weekly Google Trends Web W P
-Trends monthly Google Trends Web D P
+M5 GluonTS (Alexandrov et al., 2020a) Sales D ZS P P P P P P P ZS ZS
+Taxi GluonTS (Alexandrov et al., 2020a) Transport 30T T/T P P P P P T/T P P P P
+Taxi GluonTS (Alexandrov et al., 2020a) Transport H P P possible P P P P P P
+Uber TLC Daily GluonTS (Alexandrov et al., 2020a) Transport D T/T P P P P P T/T P P P
+Uber TLC Hourly GluonTS (Alexandrov et al., 2020a) Transport H T/T P P P P P P T/T P P P
+Wiki-Rolling GluonTS (Alexandrov et al., 2020a) Web D P P P P
+Trends hourly Google Trends Web H P P
+Trends weekly Google Trends Web W P P
+Trends monthly Google Trends Web D P P
 Requests Huawei Cloud (Joosen et al., 2023) CloudOps T ZS
@@
 Instances Huawei Cloud (Joosen et al., 2023) CloudOps T P
-ETT1 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T
-ETT1 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T
-ETT1 Informer (Zhou et al., 2020) Energy D ZS ZS
-ETT1 Informer (Zhou et al., 2020) Energy W-THU
-ETT2 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T ZS T/T ZS ZS ZS T/T T/T T/T
-ETT2 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T
-ETT2 Informer (Zhou et al., 2020) Energy D ZS ZS
-ETT2 Informer (Zhou et al., 2020) Energy W-THU
-Weather (US) Informer (Zhou et al., 2020) Weather H ZS
-Brazilian Cities Temperature Kaggle Competition Weather M P
+ETT1 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS ZS T/T
+ETT1 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS ZS T/T
+ETT1 Informer (Zhou et al., 2020) Energy D ZS ZS ZS T/T ZS
+ETT1 Informer (Zhou et al., 2020) Energy W-THU ZS T/T ZS
+ETT2 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T ZS T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS T/T
+ETT2 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS T/T
+ETT2 Informer (Zhou et al., 2020) Energy D ZS ZS ZS T/T ZS
+ETT2 Informer (Zhou et al., 2020) Energy W-THU ZS T/T ZS
+Weather (US) Informer (Zhou et al., 2020) Weather H ZS T/T
+Brazilian Cities Temperature Kaggle Competition Weather M P possible P P P P P
 Walmart Store Sales Kaggle Competition Sales W ZS ZS ZS
@@
 TSMC Stock Kaggle Competition Econ/Fin D ZS
-Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
-LargeST LargeST (Liu et al., 2023a) Transport 5T P P P
-Beijing Subway LibCity (Wang et al., 2023a) Transport 30T P P P
-HZMetro LibCity (Wang et al., 2023a) Transport 15T P P P
-Loop Seattle LibCity (Wang et al., 2023a) Transport 5T P P P
-Loop Seattle LibCity (Wang et al., 2023a) Transport H
-Loop Seattle LibCity (Wang et al., 2023a) Transport D
-Los-Loop LibCity (Wang et al., 2023a) Transport 5T P P P P
-M_DENSE LibCity (Wang et al., 2023a) Transport H
-M_DENSE LibCity (Wang et al., 2023a) Transport D ZS
-PEMS Bay LibCity (Wang et al., 2023a) Transport 5T P P P P
-PEMS03 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P
-PEMS04 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P
-PEMS07 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P
-PEMS08 LibCity (Wang et al., 2023a) Transport 5T P P T/T ZS P P
-Q-Traffic LibCity (Wang et al., 2023a) Transport 15T P P P P P
-SHMetro LibCity (Wang et al., 2023a) Transport 15T P P P P
-SZ-Taxi LibCity (Wang et al., 2023a) Transport 15T P P P
-SZ-Taxi LibCity (Wang et al., 2023a) Transport H P
-Rotterdam LibCity (Wang et al., 2023a) Transport 2T P P
-Beijing Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P P
-CDC Fluview ILINet LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T ZS T/T T/T P
-CDC Fluview WHO NREVSS LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T P
-China Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P
-Favorita Sales LOTSA_Others (Woo et al., 2024) Sales D P P P P
-Favorita Transactions LOTSA_Others (Woo et al., 2024) Sales D P P P
-GoDaddy LOTSA_Others (Woo et al., 2024) Econ/Fin M P P
-KDD Cup 2022 LOTSA_Others (Woo et al., 2024) Energy 10T P P P
-Project Tycho LOTSA_Others (Woo et al., 2024) Healthcare W P P P
-Residential Load Power LOTSA_Others (Woo et al., 2024) Energy T P P P
-Residential PV Power LOTSA_Others (Woo et al., 2024) Energy T P P P
+LargeST LargeST (Liu et al., 2023a) Transport 5T P P P P P P P P P P
+Beijing Subway LibCity (Wang et al., 2023a) Transport 30T P P P P P P P P P
+HZMetro LibCity (Wang et al., 2023a) Transport 15T P P P P P P P P P
+Loop Seattle LibCity (Wang et al., 2023a) Transport 5T P P P ZS P T/T ZS
+Loop Seattle LibCity (Wang et al., 2023a) Transport H ZS T/T ZS
+Loop Seattle LibCity (Wang et al., 2023a) Transport D ZS T/T ZS
+Los-Loop LibCity (Wang et al., 2023a) Transport 5T P P P P P P P T/T ZS P
+M_DENSE LibCity (Wang et al., 2023a) Transport H ZS T/T ZS
+M_DENSE LibCity (Wang et al., 2023a) Transport D ZS ZS T/T ZS
+PEMS Bay LibCity (Wang et al., 2023a) Transport 5T P P P P P P P P P P P P
+PEMS03 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P P P P P P P P P
+PEMS04 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P P P P P P P P P
+PEMS07 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P P P P P P P P P
+PEMS08 LibCity (Wang et al., 2023a) Transport 5T P P T/T ZS P P P P P P P P P P
+Q-Traffic LibCity (Wang et al., 2023a) Transport 15T P P P P P P P P P P P P
+SHMetro LibCity (Wang et al., 2023a) Transport 15T P P P P P P P P P P P
+SZ-Taxi LibCity (Wang et al., 2023a) Transport 15T P P P ZS T/T ZS P
+SZ-Taxi LibCity (Wang et al., 2023a) Transport H P ZS T/T ZS
+Rotterdam LibCity (Wang et al., 2023a) Transport 2T P P P T/T ZS
+Beijing Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P P P P P P P P P P P
+CDC Fluview ILINet LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T ZS T/T T/T P P P P P P P P
+CDC Fluview WHO NREVSS LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T P P P P P P P P
+China Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P P P P P P P P
+Favorita Sales LOTSA_Others (Woo et al., 2024) Sales D P P P P P P P P P P P
+Favorita Transactions LOTSA_Others (Woo et al., 2024) Sales D P P P P P P P
+GoDaddy LOTSA_Others (Woo et al., 2024) Econ/Fin M P P P P P P P P P
+KDD Cup 2022 LOTSA_Others (Woo et al., 2024) Energy 10T P P P P P P P P P P
+Project Tycho LOTSA_Others (Woo et al., 2024) Healthcare W P P P P P P P
+Residential Load Power LOTSA_Others (Woo et al., 2024) Energy T P P P P P P P P P P
+Residential PV Power LOTSA_Others (Woo et al., 2024) Energy T P P P P P P P P P P
 Restaurant LOTSA_Others (Woo et al., 2024) Sales D P
-Solar LSTNet (Lai et al., 2017) Energy 10T
-Solar LSTNet (Lai et al., 2017) Energy 5T P
-Solar LSTNet (Lai et al., 2017) Energy H P ZS ZS ZS
-Solar LSTNet (Lai et al., 2017) Energy D
-Solar LSTNet (Lai et al., 2017) Energy W-FRI
-Exchange Rate LSTNet (Lai et al., 2017) Econ/Fin D ZS ZS T/T ZS ZS T/T P
-Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales D P
-Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales W-WED
-Australian Electricity Demand Monash (Godahewa et al., 2021) Energy 30T ZS ZS T/T T/T P P P ZS P ZS P
-Bitcoin Monash (Godahewa et al., 2021) Econ/Fin D ZS T/T T/T P P ZS P ZS P
-Car Parts Monash (Godahewa et al., 2021) Sales M T/T T/T P ZS P
-CIF 2016 Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P
-COVID Deaths Monash (Godahewa et al., 2021) Healthcare D ZS ZS T/T T/T P ZS ZS P
-Covid Mobility Monash (Godahewa et al., 2021) Transport D T/T T/T P
-Elecdemand Monash (Godahewa et al., 2021) Energy 30T T/T T/T P
-Extended Web Traffic Monash (Godahewa et al., 2021) Web D T/T T/T P P P
-FRED MD Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P
-Hospital Monash (Godahewa et al., 2021) Healthcare M ZS ZS T/T T/T P ZS ZS P
-Kaggle Web Traffic Weekly Monash (Godahewa et al., 2021) Web W P P P P
-KDD Cup 2018 Monash (Godahewa et al., 2021) Nature H T/T T/T T/T P P P ZS P P
-KDD Cup 2018 Monash (Godahewa et al., 2021) Nature D
-London Smart Meters Monash (Godahewa et al., 2021) Energy 30T T/T P P P P P P P
-M1 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T P ZS P
-M1 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P P P
-M1 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P P P
-M3 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T ZS T/T ZS P
-M3 Other Monash (Godahewa et al., 2021) Econ/Fin Q T/T T/T T/T ZS
-M3 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P ZS T/T P
-M3 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P ZS T/T P
-M4 Daily Monash (Godahewa et al., 2021) Econ/Fin D T/T P T/T T/T P T/T ZS P
-M4 Hourly Monash (Godahewa et al., 2021) Econ/Fin H T/T P T/T T/T P T/T ZS P
-M4 Monthly Monash (Godahewa et al., 2021) Econ/Fin M T/T P T/T T/T ZS T/T ZS P
-M4 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q-DEC ZS P P P ZS T/T P
-M4 Weekly Monash (Godahewa et al., 2021) Econ/Fin W-SUN T/T P T/T T/T T/T ZS P
-M4 Yearly Monash (Godahewa et al., 2021) Econ/Fin A-DEC ZS P P P ZS T/T P
-NN5 Daily Monash (Godahewa et al., 2021) Econ/Fin D ZS ZS T/T T/T P ZS P ZS P
-NN5 Weekly Monash (Godahewa et al., 2021) Econ/Fin W ZS ZS T/T T/T P ZS ZS
-Oikolab Weather Monash (Godahewa et al., 2021) Climate H T/T T/T P
-Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
-Pedestrian Counts Monash (Godahewa et al., 2021) Transport H T/T ZS T/T T/T P ZS P ZS ZS P
-Rideshare Monash (Godahewa et al., 2021) Transport H T/T T/T T/T P ZS P
-Saugeen Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P ZS P ZS P
-Saugeen Monash (Godahewa et al., 2021) Nature W-THU
-Saugeen Monash (Godahewa et al., 2021) Nature M
-Solar Power (Australia) Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P
-Solar Monash (Godahewa et al., 2021) Energy W ZS P P ZS
-Solar Monash (Godahewa et al., 2021) Energy 10M P P P
-Sunspot Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P P ZS P ZS P
-Temperature Rain Monash (Godahewa et al., 2021) Nature D T/T T/T T/T P P ZS P
-Tourism Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS P T/T ZS ZS P
-Tourism Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS ZS P T/T ZS ZS P
-Tourism Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS ZS P P P T/T ZS P
-Traffic Hourly Monash (Godahewa et al., 2021) Transport H ZS P T/T T/T T/T P T/T ZS ZS ZS ZS T/T P
-Traffic Weekly Monash (Godahewa et al., 2021) Transport W ZS T/T T/T ZS P ZS ZS P
-US Births Monash (Godahewa et al., 2021) Healthcare D ZS T/T T/T P P ZS P ZS P
-US Births Monash (Godahewa et al., 2021) Healthcare W-TUE
-US Births Monash (Godahewa et al., 2021) Healthcare M
-Vehicle Trips Monash (Godahewa et al., 2021) Transport D T/T T/T P ZS P
-Weather (Australia) Monash (Godahewa et al., 2021) Climate D ZS ZS T/T T/T P ZS ZS P T/T T/T P
-Wind Farms Monash (Godahewa et al., 2021) Energy T T/T T/T P P
-Wind Farms Monash (Godahewa et al., 2021) Energy H P P
-Wind Farms Monash (Godahewa et al., 2021) Energy D P
-Wind Farms Monash (Godahewa et al., 2021) Energy 5T P
-Wind Power Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P
-Kaggle Web Traffic Daily Monash (Godahewa et al., 2021) Web D P
-Restaurant Recruit Other Sales D
-Subseasonal Other Climate D P P P
-Subseasonal Precipitation Other Climate D P P P
-USHCN Other Weather D P
-Mexico City Bikes Other Transport H P P
-Dominick Other Sales D ZS P P
-ERCOT Load Other Energy H ZS P
+Solar LSTNet (Lai et al., 2017) Energy 10T ZS T/T T/T ZS T/T
+Solar LSTNet (Lai et al., 2017) Energy 5T P possible P P P P
+Solar LSTNet (Lai et al., 2017) Energy H P ZS ZS ZS ZS P P T/T ZS P
+Solar LSTNet (Lai et al., 2017) Energy D ZS T/T ZS
+Solar LSTNet (Lai et al., 2017) Energy W-FRI ZS T/T ZS
+Exchange Rate LSTNet (Lai et al., 2017) Econ/Fin D ZS ZS T/T ZS ZS T/T P T/T ZS ZS
+Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales D P ZS T/T ZS
+Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales W-WED ZS T/T ZS
+Australian Electricity Demand Monash (Godahewa et al., 2021) Energy 30T ZS ZS T/T T/T P P P ZS P ZS P P P T/T P P ZS ZS P
+Bitcoin Monash (Godahewa et al., 2021) Econ/Fin D ZS T/T T/T P P ZS P ZS P P P T/T P P P
+Car Parts Monash (Godahewa et al., 2021) Sales M T/T T/T P ZS P ZS T/T T/T ZS ZS ZS
+CIF 2016 Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P P T/T P P ZS ZS
+COVID Deaths Monash (Godahewa et al., 2021) Healthcare D ZS ZS T/T T/T P ZS ZS P ZS T/T T/T ZS ZS ZS
+Covid Mobility Monash (Godahewa et al., 2021) Transport D T/T T/T P P T/T P P
+Elecdemand Monash (Godahewa et al., 2021) Energy 30T T/T T/T P P T/T P P
+Extended Web Traffic Monash (Godahewa et al., 2021) Web D T/T T/T P P P P T/T P P
+FRED MD Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P P P T/T P P ZS ZS P
+Hospital Monash (Godahewa et al., 2021) Healthcare M ZS ZS T/T T/T P ZS ZS P ZS T/T T/T ZS ZS ZS
+Kaggle Web Traffic Weekly Monash (Godahewa et al., 2021) Web W P P P P P P P P P P P
+KDD Cup 2018 Monash (Godahewa et al., 2021) Nature H T/T T/T T/T P P P ZS P P ZS T/T T/T T/T ZS P
+KDD Cup 2018 Monash (Godahewa et al., 2021) Nature D ZS P T/T ZS P
+London Smart Meters Monash (Godahewa et al., 2021) Energy 30T T/T P P P P P P P P P P T/T P P P P P
+M1 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T P ZS P P T/T P P ZS ZS
+M1 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P P P P P P P ZS ZS
+M1 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P P P P P P P ZS ZS
+M3 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T ZS T/T ZS P P T/T P P ZS ZS
+M3 Other Monash (Godahewa et al., 2021) Econ/Fin Q T/T T/T T/T ZS P T/T P P
+M3 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P ZS T/T P P P P P ZS ZS
+M3 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P ZS T/T P P P P P ZS ZS
+M4 Daily Monash (Godahewa et al., 2021) Econ/Fin D T/T P T/T T/T P T/T ZS P ZS T/T T/T T/T ZS P P
+M4 Hourly Monash (Godahewa et al., 2021) Econ/Fin H T/T P T/T T/T P T/T ZS P ZS T/T T/T T/T ZS P P
+M4 Monthly Monash (Godahewa et al., 2021) Econ/Fin M T/T P T/T T/T ZS T/T ZS P ZS T/T T/T T/T ZS P P
+M4 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q-DEC ZS P P P ZS T/T P ZS P T/T ZS ZS P
+M4 Weekly Monash (Godahewa et al., 2021) Econ/Fin W-SUN T/T P T/T T/T T/T ZS P ZS T/T T/T T/T ZS P
+M4 Yearly Monash (Godahewa et al., 2021) Econ/Fin A-DEC ZS P P P ZS T/T P ZS P T/T ZS ZS P
+NN5 Daily Monash (Godahewa et al., 2021) Econ/Fin D ZS ZS T/T T/T P ZS P ZS P P P T/T P P ZS ZS P
+NN5 Weekly Monash (Godahewa et al., 2021) Econ/Fin W ZS ZS T/T T/T P ZS ZS P T/T P P ZS ZS
+Oikolab Weather Monash (Godahewa et al., 2021) Climate H T/T T/T P P T/T P P
+Pedestrian Counts Monash (Godahewa et al., 2021) Transport H T/T ZS T/T T/T P ZS P ZS ZS P P P T/T T/T P P P P
+Rideshare Monash (Godahewa et al., 2021) Transport H T/T T/T T/T P ZS P P T/T T/T P P P
+Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer VisionTS TTM LLMTime Time-LMM UniTime Time-MoE Chronos-2 Flowstate(under review) LightGTS Sundial TabPFN-TS Time-MoE2 TinyTimeMixer TiRex Toto YingLong (under review) Gift-Eval TOTO2 Flowstate(T) TimesFM-ICL ROSE
+Saugeen Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P ZS P ZS P ZS P T/T T/T ZS P
+Saugeen Monash (Godahewa et al., 2021) Nature W-THU ZS T/T ZS
+Saugeen Monash (Godahewa et al., 2021) Nature M ZS T/T ZS
+Solar Power (Australia) Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P P P T/T P P P P P
+Solar Monash (Godahewa et al., 2021) Energy W ZS P P ZS ZS
+Solar Monash (Godahewa et al., 2021) Energy 10M P P P ZS
+Sunspot Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P P ZS P ZS P P P T/T P P P
+Temperature Rain Monash (Godahewa et al., 2021) Nature D T/T T/T T/T P P ZS P ZS P T/T T/T T/T ZS P P
+Tourism Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS P T/T ZS ZS P P P P ZS ZS
+Tourism Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS ZS P T/T ZS ZS P P P P ZS ZS
+Tourism Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS ZS P P P T/T ZS P P P P P ZS ZS
+Traffic Hourly Monash (Godahewa et al., 2021) Transport H ZS P T/T T/T T/T P T/T ZS ZS ZS ZS T/T P P T/T T/T P P ZS P T/T
+Traffic Weekly Monash (Godahewa et al., 2021) Transport W ZS T/T T/T ZS P ZS ZS P P T/T P P
+US Births Monash (Godahewa et al., 2021) Healthcare D ZS T/T T/T P P ZS P ZS P ZS P T/T T/T P
+US Births Monash (Godahewa et al., 2021) Healthcare W-TUE ZS T/T ZS
+US Births Monash (Godahewa et al., 2021) Healthcare M ZS T/T ZS
+Vehicle Trips Monash (Godahewa et al., 2021) Transport D T/T T/T P ZS P P T/T P P
+Weather (Australia) Monash (Godahewa et al., 2021) Climate D ZS ZS T/T T/T P ZS ZS P T/T T/T P P P T/T P P ZS P
+Wind Farms Monash (Godahewa et al., 2021) Energy T T/T T/T P P P P T/T P P P
+Wind Farms Monash (Godahewa et al., 2021) Energy H P P P P P
+Wind Farms Monash (Godahewa et al., 2021) Energy D P possible P P P P
+Wind Farms Monash (Godahewa et al., 2021) Energy 5T P possible P
+Wind Power Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P P P T/T P P P P P P
+Kaggle Web Traffic Daily Monash (Godahewa et al., 2021) Web D P P P P
+Restaurant Recruit Other Sales D ZS T/T ZS
+Subseasonal Other Climate D P P P P P P P P P P
+Subseasonal Precipitation Other Climate D P P P P P P P P P
+USHCN Other Weather D P possible P P P P P
+Mexico City Bikes Other Transport H P P possible P P P P P
+Dominick Other Sales D ZS P P P ZS ZS
+ERCOT Load Other Energy H ZS P P ZS ZS
 Weather (US) Other Weather 10T P ZS
@@
 Beijing PM 2.5 Other Nature H ZS
-TDBrain Other Healthcare 0.002S P
+TDBrain Other Healthcare 0.002S P P P
 SensorData Other IoT 0.02S P
-BDG-2 Bull ProEnFo (Wang et al., 2023b) Energy H P P P
-BDG-2 Cockatoo ProEnFo (Wang et al., 2023b) Energy H P P P
-BDG-2 Hog ProEnFo (Wang et al., 2023b) Energy H P P P
-Covid19 Energy ProEnFo (Wang et al., 2023b) Energy H P P P
-ELF ProEnFo (Wang et al., 2023b) Energy H P P
-GEF12 ProEnFo (Wang et al., 2023b) Energy H P P P
-GEF14 ProEnFo (Wang et al., 2023b) Energy H P P P
-GEF17 ProEnFo (Wang et al., 2023b) Energy H P P
-PDB ProEnFo (Wang et al., 2023b) Energy H P P P
-Spanish ProEnFo (Wang et al., 2023b) Energy H P P P P
+BDG-2 Bull ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
+BDG-2 Cockatoo ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
+BDG-2 Hog ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
+Covid19 Energy ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
+ELF ProEnFo (Wang et al., 2023b) Energy H P P P P P P
+GEF12 ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
+GEF14 ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
+GEF17 ProEnFo (Wang et al., 2023b) Energy H P P P P P P
+PDB ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
+Spanish ProEnFo (Wang et al., 2023b) Energy H P P P P P P P P P P
 Australia Rainfall Tan et al., 2021 Nature H P
@@
 Dodger TSB-UAD (Paparrizos et al., 2022) Traffic - P
-ECG TSB-UAD (Paparrizos et al., 2022) Healthcare - P
+ECG TSB-UAD (Paparrizos et al., 2022) Healthcare - P P
 IOPS TSB-UAD (Paparrizos et al., 2022) Web - P
@@
 GHL TSB-UAD (Paparrizos et al., 2022) Sensor - P
-Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
 Genesis TSB-UAD (Paparrizos et al., 2022) Sensor - P
@@
 SZSDB TSB-UAD (Paparrizos et al., 2022) Healthcare - P
-Electricity UCI ML Archive (Trindade, 2015) Energy 15T T/T P T/T P
-Electricity UCI ML Archive (Trindade, 2015) Energy H T/T P ZS ZS T/T P T/T ZS ZS ZS ZS T/T T/T
-Electricity UCI ML Archive (Trindade, 2015) Energy D ZS
-Electricity UCI ML Archive (Trindade, 2015) Energy W-FRI T/T
-Electricity UCI ML Archive (Trindade, 2015) Energy M T/T
-Electricity UCI ML Archive (Trindade, 2015) Energy W P P
+Electricity UCI ML Archive (Trindade, 2015) Energy 15T T/T P T/T P ZS T/T T/T T/T ZS P
+Electricity UCI ML Archive (Trindade, 2015) Energy H T/T P ZS ZS T/T P T/T ZS ZS ZS ZS T/T T/T ZS T/T ZS T/T T/T ZS P P T/T
+Electricity UCI ML Archive (Trindade, 2015) Energy D ZS ZS T/T ZS
+Electricity UCI ML Archive (Trindade, 2015) Energy W-FRI T/T ZS T/T T/T ZS
+Electricity UCI ML Archive (Trindade, 2015) Energy M T/T T/T
+Electricity UCI ML Archive (Trindade, 2015) Energy W P P P
 SemgHandGenderCh2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
@@
 UWaveGestureLibraryAll UCR Archive (Ismail Fawaz et al., 2019) Human - T/T
-SelfRegulationSCP1 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
+SelfRegulationSCP1 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
 UWaveGestureLibraryX UCR Archive (Ismail Fawaz et al., 2019) Human - P
 GesturePebbleZ2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
-Healthcare5000 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - T/T
+Healthcare5000 (orignally: ECG5000) UCR Archive (Ismail Fawaz et al., 2019) Healthcare - T/T P
 OSULeaf UCR Archive (Ismail Fawaz et al., 2019) Nature - T/T
@@
 Japanesevowels UCR Archive (Ismail Fawaz et al., 2019) Audio - P
-Worms UCR Archive (Ismail Fawaz et al., 2019) Nature - P
+Worms UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P
 Ham UCR Archive (Ismail Fawaz et al., 2019) Facilities - T/T
@@
 MiddlePhalanxTW UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-AtrialFibrillation UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
+AtrialFibrillation UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
 SyntheticControl UCR Archive (Ismail Fawaz et al., 2019) Generated - P
@@
 FaceFour UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-Phoneme UCR Archive (Ismail Fawaz et al., 2019) Audio - P P
+Phoneme UCR Archive (Ismail Fawaz et al., 2019) Audio - P P P P
 InlineSkate UCR Archive (Ismail Fawaz et al., 2019) Human - P
 CricketX UCR Archive (Ismail Fawaz et al., 2019) Human Body - P
-SelfRegulationSCP2 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
+SelfRegulationSCP2 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
 DistalPhalanxOutlineCorrect UCR Archive (Ismail Fawaz et al., 2019) Nature - P
@@
 ERing UCR Archive (Ismail Fawaz et al., 2019) Human - P
-StarLightCurves UCR Archive (Ismail Fawaz et al., 2019) Nature - P P
+StarLightCurves UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P P
 WordSynonyms UCR Archive (Ismail Fawaz et al., 2019) Audio - P
@@
 Herring UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
 Lightning2 UCR Archive (Ismail Fawaz et al., 2019) Nature - P
@@
 HandMovementDirection UCR Archive (Ismail Fawaz et al., 2019) Human - P
+Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer VisionTS TTM LLMTime Time-LMM UniTime Time-MoE Chronos-2 Flowstate(under review) LightGTS Sundial TabPFN-TS Time-MoE2 TinyTimeMixer TiRex Toto YingLong (under review) Gift-Eval TOTO2 Flowstate(T) TimesFM-ICL ROSE
 MoteStrain UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
@@
 BME UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-EigenWorms UCR Archive (Ismail Fawaz et al., 2019) Nature - P P
+EigenWorms UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P P
 FordB UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
@@
 CinCHealthcareTorso UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
-PigArtPressure UCR Archive (Ismail Fawaz et al., 2019) Nature - P P
+PigArtPressure UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P P
 Fish UCR Archive (Ismail Fawaz et al., 2019) Nature - P
@@
 FiftyWords UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-MotorImagery UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
+MotorImagery UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
 ToeSegmentation1 UCR Archive (Ismail Fawaz et al., 2019) Human - P
@@
 ProximalPhalanxTW UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
 DuckDuckGeese UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
@@
 ToeSegmentation2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
-PigCVP UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
+PigCVP UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
 CricketY UCR Archive (Ismail Fawaz et al., 2019) Human - P
@@
 Healthcare200 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
-Weatherbench (Daily) Weather Bench Weather D P P
-Weatherbench (Hourly) Weather Bench Weather H P P
-Weatherbench (Weekly) Weather Bench Weather W P P
-Wiki hourly Wikipedia Metrics Web H P
-Wiki daily Wikipedia Metrics Web D P P P
-Wiki weekly Wikipedia Metrics Web W P
-Wiki monthly Wikipedia Metrics Web M P
+Weatherbench (Daily) Weather Bench Weather D P P possible P P P P P
+Weatherbench (Hourly) Weather Bench Weather H P P possible P P P P P
+Weatherbench (Weekly) Weather Bench Weather W P P possible P P P P P
+Wiki hourly Wikipedia Metrics Web H P P
+Wiki daily Wikipedia Metrics Web D P P P possible P P P P P
+Wiki weekly Wikipedia Metrics Web W P P
+Wiki monthly Wikipedia Metrics Web M P P
 Global Temp Corrformer (Wu et al., 2023) Weather 10T T/T
@@
 TaxiBJ Zhang et al., 2017 Traffic and weather 30T ZS
+ERA5 Munoz-Sabater et al., 2021 Weather 3h P P
+ERA5 Munoz-Sabater et al., 2021 Weather 12h P
+ERA5 Munoz-Sabater et al., 2021 Weather D P
+ERA5 Munoz-Sabater et al., 2021 Weather W P
+ERA5 Munoz-Sabater et al., 2021 Weather M P
+ERA5 Munoz-Sabater et al., 2021 Weather Q P
+Finance (not published) Liu et al., 2025 (Sundial) Finance ? P
+IoT (not published) Liu et al., 2025 (Sundial) IoT ? P
+BOOM Cohen et al. (2025) (TOTO) Observational 10S ZS
+BOOM Cohen et al. (2025) (TOTO) Observational T ZS
+BOOM Cohen et al. (2025) (TOTO) Observational 5T ZS
+BOOM Cohen et al. (2025) (TOTO) Observational 10T ZS
+BOOM Cohen et al. (2025) (TOTO) Observational 30T ZS
+BOOM Cohen et al. (2025) (TOTO) Observational H ZS
+BOOM Cohen et al. (2025) (TOTO) Observational D

### v3

@@
-Challenges and Requirements for Benchmarking Time Series Foundation Models
+Rethinking Evaluation in the Era of Time Series Foundation Models: (Un)known Information Leakage Challenges
 Abstract
@@
 However, similar to Large Language Models (LLMs), the evaluation of TSFMs is challenging: as training corpora grow increasingly large, it becomes difficult to ensure the integrity of the test sets used for benchmarking.
-Our investigation of existing TSFM evaluation studies identifies two kinds of information leakage: (1) train-test sample overlaps arising from the multi-purpose reuse of datasets and (2) temporal overlap of correlated train and test series.
+An investigation of existing TSFM evaluation studies identifies two kinds of information leakage: (1) train-test sample overlaps arising from the multi-purpose reuse of datasets and (2) temporal overlap of correlated train and test series.
 Ignoring these forms of information leakage when benchmarking TSFMs risks producing overly optimistic performance estimates that fail to generalize to real-world settings.
@@
 Time Series Foundation Models (TSFMs) represent an emerging paradigm in forecasting, drawing inspiration from the architecture and training methodologies of foundation models in natural language processing (NLP).
-Despite being a relatively new research area, the field has evolved into a highly dynamic landscape featuring a wide variety of architectures and approaches.
-This rapid proliferation is exemplified by diverse models such as Chronos (Ansari et al., 2024a), Chronos-bolt (Ansari et al., 2024b), Chronos-2 (Ansari et al., 2025), Moirai/Moirai-MoE (Woo et al., 2024; Liu et al., 2024b), MOMENT (Goswami et al., 2024), Time-MoE (Shi et al., 2024), TimesFM (Das et al., 2024), TimesFM-2.0 (Das et al., 2024), TimesFM-2.5 (Das et al., 2024), or TiRex (Auer et al., 2025).
-In contrast to traditional time series models that are trained from scratch for specific tasks and time series, TSFMs leverage transfer learning: they are pre-trained on large collections of generic time series and can perform zero-shot forecasting without domain adaptation or fine-tuning (Liang et al., 2024).
-The reliance on massive amounts of time series training data, typically scraped from public repositories, raises the question of whether the proper evaluation of TSFMs faces similar challenges to those observed in the evaluation of Large Language Models (LLMs).
+In contrast to traditional time series models, TSFMs are pre-trained on large time series corpora, enabling zero-shot forecasting without task-specific adaptation (Liang et al., 2024).
+Over the last years, a highly dynamic landscape with a rapidly growing family of TSFMs has eveolved (Ansari et al., 2024; Auer et al., 2025; Cohen et al., 2025; Das et al., 2024).
+Yet, the very property that makes these TSFMs powerful, training globally on the world’s time series, creates structural evaluation problems, akin to issues recently observed in the evaluation of large language models (LLMs).
 In the NLP domain, training on vast portions of the internet has given rise to an “evaluation crisis" (Liao and Xiao, 2023), in which test set contamination (Mirzadeh et al., 2024; Ravaut et al., 2024; Li et al., 2024a) and memorization effects (Chang et al., 2024) have led to overly optimistic performance estimates.
-Whether current TSFM evaluations are vulnerable to analogous issues remains an open question, with important implications for fair and reliable model comparison.
-While benchmark strategies such as benchmarking on held-out data (Aksu et al., 2024) and clean train-test splits (Qiu et al., 2024) have been proposed recently, we doubt that these approaches effectively prevent information leakage in the context of TSFMs.
-Beyond the risk of direct information leakage, the increased memorization capacity of scaled foundation models poses additional challenges (see Section 3).
-This article examines the validity and reliability of current TSFM evaluation practices and discusses the risk of an evaluation crisis in the TSFM domain.
-We address this issue through two contributions: first, we conduct a comprehensive train-test set lineage analysis of 22 leading TSFMs and review all major evaluation benchmarks available as of early 2026.
-Second, we empirically demonstrate the phenomenon of temporal overlaps of correlated train and test series, a challenge specific to TSFMs, through a dedicated experiment.
-Our investigation yields two primary findings that suggest that TSFM evaluation is indeed compromised:
-  • First, while we find no evidence of unintended direct information leakage within the training and evaluation of individual TSFMs, our analysis reveals substantial ambiguity in data separation between models.
-In several cases, one model’s training set serves as another’s test set.
-This issue is further exacerbated by dataset remixing, renaming, and inadequate documentation, which already pose significant evaluation challenges for newer models (see Section 3.1).
-  • Second, we identify the temporal overlap of correlated train and test series as an indirect form of information leakage specific to TSFMs.
-External shocks, such as the COVID-19 pandemic to stock prices, as well as exogenous drivers like local weather conditions to local solar production, act as factors that directly shape or distort the causal dynamics between multiple time series (see Section 3.2).
-As a result, a TSFM trained and evaluated on disjoint time series covering the same time period can effectively memorize patterns induced by such causal relations (see Figure 3).
-Based on these findings, we argue that the rise of TSFMs necessitates new benchmarking methodologies to prevent an evaluation crisis akin to that observed in NLP.
-2 Background
-2.1 Time Series Foundation Models
-A foundation model is defined as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted or fine-tuned to a wide range of downstream tasks” (Bommasani et al., 2021).
-The term is primarily associated with Transformer-based Large Language Models (LLMs) like the ChatGPT or Llama families (OpenAI, 2023; Grattafiori et al., 2024).
-Their success is largely attributed to scaling laws (Kaplan et al., 2020), describing the practice of increasing training data, parameters, and compute to improve a model’s predictive performance.
-Although LLMs demonstrate impressive zero-shot capabilities across benchmarks (Liang et al., 2022), concerns persist that they partly rely on memorizing benchmark data (Chang et al., 2024; Carlini et al., 2022).
-Such test set contamination, where test samples appear in pre-training data like Common Crawl [^1] https://commoncrawl.org/ or The Pile [^2] https://pile.eleuther.ai/, remains common despite improved dataset curation, leading to inflated performance estimates (Li et al., 2024a; Singh et al., 2024; Grattafiori et al., 2024; Mirzadeh et al., 2024; Ravaut et al., 2024).
-Recognizing the structural similarity between language modeling and time series analysis, recent research has adapted the Transformer architecture to forecast future values of time series, ranging from the classical encoder-decoder Transformer (Ansari et al., 2024a) over Decoder-only models (Das et al., 2024) to Encoder-only models often applied in multivariate time series forecasting (Goswami et al., 2024).
-Beyond this, TSFMs are based on a variety of other architectural backbones such as LLMs Gruver et al.
-(2024); Xue and Salim (2024), Vision-based Autoencoders Chen et al.
-(2024), and mixed architectures that leverage LLM reprogramming Jin et al.
-(2024).
-Across these approaches, a key commonality lies in how data are represented: LLM-based models encode numerical values directly as language tokens like LLMTime (Gruver et al., 2024) and PromptCast (Xue and Salim, 2024), whereas some Transformer-based variants rely on specialized tokenization schemes to discretize continuous time series, as exemplified by TimeGPT-1 (Garza et al., 2024) and the T5-based Chronos (Ansari et al., 2024a).
-Furthermore, Vision-based approaches represent time series as gray-scaled images, while reprogramming-based approaches mix textual and continuous time series representations Jin et al.
-(2024).
-Similar to LLMs, all presented architectures are pre-trained on large collections of generic time series and appear to follow scaling laws (Edwards et al., 2024), enabling them to produce zero-shot forecasts and handle a wide range of frequencies and forecasting horizons.
-Only one exception—MOMENT—requires task-adaptive fine-tuning for high-quality forecasts Goswami et al.
-(2024).
-Based on an encoder-only architecture this model is capable of time series regression, classification, anomaly detection, and forecasting Goswami et al.
-(2024).
-However, due to its broad application area, the model needs task-adaptive fine-tuning of the last neural network layer that is specialized for each task.
-A key distinction between the foundation models in NLP and time series analysis concerns the definition of “zero-shot” capabilities.
-While LLMs typically adapt to entirely different tasks (Brown et al., 2020), TSFMs typically perform zero-shot forecasting by generalizing across domains, time series, input lengths, frequencies, and horizons without the need for model retraining (Ansari et al., 2024a; Das et al., 2024; Woo et al., 2024).
-2.2 Benchmarking Methodologies
-Traditional statistical or machine-learning-based models for time series forecasting generally require explicit training on the target time series.
-These local models exhibit only limited generalization capabilities; that is, they can only predict values that are identically distributed to their training data.
-Consequently, the dominant logic in time series benchmarking relies on time-based train/test splits (Hyndman and Athanasopoulos, 2021).
-This approach creates a temporal separation between training and evaluation data to prevent information leakage and is employed in major time series collections such as the Monash repository (Godahewa et al., 2021) and TFB (Qiu et al., 2024).
-To rigorously assess a model’s predictive capabilities under evolving temporal dynamics, Time-Series Cross-Validation (TSCV) is widely regarded as best practice.
-Instead of using a single time-based split, TSCV uses the idea of rolling windows (see left in Figure 1).
-This evaluation strategy does not necessarily require retraining; instead, the training set can be fixed while the inference window is rolled forward, reducing evaluation costs to inference-only computation (Hewamalage et al., 2023).
-Forecasting competitions, like the M-competitions or Kaggle, try to avoid the risk of test set contamination by soliciting forecasts for unpublished private test sets (Makridakis et al., 2022; Bojer and Meldgaard, 2021).
-These competitions are typically held at irregular intervals.
-While all of the above evaluation strategies rely on historical data for testing, the recent M6 competition introduced a live evaluation using future values of financial assets on selected dates as the target (Makridakis et al., 2024).
-Figure 1: Evaluation strategies along the time (left) or the domain dimension (right).
-While current TSFM benchmarking studies also rely on historical evaluation data, their logic diverges from the above described time-series cross-validation strategy.
-Instead of using TSCV on one series as illustrated in Figure 1 (a), they typically evaluate models using many series but with only a single time-based split per series (Das et al., 2024; Gruver et al., 2024).
-Figure 1 (b) shows the focus of most TSFM evaluations.
-Here, the focus is not on temporal depth, but on generalization across heterogeneous series, especially in zero-shot scenarios where complete datasets or domains are held out of the pre-training (Shi et al., 2024; Ansari et al., 2024a).
-This approach is focused on the generalization capabilities of TSFMs and prevents cherry-picking of test sets.
-Two prominent examples of this strategy are TSFM-Bench, which comprises 21 test sets with time series from diverse domains (e.g., energy, finance), statistical characteristics (e.g., stationary, trend), and frequencies (e.g., quarter-hourly, hourly) Li et al.
-(2025) and GIFT-Eval, which comprises 23 curated test sets with time series of seven domains, ten frequencies, and multivariate inputs (Aksu et al., 2024).
-While this series cross validation focuses on evaluating the generalization capabilities of TSFMs, it is not without drawbacks.
-Roque et al.
-(2025) showed that with just four cherry-picked test sets, out of the distribution of available test sets, 46% of models could appear as state-of-the-art purely due to selection bias.
-Notably, deep learning models were found to be more susceptible to this variance than classical baselines.
-Especially when evaluation general-purpose TSFMs, generalization is a core evaluation focus.
-Figure 1 (c) illustrates the current best practice in TSFM evaluation, that is, a combined time and series cross validation (Yu et al., 2016; Salinas et al., 2020).
-While this approach has first adopters, notably Goktas et al.
-(2025b); Shchur et al.
-(2025), it is not yet commonly used in TSFM benchmarking.
-The scaling of pre-training and evaluation datasets in TSFMs gives rise to another challenge that can also be observed in LLM evaluation.
-As LLMs are pre-trained on large datasets crawled from the internet like Common Crawl [^3] https://commoncrawl.org/ or The Pile [^4] https://pile.eleuther.ai/, they have already seen most benchmark tasks during pre-training.
-Ultimately, this led to an LLM “evaluation crisis” Liao and Xiao (2023).
-Likewise, with large and obscure pre-training datasets, defining test datasets that do not overlap with models’ pre-training datasets becomes increasingly difficult in TSFMs Aksu et al.
-(2024); Goktas et al.
-(2025a); Xu et al.
-(2025).
-As shown in the GIFT-Eval ablation studies, information leakage can strongly inflate evaluation performance of TSFMs Aksu et al.
-(2024).
-Thus, investigating information leakage in TSFMs is an important research direction.
-Addressing this, current benchmarks developed strategies to prevent information leakage.
-GIFT-Eval provides an explicit pretraining dataset containing approximately 230 billion data points and a dedicated test with 177 million data points Aksu et al.
-(2024) staying with a series cross validation (see Figure 1 (a)) but no time cross validation.
-Two other benchmarks TempusBench and Fidel-TS, use test datasets from unknown or restricted sources to prevent information leakage Goktas et al.
-(2025a); Xu et al.
-(2025).
-Particularly, TempusBench leverages a new test set with 48 uni- and multi-variate time series forecasting tasks that is currently not in any TSFM pre-training corpus Goktas et al.
-(2025a).
-While they apply a time and series cross-validation strategy, they further propose to leverage real-world live data for evaluations Goktas et al.
-(2025a).
-Fidel-TS apply a similar strategy by leveraging restricted live APIs but implement a series cross validation without time cross validation Xu et al.
-(2025).
-Overall, all current approaches to mitigate information leakage and the possibly resulting inflation of performance rely on the development of new datasets that can currently not be in the pre-training corpus of TSFMs.
-3 Challenges of Evaluating Time Series Foundation Models
-Based on the need for an examination of information leakage in current TSFMs, we decided to first conduct a TSFM train-test dataset lineage analysis (Section 3.1).
-We then further analyze the outcomes of potential indirect overlaps in time series between pre-training and test sets of TSFMs.
-3.1 Direct Information Leakage: Train-Test Sample Overlap
-To examine potential test set contamination in existing TSFM evaluations, we searched for peer-reviewed and accepted papers including the terms "Time Series Foundation Model" at leading venues in the machine learning area (TMLR, ICML, NeurIPS, ICLR, WWW) and their workshops, published up to January 2026.
-This search resulted in 22 published TSFMs that met our inclusion criteria (Table 1).
-For each of these papers, we analyzed their dataset lineage, categorizing the used datasets into pre-training, train/test and zero-shot.
+As this perspective will show, current TSFM evaluations are vulnerable to analogous issues, leading to important implications for fair and reliable forecasting evaluation in the era of TSFMs.
+While novel benchmark strategies such as benchmarking on held-out data (Aksu et al., 2024) and clean train-test splits (Qiu et al., 2024) have been proposed recently, these approaches cannot address the fundamental sources of information leakage in TSFMs.
+We identify two such sources (see Section 3): First, direct information leakage, arising from the multi-purpose reuse of public datasets across model training and evaluation pipelines.
+Second, indirect information leakage, arising from temporal overlap between correlated training and test series that often share a common causal driver — such as the COVID-19 pandemic simultaneously distorting correlated financial time series across geographies.
+Together, these sources risk turning current TSFM benchmarks into measures of memorization rather than generalization.
+Our investigation reveals that information leakage is already present in TSFM benchmarking.
+Tracing the dataset lineage of 22 published TSFMs reveals no community consensus on which data should be used for training versus evaluation: in several documented cases, one model’s pre-training corpus is another model’s test set, making direct cross-model comparison nearly impossible.
+Yet, even benchmarks that carefully avoid such direct overlap remain exposed to indirect information leakage: a model trained on global stock indices through 2020 can learn the COVID-19 crash and exploit that structure when asked to forecast any correlated series from the same period, even one it has never directly seen (see Section 3).
+The TSFM field risks repeating, in compressed time, the evaluation crisis that has undermined trust in LLM benchmarking.
+Preventing this doesn’t require incremental fixes to existing benchmarks, but a principled rethinking of what valid evaluation means when a model has been trained on the world’s time series.
+2 Time Series Foundation Models: The Current State of Evaluation
+Understanding the current state of TSFM evaluation requires first appreciating what distinguishes these models architecturally from classical forecasting methods, and how the benchmarking practices developed for those methods have been adapted to assess them.
+2.1 Time Series Foundation Models: Global Training, Zero-Shot Inference
+The foundation model paradigm, originally defined as training on broad data at scale to enable adaptation across downstream tasks (Bommasani et al., 2021), found its most prominent expression in LLMs such as the GPT and Llama families (OpenAI, 2023; Grattafiori et al., 2024), whose performance scales predictably with training data, parameters, and compute (Kaplan et al., 2020).
+Recognizing the structural sequence similarity between language modeling and time series analysis, recent work has adapted this paradigm to forecasting, producing a growing family of Transformer-based architectures that range from encoder-decoder designs (Ansari et al., 2024) to decoder-only (Das et al., 2024) and encoder-only models (Goswami et al., 2024).
+Since the emergence of TSFMs a few years ago, 22 models have been presented at top venues including TMLR, ICML, NeurIPS, ICLR, and WWW, as shown in Table 1.
 Table 1: Time series foundation models published in recent years
@@
 Time-LLM Jin et al. (2024) 2024 ICLR 24
-Chronos Ansari et al. (2024a) 2024 TMLR
+Chronos Ansari et al. (2024) 2024 TMLR
 TimesFM Das et al. (2024) 2024 ICML 24
@@
 Toto Cohen et al. (2025) 2025 NeurIPS 25
-Pre-training datasets are used to fit a model to general time series patterns, such as, trends and seasonality.
-They are employed for initial training only and, hence, should be excluded from any evaluation activities.
-Train/test datasets, also sometime called in-domain or in-distribution datasets in the context of TSFMs, are datasets used to fine-tune a model to domain-specific patterns.
-They are split at a defined point in time into training and test sets and the model is fitted on the training data and evaluated on the test data (Hyndman and Athanasopoulos, 2021; Kapoor and Narayanan, 2023).
-Zero-shot datasets are used for evaluation of TSFMs on unseen data and should therefore be excluded from any training and fine-tuning.
-In a zero-shot setting, a model is given a short historical context of an unseen time series and it is assessed how well it can predict future values of this series.
+Beyond the Transformer backbone, TSFMs draw on a range of architectural approaches: LLM-based models encode time series values directly as language tokens (Gruver et al., 2024; Xue and Salim, 2024), vision-based models represent them as grayscaled images (Chen et al., 2024), and reprogramming-based architectures mix textual and continuous representations (Jin et al., 2024).
+Despite this diversity, a key commonality runs through almost all TSFMs: pre-training on large, heterogeneous collections of time series that appear to follow scaling laws (Edwards et al., 2024), enabling zero-shot forecasting across a wide range of domains, frequencies, and horizons without retraining (MOMENT is the notable exception, requiring task-adaptive fine-tuning of its output layer (Goswami et al., 2024)).
+This zero-shot capability differs fundamentally from its NLP counterpart: while LLMs generalize to entirely different tasks (Brown et al., 2020), TSFMs generalize across domains, series, input lengths, and horizons, extending to new time series rather than new problem types (Ansari et al., 2024; Das et al., 2024; Woo et al., 2024).
+2.2 Classical and TSFM Benchmarking Practices
+Traditional statistical or machine-learning-based models for time series forecasting generally require explicit training on the target time series.
+These local models exhibit only limited generalization capabilities; that is, they can only predict values that are identically distributed to their training data.
+Consequently, the dominant logic in time series benchmarking relies on time-based train/test splits (Hyndman and Athanasopoulos, 2021).
+This approach creates a temporal separation between training and evaluation data to prevent information leakage and is employed in major time series collections such as the Monash repository (Godahewa et al., 2021) and TFB (Qiu et al., 2024).
+To rigorously assess a model’s predictive capabilities under evolving temporal dynamics, Time-Series Cross-Validation (TSCV) is widely regarded as best practice.
+Instead of using a single time-based split, TSCV uses the idea of rolling windows (see Figure 1 (a)).
+This evaluation strategy does not necessarily require retraining; instead, the training set can be fixed while the inference window is rolled forward, reducing evaluation costs to inference-only computation (Hewamalage et al., 2023).
+Forecasting competitions, like the M-competitions or Kaggle, try to avoid the risk of test set contamination by soliciting forecasts for unpublished private test sets (Bojer and Meldgaard, 2021; Makridakis et al., 2022).
+These competitions are typically held at irregular intervals.
+The recent M6 competition took this further, introducing live evaluation using future values of financial assets as the target, bypassing reliance on historical data entirely (Makridakis et al., 2024).
+Figure 1: Evaluation strategies along the time (left) or the domain dimension (right).
+While current TSFM benchmarking studies also rely on historical data for evaluation, their logic diverges from the above described time-series cross-validation strategy.
+Instead of using TSCV on one series as illustrated in Figure 1 (a), they typically evaluate models using many series but with only a single time-based split per series (Das et al., 2024; Gruver et al., 2024).
+Figure 1 (b) shows the focus of most TSFM evaluations.
+Here, the focus is not on temporal depth, but on generalization across heterogeneous series, especially in zero-shot scenarios where complete datasets or domains are held out of the pre-training (Shi et al., 2024; Ansari et al., 2024).
+This approach is focused on the generalization capabilities of TSFMs and prevents cherry-picking of test sets.
+Two prominent examples of this strategy are TSFM-Bench, which comprises 21 test sets with time series from diverse domains (e.g., energy, finance), statistical characteristics (e.g., stationary, trend), and frequencies (e.g., quarter-hourly, hourly) Li et al.
+(2025) and GIFT-Eval, which comprises 23 curated test sets with time series of seven domains, ten frequencies, and multivariate inputs (Aksu et al., 2024).
+While this series cross validation focuses on evaluating the generalization capabilities of TSFMs, it is not without drawbacks.
+Roque et al.
+(2025) showed that the final ranking of models depends heavily on which series are selected as test sets.
+With just four cherry-picked test sets, out of the distribution of all available test sets, 46% of the benchmarked models could appear as state-of-the-art purely due to selection bias.
+Notably, deep learning models were found to be more susceptible to this variance than classical baselines.
+Especially when evaluating general-purpose TSFMs, generalization is a core evaluation focus.
+Figure 1 (c) illustrates the current best practice in TSFM evaluation, that is, a combined time and series cross validation (Yu et al., 2016; Salinas et al., 2020).
+While this approach has first adopters, notably Goktas et al.
+(2025); Shchur et al.
+(2025), it is not yet commonly used in TSFM benchmarking.
+2.3 Information Leakage in TSFM Benchmarking: A Recognized Challenge
+As pre-training corpora grow larger and more heterogeneous, ensuring that evaluation data has not been seen during training becomes increasingly difficult.
+The same dynamic has been observed in LLM evaluation: models pre-trained on large internet crawls such as Common Crawl [^1] https://commoncrawl.org/ or The Pile [^2] https://pile.eleuther.ai/ had already seen many benchmark tasks during pre-training (Carlini et al., 2022; Chang et al., 2024), ultimately contributing to what has been described as an LLM “evaluation crisis” (Liao and Xiao, 2023).
+TSFMs face the same structural problems, and the community has begun to respond.
+As shown in the GIFT-Eval ablation studies, information leakage can strongly inflate evaluation performance (Aksu et al., 2024), and several recent benchmarks have developed explicit strategies to limit its impact (Aksu et al., 2024; Goktas et al., 2025; Xu et al., 2025).
+GIFT-Eval pairs an explicit pre-training corpus of approximately 230 billion data points with a dedicated test set of 177 million data points (Aksu et al., 2024).
+TempusBench draws on 48 forecasting tasks from sources currently absent from any known pre-training corpus, and further proposes live evaluation on real-world data (Goktas et al., 2025).
+Fidel-TS takes a similar approach through restricted live APIs (Xu et al., 2025).
+Yet, none of these efforts have been preceded by a systematic analysis of where information leakage actually originates from, nor of how many distinct forms it takes.
+Without that foundation, benchmark design remains reactive rather than principled.
+3 Two Distinct Pathways of Information Leakage in TSFM Evaluation
+The integrity of any benchmark rests on a strict separation between training and evaluation data.
+For TSFMs, this separation is threatened along two distinct dimensions.
+The first is direct: datasets used for pre-training or fine-tuning reappear as test sets, undermining evaluation.
+The second is indirect and specific to time series: temporally overlapping but nominally disjoint series can be statistically correlated, allowing information about the test period to leak into the training signal without any test data point ever being seen directly.
+Evidence for both forms of leakage can already be found in current TSFM evaluations.
+3.1 When Pre-Training Flexibility Comes at the Cost of Benchmark Comparability
+Each TSFM team independently selects which datasets to use for pre-training, fine-tuning, and evaluation.
+Rightly so, since the freedom to train on diverse, large-scale corpora is central to the foundation model paradigm.
+Yet this very flexibility carries a structural side effect that has received little attention: once a dataset has been used for pre-training or fine-tuning by any model, it is no longer admissible as a foundation for benchmarking models against others.
+With each new model trained on a new combination of datasets, the set of valid datasets for benchmarking is reduced.
+A lineage analysis of 22 TSFMs published at leading machine learning venues (TMLR, ICML, NeurIPS, ICLR, WWW) and their workshops through January 2026 (Table 1) makes this concrete.
+For each model, every dataset was categorized into one of three roles: pre-training (used only for initial model training and thus excluded from evaluation), train/test (split temporally for fine-tuning and in-domain evaluation), or zero-shot (reserved exclusively for held-out evaluation on unseen data).
 Figure 2: Lineage of dataset collections (left) used for training and evaluating recent Time Series Foundation Models (right).
@@
 Smaller datasets (used n<10) are shown as "Other".
-Our lineage analysis, the details are described in Supplementary Information Section D and the results are visualized in Figure 2, reveals that there is no consensus between model developers which publicly available datasets should be used for what purpose.
-In fact, each TSFM team uses a different combination of pre-training, train/test, and zero-shot datasets.
-Whereas we find no indication that any individual TSFM suffers from information leakage within its own training and evaluation, the widespread practice of re-purposing datasets introduces substantial challenges for benchmarking across TSFMs.
-Once a dataset has been used for pre-training or fine-tuning by any model, it must be regarded as contaminated and is no longer admissible for TSFM benchmarking.
-Our lineage analysis shows that this is not merely a theoretical concern, but a concrete and pervasive practical issue.
-No dataset has been universally used across all TSFMs for train/test or zero-shot evaluation (see Supplementary Information Section D).
-And only 6% of the analyzed datasets have never been used for pre-training or fine-tuning, making them the only suitable candidates for genuine zero-shot model benchmarking.
-In Supplementary Information Section B, we summarize literature showing that leakage between training and test sets can easily lead to over 50% better test scores.
-Several documented leakage cases suggest that the often unintended overlaps between train and test data are not always identified during the peer-review process Saravanan et al.
-(2024); Montet et al.
-(2025); Li et al.
-(2024b).
-This is not surprising as exactly determining on which datasets a model has been trained requires careful reading of the original paper, its appendix, and sometimes even analyzing published source code.
-Doing this for a dozen or more models when performing a benchmarking study requires substantial effort from the authors and reviewers as well.
-As more TSFMs are published, authors increasingly face the challenge of navigating complicated dataset relationships.
-For example, the authors of the Tirex model discovered that a dataset called "Solar" appears in parts in the Gift-Eval test data Aksu et al.
-(2024) as well as in the Chronos pre-training data, which is used by several models including Tirex, Toto, and Sundial Ansari et al.
-(2024a); Auer et al.
-(2025); Liu et al.
-(2025); Cohen et al.
-(2025).
-As a precaution, Tirex and Toto excluded the dataset, though Tirex explicitly notes the difficulty of verifying whether the data is actually identical.
-Through our lineage analysis, we found that in the Chronos pre-training data, the "Solar" dataset includes solar power data from all US states at two frequencies: 5T (5 minutes) and 1H (hourly), where the 5T data was sourced directly from the original data provider [^5] https://www.nlr.gov/grid/solar-power-data.
-Gift-Eval, however, sourced only the Alabama subset as a test set, aggregated to 10T, 1H, 1D (daily), and 1W (weekly), through the LSTNet dataset repository [^6] https://github.com/laiguokun/multivariate-time-series-data/tree/master/solar-energy.
+The lineage map (Figure 2; full dataset-level analysis in Section 6) confirms this: across 401 datasets identified, each TSFM team assembled a distinct combination of pre-training, train/test, and zero-shot datasets, and no single dataset has been universally reserved for evaluation.
+Only 6% of these 401 datasets have never appeared in any model’s pre-training or fine-tuning corpus.
+These are the only candidates that could, in principle, support genuine zero-shot comparisons across the full set of published TSFMs.
+The pervasiveness of this problem is best illustrated through concrete examples.
+The Australian Electricity Demand dataset from the Monash collection (Godahewa et al., 2021) has been used for pre-training (e.g., Lag-Llama, Timer), train/test evaluation (e.g., Moirai), and zero-shot forecasting (e.g., Chronos, TimesFM), making it practically useless for any cross-model comparison.
+Similarly, the Informer collection (Zhou et al., 2021), widely treated as a zero-shot benchmark, contains individual series such as ETTh1, ETTh2, and ETTm1 that have also been used for pre-training (e.g., in Lag-Llama) or train/test evaluation (e.g., in UniTime).
+These overlaps are often difficult to detect, because datasets are routinely remixed, renamed, and redistributed across repositories.
+For instance, the dataset “Elecdemand” from the Monash Repository is a scaled subset ($1/1000$) of “Australian Electricity Demand”.
+The dataset “ElectricityLoadDiagrams20112014” appears under the name “Electricity” in both the Autoformer and Monash collections (Godahewa et al., 2021; Wu et al., 2022) and as “ECL” in the Informer collection (Zhou et al., 2021).
+Even datasets sharing the same name may contain different underlying data: the car ride shares dataset is included from the same year 2015 in both the Monash and GluonTS collections.
+Although both repositories cite the same GitHub source [^3] https://github.com/fivethirtyeight/uber-tlc-foil-response, a close examination reveals that they contain different time series of distinct For-Hire Vehicles (Uber in GluonTS vs.
+Lyft in Monash).
+Consequently, it is necessary to always compare the underlying data itself, keeping in mind that it could have been transformed in various ways (scaling, handling missing values, renaming, resampling, etc.).
+A particularly telling example of how dataset provenance can become convoluted through multiple transformations and repositories involves the “Solar” dataset.
+The Chronos pre-training corpus includes solar power data from all US states at two frequencies: 5T (5 minutes) and 1H (hourly), sourced directly from the original data provider [^4] https://www.nlr.gov/grid/solar-power-data.
+GIFT-Eval, however, includes only the Alabama subset as a test set, aggregated to 10T, 1H, 1D (daily), and 1W (weekly), sourced through the LSTNet dataset repository [^5] https://github.com/laiguokun/multivariate-time-series-data/tree/master/solar-energy.
 LSTNet also sourced the data from the original provider but, as noted, included only Alabama rather than all states.
 In addition, the Alabama solar subset appears in the Monash Repository [^6] https://forecastingdata.org/ at frequencies 10T and 1W.
-This modified version of the source data was used as pretraining data by the models TinyTimeMixers and Lag-Llama Ekambaram et al.
-(2024); Rasul et al.
-(2024).
-While TinyTimeMixers and Lag-Llama were published before Gift-Eval, the authors of Sundial validate against Gift-Eval and state they excluded any overlaps without explicitly listing which datasets they excluded.
-This example illustrates how dataset provenance can become convoluted through multiple transformations and repositories, making it difficult for researchers to identify overlaps.
-While in the previous example the dataset name was constantly "Solar", relying on dataset names to avoid train-test overlaps is not sufficient.
-The identification of suitable datasets is complicated by dataset remixing and renaming.
-For example, the dataset "Elecdemand" from the Monash Repository is a scaled subset ($1/1000$) of the “Australian Electricity Demand” dataset.
-Similarly, the dataset "ElectricityLoadDiagrams20112014" appears under various names, such as "Electricity" in the Autoformer and Monash collections (Godahewa et al., 2021; Wu et al., 2022) and "ECL" in the Informer (Zhou et al., 2021) collection.
-Dataset collections sometimes also contain different observations of the the same variables.
-For instance, the car ride shares dataset is included with same frequency and from the same year 2015 in both the Monash and GluonTS collections.
-Although both repositories cite the same GitHub source [^8] https://github.com/fivethirtyeight/uber-tlc-foil-response, a close examination reveals that they contain different time series of distinct For-Hire Vehicles (Uber in GluonTS vs Lyft in Monash).
-Consequently it is necessary to always compare the underlying data itself, keeping in mind it could have been transformed in various ways (scaling, handling missing values, renaming, resampling, etc.).
-3.2 Indirect Information Leakage: Temporal Overlap of Correlated Train and Test Series
-Direct overlap in data points between training and test sets is not the only cause for information leakage.
-In the context of time series data, a more subtle case of information leakage can happen due to indirect dependencies between data points in the training and test sets (Kapoor and Narayanan, 2023).
-Often, these indirect dependencies stem from confounding bias, that is, an exogenous factor influencing observations from both datasets.
-For example, global events or crises like COVID-19 or the European energy crisis, caused by Russia’s invasion of Ukraine, can influence a large number of time series from different domains.
-However, not only global phenomena but also local factors can introduce such dependencies.
-Weather conditions at a specific location, for instance, directly influence solar energy production data, creating a causal relationship between meteorological measurements and power generation time series.
-Similarly, local traffic patterns may affect both air quality measurements and delivery time predictions within the same geographic area.
-A key strength of TSFMs lies in their ability to learn subtle patterns in a flexible, data-driven manner and to apply these patterns to forecast future values of previously unseen time series.
-At large scale, however, TSFMs may inadvertently memorize patterns observed in one or more training time series and transfer them to series in the test set from the same time period.
-As long as training and test series are correctly ordered in time and do not overlap, this behavior does not constitute a serious issue.
-However, if training and test series overlap temporally and have a casual relationship (confounding or direct causation), such pattern transfer can result in information leakage.
-In the example illustrated in Figure 3, a highly flexible TSFM may learn the characteristic shape of the stock market decline caused by the global COVID-19 pandemic from the time series of the DAX index and reproduce it when tasked with forecasting the S&P 500 index over the same period.
+This modified version was used as pretraining data by TinyTimeMixers and Lag-Llama (Ekambaram et al., 2024; Rasul et al., 2024).
+While TinyTimeMixers and Lag-Llama were published before GIFT-Eval, the authors of Sundial validate against GIFT-Eval and state they excluded any overlaps without explicitly listing which datasets they excluded (Aksu et al., 2024; Liu et al., 2025).
+The authors of TiRex and Toto discovered this entanglement and excluded the dataset as a precaution, though TiRex explicitly notes the difficulty of verifying whether the data across repositories is actually identical (Auer et al., 2025; Cohen et al., 2025).
+The consequences of undetected contamination can be severe.
+As summarized in Supplementary Information Section A, documented cases of leakage between training and test sets have led to over 50% better test scores.
+Several such cases suggest that overlaps between train and test data are not always identified during the peer-review process (Saravanan et al., 2024; Montet et al., 2025; Li et al., 2024b).
+This is not surprising: exactly determining on which datasets a model has been trained requires careful reading of the original paper, its appendix, and sometimes even analyzing published source code.
+During the lineage analysis, a potential overlap surfaced between the pre-training and test set in TimesFM (Das et al., 2024), as the traffic hourly dataset appears to have been used during pretraining while also being included in the reported Monash evaluation benchmark.
+This anecdotal example underscores how difficult it is to maintain a reliable overview of datasets; a challenge that only intensifies as the number of published TSFMs grows.
+3.2 When Temporal Overlap Turns Correlation Into Leakage
+Time series carry a special structural property: temporal correlation between nominally independent series.
+This correlation often has causal roots, yet TSFMs do not learn causal relationships; they learn from observed statistical patterns alone.
+When an exogenous shock affects multiple domains simultaneously — a pandemic, a financial crisis, a geopolitical shift — it imprints a recognizable signature across many series at once.
+The same effect operates at a local scale: weather conditions at a specific location directly shape solar energy production (Aryandoust et al., 2022), and local traffic patterns simultaneously influence air quality measurements and delivery time predictions.
+A model trained on any of these series during the affected period has, in effect, learned something about all of them, even those it never encountered directly.
+While spurious correlation without any shared cause could in principle produce a similar effect, it may also lead to worse predictions, as the model transfers patterns that bear no true relationship to the test series.
+But the systematic correlation arising from genuine common drivers is the more prevalent and serious concern.
 Figure 3: Indirect Information Leakage - Temporal Overlap of Correlated Train Series A and Test Series B.
-This temporal overlap of correlated train-test samples introduces a problematic dependency between the training and test data, violating the assumption of independence that is crucial for robust evaluation.
-As a result, the model gains access to information during training that would be inaccessible in a real-world scenario at the time of prediction, leading to inflated performance metrics.
-Therefore the explicit time period of a pattern and its actually encoded information gain importance for globally trained models such as TSFMs.
-This issue is not merely theoretical, but has been demonstrated empirically.
-Rodrigo and Ortiz (2024) conducted an experiment isolating the influence of temporal overlap information leakage.
-They build a setup with highly correlated time series, such as metro, bus, road, and train usage in Madrid for the year 2024.
-They trained two models in a global manner with all series.
-One model was trained with a common train-test split date for all series.
-For the other model only the metro time-series was cutted at the train-test split date, keeping the other series fully in the train set.
-After training, they only predicted the test period of the metro series, which was never part of the training dataset.
-The second model with full access to the other correlated series during training achieved a Mean Absolute Error (MAE) of 248,24 compared to 439,14 to the first model with a common train-test split date across all series.
-This corresponds to a performance improvement of approximately 43.47% attributable solely to this indirect form of leakage, confirming that access to correlated and temporally overlapping external information significantly distorts evaluation results even without direct access to the target series.
+The evaluation implication follows directly.
+A key strength of TSFMs lies in their ability to learn subtle patterns in a data-driven manner and transfer them to previously unseen series.
+At large scale, however, this same capability becomes a liability for evaluation.
+When correlated training and test series overlap temporally, a sufficiently flexible TSFM can transfer learned statistical structure from the former to the latter.
+Crucially, this does not reflect genuine forecasting skill: the model is not learning the generative dynamics of the test series, but exploiting correlation produced by a shared external driver.
+As illustrated in Figure 3, a TSFM trained on the DAX index through the COVID-19 crash absorbs the temporal signature of that event.
+When subsequently asked to forecast the S&P 500 over the same window, it can reproduce that shape — not by understanding US market dynamics, but because both indices were driven by the same global shock and are therefore highly correlated during that period.
+The key distinction from classical benchmarking is therefore not only which series appeared in training, but when those series were observed: contemporaneous observation of correlated series embeds information about the test period even when no test data point was ever directly seen.
+This temporal overlap violates the independence assumption fundamental to unbiased evaluation, inflating performance metrics in ways that do not reflect real-world forecasting ability.
+This concern is not merely theoretical.
+Rodrigo and Ortiz (2024) provided a clean empirical demonstration using highly correlated public transport series from Madrid: metro, bus, road, and train ridership throughout 2024.
+The metro series was withheld entirely from training and served as the zero-shot forecasting target.
+For all remaining series, two training configurations were compared: in the first, a common train-test split date was applied, so that none of the correlated series extended into the metro test period; in the second, all non-metro series were kept in training up to the end of the evaluation window, overlapping temporally with the metro test period.
+Both models were then evaluated exclusively on the metro test period.
+The model whose correlated training series overlapped with the test period achieved a Mean Absolute Error (MAE) of 248.24 compared to 439.14 for the properly split model, a performance improvement of approximately 43% attributable solely to this indirect form of leakage.
+No metro data point was ever seen during training; the advantage came entirely from the shared temporal dynamics of the correlated series.
+The Madrid experiment used a near-ideal setting for leakage: virtually all training series were highly correlated with the test target.
+The more pressing question for TSFM evaluation is whether temporal leakage remains detectable when the leaked signal is diluted among a large, heterogeneous training corpus.
+A controlled replication around the COVID-19 stock market crash of early 2020 suggests the answer is yes.
+To isolate the effect cleanly, a Time-Series Transformer (TST) (Zerveas et al., 2020) was trained from scratch, rather than relying on an existing TSFM that could already carry this form of leakage.
+The architecture shares multi-variate characteristics with recent models such as MOMENT or Chronos-2 (Goswami et al., 2024; Ansari et al., 2025) while remaining small enough for fast training saturation.
+Two almost identical training sets were constructed [^7] Details on the general setup, used datasets, scaling, hyperparameter-tuning and results can be found in Supplementary Information Section B.: both contained the same non-financial series (NN5 Daily and Tourism Monthly from the Monash collection) plus seven major stock indices (Bovespa, CAC40, DAX, DowJones, FTSE 100, Nikkei 225, TSX) [^8] All stock index data was retrieved from the website investing.com, differing only in the temporal range of the stock data.
+The first set cut the stock series at December 2019, before the pandemic; the second extended them through December 2020, encompassing the crash and partial recovery.
+In both cases, the S&P 500 was withheld entirely as the zero-shot test target.
+Stock prices were chosen deliberately: they are notoriously hard to model beyond short horizons (Goyal et al., 2024), ensuring that any systematic performance gap reflects leaked external information rather than improved modeling of intrinsic dynamics.
 Figure 4: Zero-shot forecasting for S&P 500 stock index during the Covid shock in the year 2020.
@@
 In both cases S&P 500 is not part of the training data.
-While the study by Rodrigo and Ortiz (2024) used time series where the complete training data is highly correlated and temporally overlapping with the complete test data, we replicate an TSFM training process in a more realistic scenario.
-We incorporate diverse data where only a small percentage of observations is affected by indirect leakage.
-To demonstrate the impact of temporal train-test leakage on model performance under these conditions, we conducted an experiment using a trained-from-scratch Time-Series Transformer (TST) (Zerveas et al., 2020).
-As we have seen in Section 3.1, any other pre-trained TSFM could be already affected by this form of information leakage.
-The encoder-only model shares multi-variate characteristics with recent models such as MOMENT or Chronos-2 (Goswami et al., 2024; Ansari et al., 2025) but is significantly smaller, enabling faster training saturation.
-We trained the model using 10 different random seeds for two different training sets [^9] Details on the general setup, used datasets, scaling, hyperparameter-tuning and results can be found in Supplementary Information Section C: one with a temporal overlap to the test set during the COVID-19 pandemic and one without such overlap.
-The first training set comprised the NN5 Daily and Tourism Monthly (Monash collection), and the American (Dow Jones), Brazilian (Bovespa), British (FTSE 100), Canadian (TSX), French (CAC 40), German (DAX) and Japanese (Nikkei 225) stock indices from 2015 to the end of 2019.
-The only difference in the second training set was that the stock index data spanned from 2016 to the end of 2020, comprising observations affected by the COVID-19 outbreak.
-No data points were shared directly between the training and test sets.
-Table 2: Per-seed S&P 500 MAE breakdown for indirect leak from other countries stock indices vs no-leak experiments.
-The percentage difference shows how much better the model with leakage data performs than the non-leakage model.
-For each seed, the better MAE in displayed in bold.
-Seed MAE Before Split MAE After Split
-No Leak Indirect Leak % Diff No Leak Indirect Leak % Diff
-410 230.53 227.72 1.22% 157.85 199.33 -26.28%
-1425 219.68 178.09 18.93% 120.77 82.53 31.67%
-1680 193.52 187.19 3.27% 80.74 81.22 -0.60%
-1825 247.76 156.12 36.99% 110.47 70.36 36.31%
-2287 231.19 206.28 10.78% 102.42 100.54 1.84%
-3658 219.23 156.91 28.43% 117.25 84.62 27.84%
-4013 224.75 193.16 14.06% 114.19 97.98 14.20%
-4507 206.45 213.70 -3.51% 93.10 84.01 9.77%
-8936 189.40 199.38 -5.27% 103.59 153.67 -48.35%
-9675 206.30 182.99 11.30% 82.35 132.44 -60.82%
-Mean 216.88 190.15 11.62% 108.27 108.67 -1.44%
-Std Dev 17.15 21.88 21.10 38.83
-We selected stock prices as our target variable as they are specifically hard to model appropriately in a long-term valid manner Goyal et al.
-(2024), suggesting that any performance gaps must come from external leaked information.
-For both models, we used the S&P 500 index as a zero-shot test set, which was not part of both datasets.
-Additionally, we chose a time-frame where a common cause had influence on many different time series and chose therefore the COVID-19 outbreak.
-In Figure 3, time series 1 shows the DAX index and time series 2 represents the S&P 500 index.
-Note how the COVID-19 outbreak is visible in both series.
-Figure 4 displays the mean and standard deviation of the predictions for both models across the ten runs, and Table 2 shows the specific values for each seed.
-For almost all seeds, the model trained on data with temporal overlap to the test set predicted the initial market crash following the outbreak significantly better than the non-leakage model.
-In total, the temporal overlap model performed better eight out of ten times.
-Up to the first of May, the leakage model achieved an average MAE of about 190 compared to about 217 for the non-leakage model.
-For the remaining part, both models had a very similar average MAE, with the non-leakage model being slightly better (108.3 compared to 108.7 for the non-leakage model.
-When looking at the specific runs, shown in Table 2, the performance difference between the two models is up to about 37% in favor of the leakage model during the shock.
-Despite the temporal overlap representing only a small fraction of the overall training data, specifically seven time series out of 484 (or 1,600 out of 205,859 observations in total) (Table 6), this performance gap underscores the significant impact the small temporal overlap of correlated series can have on the predictive performance of a model.
-4 TSFM Benchmark Requirements
+Figure 4 displays the mean and standard deviation of the predictions across ten random seeds for both models, with per-seed results provided in Supplementary Information Table 4.
+The results are unambiguous: during the crash period, the model trained on data with temporal overlap predicted the initial market decline significantly better than the clean model, outperforming it in eight out of ten runs with individual improvements reaching up to approximately 37%.
+Up to the first of May, the temporal overlap model achieved an average MAE of about 190 compared to about 217 for the non-leakage model.
+For the remaining part, both models converged to very similar average performance (108.3 vs.
+108.7), confirming that the advantage was specific to the period of shared causal influence rather than reflecting a general training benefit.
+This performance gap is all the more striking given that the leaked signal comprised just seven time series out of 484, or roughly 1,600 out of over 205,000 total training observations (Table 3).
+If a small transformer trained from scratch on fewer than 500 series can exploit temporal correlation for a measurable advantage, production-scale TSFMs trained on millions of series are likely more susceptible, not less, given their greater capacity to memorize subtle patterns and the higher probability that their pre-training corpora contain series that are temporally overlapped and correlated with any given test set.
+Both the Madrid transport experiment (Rodrigo and Ortiz, 2024) and the COVID stock market experiment involve indirect leakage between series of the same frequency.
+Yet a further, largely undiscussed complication concerns the relationship between different frequency representations of the same underlying process.
+The community has not yet reached consensus on this matter: some treat different sampled frequencies as practically independent (Auer et al., 2025; Shchur et al., 2025; Das et al., 2024), while others leverage similar patterns across frequencies architecturally and consider them dependent (Liu et al., 2024b; Graf et al., 2025).
+We argue that frequency variants of the same underlying process should be treated as dependent, for two reasons.
+First, time series sampled at different frequencies from the same source retain very high correlation when mapped to common timestamps.
+Second, if a TSFM is capable of capturing the “bigger picture” of a time series (and this is precisely the promise of foundation models), it should recognize the same trends and seasonal patterns regardless of temporal resolution.
+Training on one frequency and testing on another from the same ground truth should therefore be considered a form of indirect leakage.
+4 Two Requirements for Information-Leakage-Free TSFM Evaluation
 The essence of time series forecasting is to predict the future.
@@
 The only pitfall to avoid is adhering to the correct ordering of training and test observations.
-However, the paradigm shift behind TSFMs introduces new possibilities for information leakage.
+Figure 5: Information leakage risks and challenges arising from the advent of TSFMs, and the requirements needed to address them.
+However, the paradigm shift behind TSFMs shatters this assumption.
 Unlike local models, TSFMs learn patterns globally.
 Consequently, strict temporal ordering within a single time series is no longer sufficient to guarantee a fair evaluation.
-To rigorously assess these models, we must address leakage across two distinct dimensions: the sample dimension and the temporal dimension (see Figure 5).
-Figure 5: Information leakage risks and challenges arising from the advent of TSFMs, and the requirements needed to address them.
-Flexibility in Pre-Training Data (Training Dimension) The current lack of transparency surrounding TSFM pre-training corpora creates a substantial risk of information leakage as shown in Section 3.1.
-This growing confusion is evident in newer benchmarks and models that report results contaminated by data leakage—whether unintentional or intentional Aksu et al.
-(2024); Auer et al.
-(2025); Li et al.
-(2024b).
-Attempting to mitigate this risk by restricting pre-training data is both counterproductive and ineffective: such restrictions artificially limit the potential of foundation models and, given the opacity of large-scale pre-training corpora, our lineage analysis shows that it is almost impossible to rule out train-test sample overlaps.
-We therefore argue for full flexibility in pre-training data.
-A robust evaluation framework should be agnostic to pre-training data choices, allowing the use of any existing data to fully exploit the predictive capabilities of TSFMs.
-This perspective shifts the responsibility for preventing leakage from the TSFM developer to the benchmark designer, thereby necessitating evaluation datasets that are genuinely novel and capable of assessing generalization rather than memorization.
-Strict Global Post-Training Test Periods (Inference Dimension) Indirect temporal dependencies between time series, originating from direct or indirect causal relationships, create another more subtle risk for information leakage.
-Preventing this risk by enforcing global cut-off dates for pre-training data to separate training from test data is not a viable solution to this challenge.
-First, it would require all researchers to tailor their training protocol to exactly this date and second, it prevents new time series patterns of current time periods after the cut-off from being trained into the models.
-We therefore argue for strict global post-training test periods.
-These test periods need to be flexibly determined based on the global last pre-training time point of all evaluated models, instead of forcing a pre-training cut-off.
-For a robust evaluation framework a strict temporal barrier must be enforced across the entire universe of data: if the last pre-training timestamp across all evaluated models is $t$, only observations from $t+1$ onwards can be present in the test set of a benchmark, ensuring that TSFMs rely solely on past dynamics.
-Addressing these two dimensions (unrestricted training data and strict temporal inference separation) is non-trivial and often overlooked in current evaluation standards.
-We ask the community to develop ideas for new benchmark methodologies which are grounded in these two requirements.
-Possible directions could be constantly newly sourced test data, synthetic test data, or fast-paced competitions on unreleased and unbiased test data, just to name a few.
+To rigorously assess these models, we must address leakage across two distinct dimensions: through what was in the training data (the sample dimension) and through when the training data was observed (the temporal dimension) (see Figure 5).
+Requirement 1: Full flexibility in pre-training data.
+Attempting to prevent contamination by restricting what data TSFMs may train on is both counterproductive and unenforceable.
+Restrictions artificially limit the potential of foundation models, and the lineage analysis presented here demonstrates that verifying the absence of train-test overlap across opaque, large-scale pre-training corpora is practically impossible.
+The responsibility for preventing leakage must therefore shift from model developers to benchmark designers.
+A robust evaluation framework should be agnostic to pre-training choices, allowing any available data to be used for training.
+This means that test sets must be genuinely novel: data that could not plausibly have appeared in any pre-training corpus.
+The benchmark, not the model, must guarantee separation.
+Requirement 2: Strict global post-training test periods.
+Indirect temporal leakage cannot be prevented by controlling which series appear in training and test sets; it arises from correlation between any temporally overlapping series.
+A hard temporal barrier is therefore needed: if the latest pre-training timestamp across all evaluated models is $t$, then every observation in the test set must come from $t+1$ or later.
+This barrier must be determined globally, based on the actual training cutoff of the latest model in the comparison, rather than imposed as a fixed date.
+Fixing a universal cutoff would require all researchers to align their training protocols to the same date and would prevent models from learning current patterns, both of which are undesirable.
+A flexible, model-aware temporal barrier $t_{now}$ avoids these drawbacks while ensuring that no test observation $t_{now}+1$ could have been influenced by any training observation through shared causal drivers.
+Together, these two requirements form a coherent basement for new frameworks.
+Requirement 1 addresses the sample dimension of leakage by making test data provably novel.
+Requirement 2 addresses the temporal dimension by ensuring strict temporal precedence of all training over all test data.
+Neither requirement alone is sufficient; both are necessary.
+Realizing both demands not incremental adjustments to existing evaluation practices, but a structural rethinking of how TSFM benchmarks are designed and maintained.
 5 Conclusion
-Remarkably, TSFMs appear to follow a development trajectory similar to that of LLMs in terms of both data and model size.
-Scaling laws seem to apply to TSFMs, and there are notable similarities in the evaluation challenges they face, particularly concerning potential information leakage.
-We highlight key evaluation challenges that arise from the emergence of TSFMs.
-By investigating the data lineage of 22 prominent TSFMs and reflecting on common TSFM training and evaluation strategies, we identified two significant sources of information leakage: direct information leakage by train-test sample overlaps through multi-purpose use of datasets and indirect information leakage by temporal overlap of correlated train and test series.
-Current benchmark approaches from time series forecasting or machine learning do not mitigate all information leakage risks.
-Based on our investigations and existing methodologies, we propose two key requirements for information leakage free evaluations in the age of TSFMs: flexibility in pre-training data and strict global post-training test periods.
-Beyond the development of new benchmark methodologies, we identify several directions for future research .
-To date, the effect size of test-set contamination—when a test dataset is present in the pre-training corpus—has only been shown to be significant in isolated cases.
+We call on the community to develop new benchmark methodologies grounded in both requirements proposed here: full flexibility in pre-training data and strict global post-training test periods.
+Possible directions include constantly newly sourced test data, synthetic test data, or fast-paced competitions on unreleased and unbiased test data.
+While these two requirements define the conditions for trustworthy evaluation, several critical questions remain open.
+To date, the effect size of test-set contamination, when a test dataset is present in the pre-training corpus, has only been shown to be significant in isolated cases.
 Likewise, the magnitude of temporal leakage arising from correlated and overlapping time series remains largely unknown in real TSFM settings and is likely to vary across domains, types of global events, and other contextual factors.
@@
 6 Data Availability
-The resulting table of the TSFM data-lineage analysis is supplied in the Supplementary Information Section D.
-For the global pattern experiment in Section 3.2, the used data is available through the source code via zenodo: https://doi.org/10.5281/zenodo.18635478
+The resulting table of the TSFM data-lineage analysis is supplied in the Github Repository under this link.
+For the indirect information leakage experiment in Section 3.2, the used data is available through the source code as well via Github: https://github.com/DAG-UPB/rethinking-tsfm-evaluation
 7 Code Availability
-The code to reproduce the experiment in Section 3.2 is provided via zenodo: https://doi.org/10.5281/zenodo.18635478
-Appendix A Time Series Foundation Model Analysis
-A.1 Multi-Purpose Use of Train and Test Data
-This section showcases the varying training and testing dataset strategies of selected TSFMs.
-Table 3 reveals that there is virtually no standardized approach among TSFMs, with each model employing different combinations of datasets for train/test (or in-domain) datasets.
-While some datasets like Electricity (H) and Traffic Hourly (H) appear across multiple models, the coverage remains fragmented.
-Table 4 further illustrates this inconsistency in zero-shot evaluation practices.
-Even among datasets with five or more overlaps, the specific models testing on each dataset vary considerably.
-For instance, ETT2 (15T) is evaluated by Chronos, Moirai, and Lag-Llama, while Electricity (H) appears in Moirai, Moirai-MoE, and LLMTime.
-Another notable case for the multi-purpose use of datasets is the subset Australian Electricity Demand from the Monash dataset collection (Godahewa et al., 2021), which is used for pre-training (e.g., Lag-Llama, Timer), train/test evaluation (e.g., Moirai), and zero-shot forecasting (e.g., Chronos, TimesFM).
-This multi-purpose use makes the dataset practically useless for benchmarking between models (e.g., it cannot be used to compare Chronos to Lag-Llama).
-To understand the problem of overlaps better, we provide Figure 6.
-Also an interesting case is the Informer collection (Zhou et al., 2021), which is primarily used for zero-shot evaluations.
-However, some individual time series of the Informer collection, such as ETTh1, ETTh2, and ETTm1, have also been used for pre-training (e.g., in LagLlama) or train/test evaluation (e.g., in UniTime).
-This means that LagLlama or UniTime cannot be benchmarked against any other model using the Informer collection.
-Taken together with investigations of Roque et al.
-(2025) that the selection of evaluation datasets can highly influence the final evaluation results, this strengthens the need for a more uniform benchmark approach.
-Figure 6: Test set contaminations when comparing TSFM A with TSFM B with overlapping pre-training and test sets on the test set of TSFM A.
-Table 3: Comparison of train/test evaluations of selected different models ordered by the total number of overlaps. For full results see Section D.
-Dataset Chronos Moirai MOMENT Time-LMM
-Electricity (H) T/T - T/T T/T
-Traffic Hourly (H) - T/T T/T T/T
-ETT1 (H) - - T/T T/T
-Weather (Australia) (D) - T/T - T/T
-M4 Monthly (M) T/T T/T - -
-M4 Hourly (H) T/T T/T - -
-M4 Daily (D) T/T T/T - -
-M4 Weekly (W-SUN) T/T T/T - -
-ETT1 (15T) - - T/T T/T
-ETT2 (15T) - - T/T T/T
-ETT2 (H) - - T/T T/T
-CDC Fluview ILINet (W) - - T/T T/T
-Table 4: Comparison of zero-shot evaluations of selected different models ordered by the total number of overlaps.
-Only datasets with 5 or more overlaps are shown.
-Models which have not included any of the shown datasets into their train/test evaluation are not displayed.
-For full results see Section D.
-Dataset Chronos Moirai Moirai-MoE Lag-Llama LLMTime
-Jena Weather (10T) - ZS ZS - ZS
-ETT2 (15T) ZS ZS - ZS -
-Electricity (H) - ZS ZS - ZS
-ETT2 (H) ZS ZS - - -
-ETT1 (H) ZS ZS - - -
-ETT1 (15T) ZS ZS - - -
-Traffic Hourly (H) ZS - - - ZS
-Appendix B Leakage Investigations
-B.1 Effects of Overlaps in Train-Test Samples
-This section collects different examples where the train-test sample overlaps unintentionally led to information leakage in recent benchmark papers.
-In one case, the benchmark creators, due to the TSFMs’ incomprehensible datasets, unintentionally included three evaluation datasets as test sets that had already been used in the pretraining of TimesFM, UniTS, and TTM (Li et al., 2024b).
-Based on our analysis of the benchmark results, this leads to an advantage of 47% - 184% lower mean squared error (MSE) rate compared to best models not pre-trained on the leaking datasets.
+The code to reproduce the indirect information leakage experiment in Section B is provided via Github: https://github.com/DAG-UPB/rethinking-tsfm-evaluation
+Appendix A Leakage Investigations
+A.1 Cases of Overlaps in Train-Test Samples
+This section summarizes individual cases where train-test sample overlaps unintentionally led to information leakage in recent benchmarking studies.
+Through our lineage analysis, we could easily compare the used benchmark datasets in the study against the TSFM training datasets (for full lineage table see Section 6).
+In one case, the benchmark creators, due to the TSFMs’ incomprehensible datasets, unintentionally included three evaluation datasets as test sets that had already been used for the pretraining of TimesFM, UniTS, and TTM (Li et al., 2024b).
+Based on our analysis of the benchmark results, this lead to an 47% - 184% lower mean squared error (MSE) rate compared to best models not pre-trained on the leaking datasets.
 The advantage of the best TSFM on non-leaked datasets is only between 0.3% and 14%.
-The performance benefit through information leakage appears to align with findings from the Moirai leakage example (see Section B.2).
-In another example a peer-reviewed paper benchmarked TSFM models on the Electricity dataset which is included in every tested TSFM pretraining data Saravanan et al.
+This performance benefit through information leakage appears to align with findings from the Moirai leakage example (see Section A.2).
+In another example a peer-reviewed paper benchmarked TSFM models on the Electricity dataset, which has been used for pretraining of every of tested TSFMs Saravanan et al.
 (2024).
@@
 (2025).
-This highlights the increasing difficulties for benchmarking TSFMs.
-B.2 Intended Information-Leakage Experiment for Moirai Model
-Table 5: Information leakage in TSFM comparing Moirai with (MoiLeak) and without (Moi) information leakage during training phase for different model sizes.
+Taken together, these illustrative case highlight the increasing difficulties for finding suitable benchmarking data for TSFMs.
+A.2 Intended Information-Leakage Experiment for Moirai Model
+Table 2: Information leakage in TSFM comparing Moirai with (Moirai Leakage) and without (Moirai) information leakage during training phase for different model sizes.
 Metric is MAPE.
@@
 L 0.78 0.82 0.71 0.60 2.14 2.24
-The empirical analysis using the Moirai TSFM by Aksu et al.
-(2024) demonstrates the effects of information leakage between training and test sets on the predictive performance of TSFMs.
-The authors prepared a new pre-training and held-out evaluation dataset without any overlap and trained the Moirai architecture on it.
-A different Moirai model from a previous publication was taken which pre-training dataset contained 0.1% data of the newly defined held-out evaluation dataset leading to a deliberate small portion of information-leakage.
-The investigated data includes nine datasets tested on three different forecast horizons (short, medium, long).
-Not for all horizons data was available, so we filtered the results on 4 datasets where the results are available for all horizons to reasonably compare the metrics and model sizes, as shown in Figure 7.
-On short horizons, both models, leakage-free and leakage model, perform better with increasing model size while Moirai Leakage has on average 7.6% points lower MAPE score.
-At medium horizons, the leakage-free MoiraiL generally underperforms compared to other model sizes, exhibiting an average performance deficit of 32% points compared to the Moirai LeakageL model.
-On long horizons, all models perform significantly worse.
-However, Moirai Leakage achieves an average MAPE score that is 29% points lower than the leakage-free Moirai.
-Across all horizons, the biggest difference is between MoiraiL and Moirai LeakageL.
-These results vividly demonstrate the significant impact that train-test sample overlap has on the performance of TSFMs. Larger model sizes especially seem prone to this.
+Aksu et al.
+(2024) performed an empirical analysis using the Moirai TSFM to demonstrate the effects of information leakage between training and test sets.
+The authors prepared a new pre-training and held-out evaluation dataset without any overlaps and trained the Moirai architecture on it.
+In addition, they took an already pretrained Moirai model from a previous publication, whose pre-training dataset contained 0.1% of the newly defined held-out evaluation data, resulting in a deliberately small amount of information leakage.
+The leaked data includes nine datasets and three different forecast horizons (short, medium, and long), though not all combinations of datasets and horizons are present.
+Figure 6 summarizes the MAPE of both models.
+On short horizons, Moirai Leakage has on average an 8 percentage points lower MAPE score across all model sizes.
+At medium horizons, Moirai Leakage exhibits an average accuracy advantage of 15 percentage points.
+And on long horizons, Moirai Leakage achieves an average MAPE score that is 29 percentage points lower than the leakage-free Moirai.
+Evaluation across model sizes reveals that larger models benefit more substantially from data leakage.
+The small model (S) shows a modest average improvement of 4 percentage points, while the base model (B) demonstrates a more pronounced advantage of 13 percentage points.
+Most notably, the large model (L) exhibits the strongest leakage effect with an average MAPE reduction of 34 percentage points across all forecasting horizons, indicating that model capacity amplifies the impact of training-test data contamination.
+These results vividly demonstrate the significant impact that train-test sample overlap can have on the performance of TSFMs and show that larger models are especially prone to this type of leakage.
 Figure 6: Comparing the model Moirai once with information leakage (Moirai Leakage), where the model was exposed to the test data during pre-training and the same model trained without information leakage (Moirai).
@@
 (2024).
-B.3 Potential Overlaps of Correlated Train and Test Series in Dataset Collections
+A.3 Potential Overlaps of Correlated Train and Test Series in Dataset Collections
 It may be theoretically possible that existing model evaluation results have already been influenced by overlaps of correlated train and test series.
@@
 As an example, the proposed Australian Electricity Demand test time frame (April 2015) lays in the same period as the training period of the "Weather (Australia)" training set (1900-2021).
-The Moirai and Moirai-MoE models use both datasets for train/test forecasts, meaning if their are any global patterns in the weather data which could influence the Australian Electricity Demand data, this would be a kind of information leakage, at least on theoretical basis.
-Similarly, the Oikolab Weather dataset ranges from 2010 January to 2021 May, which is weather data located in Melbourne.
+The Moirai and Moirai-MoE models use both datasets for train/test forecasts, meaning if there are any global patterns in the weather data which could influence the Australian Electricity Demand data, this would be a kind of information leakage, at least on theoretical basis.
+Similarly, the Oikolab Weather dataset ranges from January 2010 to May 2021, which is weather data located in Melbourne.
 The test set of the dataset Pedestrian Count (also Melbourne) is located at the end of April 2020.
@@
 No data cleaning was done.
-Appendix C Temporal Overlap of Correlated Series Experiment
-This section provides implementation details of the experiment conducted to show the influence of the temporal overlap of correlated train and test series in a TSFM setting.
-C.1 Data visualizations and statistics
+Appendix B Temporal Overlap of Correlated Series Experiment
+B.1 Data visualizations and statistics
 Figure 8 shows two example time series from major stock indices, the DAX and S&P 500.
@@
 This visual comparison highlights how global financial markets often move in tandem, reflecting shared responses to worldwide economic events.
-Figure 10 illustrates the relationship between training data and test data that overlap temporally.
-It is important to note that while the training and test periods overlap in time, they do not share the same data points, the overlap is purely temporal, meaning both datasets cover in a small part the same time period but contain different observations or series as can be seen in Table 6.
+Figure 9 illustrates the relationship between training and test data that overlap temporally.
+It is important to note that although the training and test periods overlap in time, they do not share the same data points.
+The overlap is purely temporal, meaning that both datasets cover a small portion of the same time period but contain different observations or series, as shown in Table 3.
 Figure 8: Comparison of DAX and S&P 500 points over time
@@
 Table 3: Summary of time series and data points with and without leakage
-C.2 Scaling and Model Tuning
-Min-max scaling was applied separately to each dataset to account for differences in scale across datasets and ensure optimal model performance.
+Table 4: Per-seed S&P 500 MAE breakdown for indirect leak from other countries stock indices vs no-leak experiments.
+The percentage difference shows how much better the model with leakage data performs than the non-leakage model.
+For each seed, the better MAE is displayed in bold.
+Seed MAE Before Split MAE After Split
+No Leak Indirect Leak % Diff No Leak Indirect Leak % Diff
+410 230.53 227.72 1.22% 157.85 199.33 -26.28%
+1425 219.68 178.09 18.93% 120.77 82.53 31.67%
+1680 193.52 187.19 3.27% 80.74 81.22 -0.60%
+1825 247.76 156.12 36.99% 110.47 70.36 36.31%
+2287 231.19 206.28 10.78% 102.42 100.54 1.84%
+3658 219.23 156.91 28.43% 117.25 84.62 27.84%
+4013 224.75 193.16 14.06% 114.19 97.98 14.20%
+4507 206.45 213.70 -3.51% 93.10 84.01 9.77%
+8936 189.40 199.38 -5.27% 103.59 153.67 -48.35%
+9675 206.30 182.99 11.30% 82.35 132.44 -60.82%
+Mean 216.88 190.15 11.62% 108.27 108.67 -1.44%
+Std Dev 17.15 21.88 21.10 38.83
+B.2 Data Preprocessing and Model Tuning
+Min-max scaling was applied separately to each dataset to account for differences in scale across datasets.
 Each scaler was fitted on the respective training data and subsequently used to transform the remaining data.
@@
 The Time-Series Transformer was implemented via the tsai library [^10] https://timeseriesai.github.io/tsai.
-The tuning was performed via Optuna [^12] https://github.com/optuna/optuna for 50 trials to minimize the MAE on the validation set.
+Hyperparameter tuning was performed via Optuna [^11] https://github.com/optuna/optuna for 50 trials to minimize the MAE on the validation set.
 The training set consisted of the first 80% of NN5 Daily and Tourism Monthly datasets as well as January 1st, 2015 until June 30th, 2019 for no-leak and January 1st 2016 until June 30th, 2020 of the stock index data for leak.
@@
 Batch size - - - 16, 32, 64, 128
-Epochs 5 30 5 -
+Epochs 2 20 1 -
 Dropout 0.0 0.7 - -
@@
 Number of layers 2 8 1 -
+Number of heads - - - 4, 8, 16, 32
 Model dimension - - - 64, 128, 256, 512, 1024
-Number of heads - - - 4, 8, 16, 32
 Dimension of feedforward network model - - - 32, 64, 128, 256, 512
 Table 5: Optuna search space for tuning
-Parameter No-leak Leak
-Learning rate 1.018e-03 1.431e-03
-Batch size 128 128
-Epochs 30 30
-Dropout 0.5 0.3
-FC Dropout 0.27 0.004
-Number of layers 2 4
-Model dimension 128 128
-Number of heads 32 4
-Dimension of feedforward network model 32 512
-Table 8: Final model parameters
-Appendix D TSFM Data Lineage
-Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer VisionTS TTM LLMTime Time-LMM UniTime Time-MoE Chronos-2 Flowstate(under review) LightGTS Sundial TabPFN-TS Time-MoE2 TinyTimeMixer TiRex Toto YingLong (under review) Gift-Eval TOTO2 Flowstate(T) TimesFM-ICL ROSE
-Jena Weather Autoformer (Wu et al., 2021) Nature 10T ZS ZS T/T T/T ZS ZS ZS ZS T/T ZS T/T ZS T/T ZS T/T
-Jena Weather Autoformer (Wu et al., 2021) Nature H ZS T/T ZS
-Jena Weather Autoformer (Wu et al., 2021) Nature D ZS ZS T/T ZS
-BizITObs - Application AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S ZS ZS T/T ZS
-BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps 5T ZS ZS T/T ZS
-BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps H ZS T/T ZS
-BizITObs - Service AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S ZS T/T ZS
-BDG-2 Bear BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
-BDG-2 Fox BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
-BDG-2 Panther BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
-BDG-2 Rat BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
-Borealis BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
-Buildings900K BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
-IDEAL BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
-Low Carbon London BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P
-Sceaux BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
-SMART BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
-CMIP6 ClimateLearn (Nguyen et al., 2023) Climate 6H P P P P P P P P
-ERA5 ClimateLearn (Nguyen et al., 2023) Climate H P P P P P P P P P
-Alibaba Cluster Trace 2018 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P P P P P
-Azure vM Traces 2017 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P P P P P P P P
-Borg Cluster Data 2011 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P P P P P P P P
-Air Passengers Darts Dataset (Herzen et al., 2021) Econ/Fin M ZS ZS
-Aus Beer Darts Dataset (Herzen et al., 2021) Econ/Fin Q ZS ZS
-Gas Rate CO2 Darts Dataset (Herzen et al., 2021) Econ/Fin Y ZS ZS
-Monthly Milk Darts Dataset (Herzen et al., 2021) Econ/Fin M ZS ZS
-Wine Darts Dataset (Herzen et al., 2021) Sales M ZS ZS
-Wooly Darts Dataset (Herzen et al., 2021) Econ/Fin Q ZS ZS
-Heart Rate Darts Dataset (Herzen et al., 2021) Nature 0.5S ZS ZS
-Sunspots Darts Dataset (Herzen et al., 2021) Nature M ZS
-Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T ZS T/T ZS
-Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H ZS T/T ZS
-Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T ZS T/T ZS
-Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H ZS T/T ZS
-FRED FRED (Oreshkin et al., 2021) Econ/Fin D P
-FRED FRED (Oreshkin et al., 2021) Econ/Fin Y P
-FRED FRED (Oreshkin et al., 2021) Econ/Fin W P
-FRED FRED (Oreshkin et al., 2021) Econ/Fin M P
-FRED FRED (Oreshkin et al., 2021) Econ/Fin Q P
-M5 GluonTS (Alexandrov et al., 2020a) Sales D ZS P P P P P P P ZS ZS
-Taxi GluonTS (Alexandrov et al., 2020a) Transport 30T T/T P P P P P T/T P P P P
-Taxi GluonTS (Alexandrov et al., 2020a) Transport H P P possible P P P P P P
-Uber TLC Daily GluonTS (Alexandrov et al., 2020a) Transport D T/T P P P P P T/T P P P
-Uber TLC Hourly GluonTS (Alexandrov et al., 2020a) Transport H T/T P P P P P P T/T P P P
-Wiki-Rolling GluonTS (Alexandrov et al., 2020a) Web D P P P P
-Trends hourly Google Trends Web H P P
-Trends weekly Google Trends Web W P P
-Trends monthly Google Trends Web D P P
-Requests Huawei Cloud (Joosen et al., 2023) CloudOps T ZS
-Function Delay Huawei Cloud (Joosen et al., 2023) CloudOps T P
-Platform Delay Huawei Cloud (Joosen et al., 2023) CloudOps T ZS
-CPU Usage Huawei Cloud (Joosen et al., 2023) CloudOps T P
-Memory Usage Huawei Cloud (Joosen et al., 2023) CloudOps T P
-CPU Limit Huawei Cloud (Joosen et al., 2023) CloudOps T P
-Memory Limit Huawei Cloud (Joosen et al., 2023) CloudOps T P
-Instances Huawei Cloud (Joosen et al., 2023) CloudOps T P
-ETT1 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS ZS T/T
-ETT1 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS ZS T/T
-ETT1 Informer (Zhou et al., 2020) Energy D ZS ZS ZS T/T ZS
-ETT1 Informer (Zhou et al., 2020) Energy W-THU ZS T/T ZS
-ETT2 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T ZS T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS T/T
-ETT2 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS T/T
-ETT2 Informer (Zhou et al., 2020) Energy D ZS ZS ZS T/T ZS
-ETT2 Informer (Zhou et al., 2020) Energy W-THU ZS T/T ZS
-Weather (US) Informer (Zhou et al., 2020) Weather H ZS T/T
-Brazilian Cities Temperature Kaggle Competition Weather M P possible P P P P P
-Walmart Store Sales Kaggle Competition Sales W ZS ZS ZS
-Istanbul Traffic Kaggle Competition Transport H ZS ZS ZS
-Turkey Power Kaggle Competition Energy H ZS ZS ZS ZS
-TSMC Stock Kaggle Competition Econ/Fin D ZS
-LargeST LargeST (Liu et al., 2023a) Transport 5T P P P P P P P P P P
-Beijing Subway LibCity (Wang et al., 2023a) Transport 30T P P P P P P P P P
-HZMetro LibCity (Wang et al., 2023a) Transport 15T P P P P P P P P P
-Loop Seattle LibCity (Wang et al., 2023a) Transport 5T P P P ZS P T/T ZS
-Loop Seattle LibCity (Wang et al., 2023a) Transport H ZS T/T ZS
-Loop Seattle LibCity (Wang et al., 2023a) Transport D ZS T/T ZS
-Los-Loop LibCity (Wang et al., 2023a) Transport 5T P P P P P P P T/T ZS P
-M_DENSE LibCity (Wang et al., 2023a) Transport H ZS T/T ZS
-M_DENSE LibCity (Wang et al., 2023a) Transport D ZS ZS T/T ZS
-PEMS Bay LibCity (Wang et al., 2023a) Transport 5T P P P P P P P P P P P P
-PEMS03 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P P P P P P P P P
-PEMS04 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P P P P P P P P P
-PEMS07 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P P P P P P P P P
-PEMS08 LibCity (Wang et al., 2023a) Transport 5T P P T/T ZS P P P P P P P P P P
-Q-Traffic LibCity (Wang et al., 2023a) Transport 15T P P P P P P P P P P P P
-SHMetro LibCity (Wang et al., 2023a) Transport 15T P P P P P P P P P P P
-SZ-Taxi LibCity (Wang et al., 2023a) Transport 15T P P P ZS T/T ZS P
-SZ-Taxi LibCity (Wang et al., 2023a) Transport H P ZS T/T ZS
-Rotterdam LibCity (Wang et al., 2023a) Transport 2T P P P T/T ZS
-Beijing Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P P P P P P P P P P P
-CDC Fluview ILINet LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T ZS T/T T/T P P P P P P P P
-CDC Fluview WHO NREVSS LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T P P P P P P P P
-China Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P P P P P P P P
-Favorita Sales LOTSA_Others (Woo et al., 2024) Sales D P P P P P P P P P P P
-Favorita Transactions LOTSA_Others (Woo et al., 2024) Sales D P P P P P P P
-GoDaddy LOTSA_Others (Woo et al., 2024) Econ/Fin M P P P P P P P P P
-KDD Cup 2022 LOTSA_Others (Woo et al., 2024) Energy 10T P P P P P P P P P P
-Project Tycho LOTSA_Others (Woo et al., 2024) Healthcare W P P P P P P P
-Residential Load Power LOTSA_Others (Woo et al., 2024) Energy T P P P P P P P P P P
-Residential PV Power LOTSA_Others (Woo et al., 2024) Energy T P P P P P P P P P P
-Restaurant LOTSA_Others (Woo et al., 2024) Sales D P
-Solar LSTNet (Lai et al., 2017) Energy 10T ZS T/T T/T ZS T/T
-Solar LSTNet (Lai et al., 2017) Energy 5T P possible P P P P
-Solar LSTNet (Lai et al., 2017) Energy H P ZS ZS ZS ZS P P T/T ZS P
-Solar LSTNet (Lai et al., 2017) Energy D ZS T/T ZS
-Solar LSTNet (Lai et al., 2017) Energy W-FRI ZS T/T ZS
-Exchange Rate LSTNet (Lai et al., 2017) Econ/Fin D ZS ZS T/T ZS ZS T/T P T/T ZS ZS
-Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales D P ZS T/T ZS
-Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales W-WED ZS T/T ZS
-Australian Electricity Demand Monash (Godahewa et al., 2021) Energy 30T ZS ZS T/T T/T P P P ZS P ZS P P P T/T P P ZS ZS P
-Bitcoin Monash (Godahewa et al., 2021) Econ/Fin D ZS T/T T/T P P ZS P ZS P P P T/T P P P
-Car Parts Monash (Godahewa et al., 2021) Sales M T/T T/T P ZS P ZS T/T T/T ZS ZS ZS
-CIF 2016 Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P P T/T P P ZS ZS
-COVID Deaths Monash (Godahewa et al., 2021) Healthcare D ZS ZS T/T T/T P ZS ZS P ZS T/T T/T ZS ZS ZS
-Covid Mobility Monash (Godahewa et al., 2021) Transport D T/T T/T P P T/T P P
-Elecdemand Monash (Godahewa et al., 2021) Energy 30T T/T T/T P P T/T P P
-Extended Web Traffic Monash (Godahewa et al., 2021) Web D T/T T/T P P P P T/T P P
-FRED MD Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P P P T/T P P ZS ZS P
-Hospital Monash (Godahewa et al., 2021) Healthcare M ZS ZS T/T T/T P ZS ZS P ZS T/T T/T ZS ZS ZS
-Kaggle Web Traffic Weekly Monash (Godahewa et al., 2021) Web W P P P P P P P P P P P
-KDD Cup 2018 Monash (Godahewa et al., 2021) Nature H T/T T/T T/T P P P ZS P P ZS T/T T/T T/T ZS P
-KDD Cup 2018 Monash (Godahewa et al., 2021) Nature D ZS P T/T ZS P
-London Smart Meters Monash (Godahewa et al., 2021) Energy 30T T/T P P P P P P P P P P T/T P P P P P
-M1 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T P ZS P P T/T P P ZS ZS
-M1 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P P P P P P P ZS ZS
-M1 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P P P P P P P ZS ZS
-M3 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T ZS T/T ZS P P T/T P P ZS ZS
-M3 Other Monash (Godahewa et al., 2021) Econ/Fin Q T/T T/T T/T ZS P T/T P P
-M3 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P ZS T/T P P P P P ZS ZS
-M3 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P ZS T/T P P P P P ZS ZS
-M4 Daily Monash (Godahewa et al., 2021) Econ/Fin D T/T P T/T T/T P T/T ZS P ZS T/T T/T T/T ZS P P
-M4 Hourly Monash (Godahewa et al., 2021) Econ/Fin H T/T P T/T T/T P T/T ZS P ZS T/T T/T T/T ZS P P
-M4 Monthly Monash (Godahewa et al., 2021) Econ/Fin M T/T P T/T T/T ZS T/T ZS P ZS T/T T/T T/T ZS P P
-M4 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q-DEC ZS P P P ZS T/T P ZS P T/T ZS ZS P
-M4 Weekly Monash (Godahewa et al., 2021) Econ/Fin W-SUN T/T P T/T T/T T/T ZS P ZS T/T T/T T/T ZS P
-M4 Yearly Monash (Godahewa et al., 2021) Econ/Fin A-DEC ZS P P P ZS T/T P ZS P T/T ZS ZS P
-NN5 Daily Monash (Godahewa et al., 2021) Econ/Fin D ZS ZS T/T T/T P ZS P ZS P P P T/T P P ZS ZS P
-NN5 Weekly Monash (Godahewa et al., 2021) Econ/Fin W ZS ZS T/T T/T P ZS ZS P T/T P P ZS ZS
-Oikolab Weather Monash (Godahewa et al., 2021) Climate H T/T T/T P P T/T P P
-Pedestrian Counts Monash (Godahewa et al., 2021) Transport H T/T ZS T/T T/T P ZS P ZS ZS P P P T/T T/T P P P P
-Rideshare Monash (Godahewa et al., 2021) Transport H T/T T/T T/T P ZS P P T/T T/T P P P
-Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer VisionTS TTM LLMTime Time-LMM UniTime Time-MoE Chronos-2 Flowstate(under review) LightGTS Sundial TabPFN-TS Time-MoE2 TinyTimeMixer TiRex Toto YingLong (under review) Gift-Eval TOTO2 Flowstate(T) TimesFM-ICL ROSE
-Saugeen Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P ZS P ZS P ZS P T/T T/T ZS P
-Saugeen Monash (Godahewa et al., 2021) Nature W-THU ZS T/T ZS
-Saugeen Monash (Godahewa et al., 2021) Nature M ZS T/T ZS
-Solar Power (Australia) Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P P P T/T P P P P P
-Solar Monash (Godahewa et al., 2021) Energy W ZS P P ZS ZS
-Solar Monash (Godahewa et al., 2021) Energy 10M P P P ZS
-Sunspot Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P P ZS P ZS P P P T/T P P P
-Temperature Rain Monash (Godahewa et al., 2021) Nature D T/T T/T T/T P P ZS P ZS P T/T T/T T/T ZS P P
-Tourism Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS P T/T ZS ZS P P P P ZS ZS
-Tourism Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS ZS P T/T ZS ZS P P P P ZS ZS
-Tourism Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS ZS P P P T/T ZS P P P P P ZS ZS
-Traffic Hourly Monash (Godahewa et al., 2021) Transport H ZS P T/T T/T T/T P T/T ZS ZS ZS ZS T/T P P T/T T/T P P ZS P T/T
-Traffic Weekly Monash (Godahewa et al., 2021) Transport W ZS T/T T/T ZS P ZS ZS P P T/T P P
-US Births Monash (Godahewa et al., 2021) Healthcare D ZS T/T T/T P P ZS P ZS P ZS P T/T T/T P
-US Births Monash (Godahewa et al., 2021) Healthcare W-TUE ZS T/T ZS
-US Births Monash (Godahewa et al., 2021) Healthcare M ZS T/T ZS
-Vehicle Trips Monash (Godahewa et al., 2021) Transport D T/T T/T P ZS P P T/T P P
-Weather (Australia) Monash (Godahewa et al., 2021) Climate D ZS ZS T/T T/T P ZS ZS P T/T T/T P P P T/T P P ZS P
-Wind Farms Monash (Godahewa et al., 2021) Energy T T/T T/T P P P P T/T P P P
-Wind Farms Monash (Godahewa et al., 2021) Energy H P P P P P
-Wind Farms Monash (Godahewa et al., 2021) Energy D P possible P P P P
-Wind Farms Monash (Godahewa et al., 2021) Energy 5T P possible P
-Wind Power Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P P P T/T P P P P P P
-Kaggle Web Traffic Daily Monash (Godahewa et al., 2021) Web D P P P P
-Restaurant Recruit Other Sales D ZS T/T ZS
-Subseasonal Other Climate D P P P P P P P P P P
-Subseasonal Precipitation Other Climate D P P P P P P P P P
-USHCN Other Weather D P possible P P P P P
-Mexico City Bikes Other Transport H P P possible P P P P P
-Dominick Other Sales D ZS P P P ZS ZS
-ERCOT Load Other Energy H ZS P P ZS ZS
-Weather (US) Other Weather 10T P ZS
-Bike Sharing Other Transport H ZS ZS
-CC Carbon Other Energy 2T ZS
-Air Quality Other Nature H P
-Beijing PM 2.5 Other Nature H ZS
-TDBrain Other Healthcare 0.002S P P P
-SensorData Other IoT 0.02S P
-BDG-2 Bull ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
-BDG-2 Cockatoo ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
-BDG-2 Hog ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
-Covid19 Energy ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
-ELF ProEnFo (Wang et al., 2023b) Energy H P P P P P P
-GEF12 ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
-GEF14 ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
-GEF17 ProEnFo (Wang et al., 2023b) Energy H P P P P P P
-PDB ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
-Spanish ProEnFo (Wang et al., 2023b) Energy H P P P P P P P P P P
-Australia Rainfall Tan et al., 2021 Nature H P
-Beijing PM 2.5 Tan et al., 2021 Nature H P
-BenzeneConcentration Tan et al., 2021 Nature H P
-IEEEPPG Tan et al., 2021 Healthcare 0.008S P
-BIDMC32HR Tan et al., 2021 Healthcare - P
-Dodger TSB-UAD (Paparrizos et al., 2022) Traffic - P
-ECG TSB-UAD (Paparrizos et al., 2022) Healthcare - P P
-IOPS TSB-UAD (Paparrizos et al., 2022) Web - P
-KDD21 TSB-UAD (Paparrizos et al., 2022) Multiple - P
-MGAB TSB-UAD (Paparrizos et al., 2022) Synthetic - P
-NAB TSB-UAD (Paparrizos et al., 2022) Web - P
-SensorScope TSB-UAD (Paparrizos et al., 2022) Nature - P
-YAHOO TSB-UAD (Paparrizos et al., 2022) Web - P
-NASA-MSL TSB-UAD (Paparrizos et al., 2022) Sensor - P
-NASA-SMAP TSB-UAD (Paparrizos et al., 2022) Sensor - P
-Daphnet TSB-UAD (Paparrizos et al., 2022) Gait - P
-GHL TSB-UAD (Paparrizos et al., 2022) Sensor - P
-Genesis TSB-UAD (Paparrizos et al., 2022) Sensor - P
-MITDB TSB-UAD (Paparrizos et al., 2022) Healthcare - P
-OPP TSB-UAD (Paparrizos et al., 2022) Sensor - P
-Occupancy TSB-UAD (Paparrizos et al., 2022) Sensor - P
-SMD TSB-UAD (Paparrizos et al., 2022) Web - P
-SZSDB TSB-UAD (Paparrizos et al., 2022) Healthcare - P
-Electricity UCI ML Archive (Trindade, 2015) Energy 15T T/T P T/T P ZS T/T T/T T/T ZS P
-Electricity UCI ML Archive (Trindade, 2015) Energy H T/T P ZS ZS T/T P T/T ZS ZS ZS ZS T/T T/T ZS T/T ZS T/T T/T ZS P P T/T
-Electricity UCI ML Archive (Trindade, 2015) Energy D ZS ZS T/T ZS
-Electricity UCI ML Archive (Trindade, 2015) Energy W-FRI T/T ZS T/T T/T ZS
-Electricity UCI ML Archive (Trindade, 2015) Energy M T/T T/T
-Electricity UCI ML Archive (Trindade, 2015) Energy W P P P
-SemgHandGenderCh2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
-GestureMidAirD2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
-UWaveGestureLibraryAll UCR Archive (Ismail Fawaz et al., 2019) Human - T/T
-SelfRegulationSCP1 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
-UWaveGestureLibraryX UCR Archive (Ismail Fawaz et al., 2019) Human - P
-GesturePebbleZ2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
-Healthcare5000 (orignally: ECG5000) UCR Archive (Ismail Fawaz et al., 2019) Healthcare - T/T P
-OSULeaf UCR Archive (Ismail Fawaz et al., 2019) Nature - T/T
-MedicalImages UCR Archive (Ismail Fawaz et al., 2019) Healthcare - T/T
-Haptics UCR Archive (Ismail Fawaz et al., 2019) Human - P
-LargeKitchenAppliances UCR Archive (Ismail Fawaz et al., 2019) Power - P
-Japanesevowels UCR Archive (Ismail Fawaz et al., 2019) Audio - P
-Worms UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P
-Ham UCR Archive (Ismail Fawaz et al., 2019) Facilities - T/T
-DistalPhalanxTW UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-ProximalPhalanxOutlineCorrect UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-SemgHandMovementCh2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
-RefrigerationDevices UCR Archive (Ismail Fawaz et al., 2019) Power - P
-FreezerRegularTrain UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-PigAirwayPressure UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-TwoLeadECG UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
-GunPointMaleversusFemale UCR Archive (Ismail Fawaz et al., 2019) Human - P
-Trace UCR Archive (Ismail Fawaz et al., 2019) Power - P
-SmoothSubspace UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-MiddlePhalanxTW UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-AtrialFibrillation UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
-SyntheticControl UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-ShapesAll UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-Human UCR Archive (Ismail Fawaz et al., 2019) BodyZSerticalSignal - P
-PLAID UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-AllGestureWiimoteX UCR Archive (Ismail Fawaz et al., 2019) Human - P
-Heartbeat UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
-Wafer UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-FaceFour UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-Phoneme UCR Archive (Ismail Fawaz et al., 2019) Audio - P P P P
-InlineSkate UCR Archive (Ismail Fawaz et al., 2019) Human - P
-CricketX UCR Archive (Ismail Fawaz et al., 2019) Human Body - P
-SelfRegulationSCP2 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
-DistalPhalanxOutlineCorrect UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-ChlorineConcentration UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-Chinatown UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
-GestureMidAirD1 UCR Archive (Ismail Fawaz et al., 2019) Human - P
-MiddlePhalanxOutlineAgeGroup UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-UMD UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-Crop UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-PenDigits UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-GesturePebbleZ1 UCR Archive (Ismail Fawaz et al., 2019) Human - P
-Handwriting UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-Mallat UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-ERing UCR Archive (Ismail Fawaz et al., 2019) Human - P
-StarLightCurves UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P P
-WordSynonyms UCR Archive (Ismail Fawaz et al., 2019) Audio - P
-PEMS-SF UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
-FaceDetection UCR Archive (Ismail Fawaz et al., 2019) Human - P
-Computers UCR Archive (Ismail Fawaz et al., 2019) Power - P
-ArrowHead UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-Wine UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-Coffee UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-Earthquakes UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-Herring UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-Lightning2 UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-Beef UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-MiddlePhalanxOutlineCorrect UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-HealthcareFiveDays UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
-Yoga UCR Archive (Ismail Fawaz et al., 2019) Human - P
-Adiac UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-HandMovementDirection UCR Archive (Ismail Fawaz et al., 2019) Human - P
-Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer VisionTS TTM LLMTime Time-LMM UniTime Time-MoE Chronos-2 Flowstate(under review) LightGTS Sundial TabPFN-TS Time-MoE2 TinyTimeMixer TiRex Toto YingLong (under review) Gift-Eval TOTO2 Flowstate(T) TimesFM-ICL ROSE
-MoteStrain UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-Rock UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-Strawberry UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-InsectWingbeatSound UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-DodgerLoopWeekend UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
-MixedShapesSmallTrain UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-EthanolConcentration UCR Archive (Ismail Fawaz et al., 2019) Power - P
-OliveOil UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-Meat UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-MelbournePedestrian UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
-Car UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-FaceAll UCR Archive (Ismail Fawaz et al., 2019) Human - P
-FacesUCR UCR Archive (Ismail Fawaz et al., 2019) Human - P
-AllGestureWiimoteY UCR Archive (Ismail Fawaz et al., 2019) Human - P
-NATOPS UCR Archive (Ismail Fawaz et al., 2019) Human - P
-SemgHandSubjectCh2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
-ShakeGestureWiimoteZ UCR Archive (Ismail Fawaz et al., 2019) Human - P
-Cricket UCR Archive (Ismail Fawaz et al., 2019) Human - P
-BME UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-EigenWorms UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P P
-FordB UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-NonInvasiveFetalHealthcareThorax1 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
-UWaveGestureLibrary UCR Archive (Ismail Fawaz et al., 2019) Human - P
-CinCHealthcareTorso UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
-PigArtPressure UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P P
-Fish UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-SonyAIBORobotSurface2 UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-FiftyWords UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-MotorImagery UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
-ToeSegmentation1 UCR Archive (Ismail Fawaz et al., 2019) Human - P
-PhonemeSpectra UCR Archive (Ismail Fawaz et al., 2019) Audio - P
-FreezerSmallTrain UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-TwoPatterns UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-ShapeletSim UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-Plane UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-GestureMidAirD3 UCR Archive (Ismail Fawaz et al., 2019) Human - P
-DiatomSizeReduction UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-Human UCR Archive (Ismail Fawaz et al., 2019) BodyHorizontalSignal - P
-CricketZ UCR Archive (Ismail Fawaz et al., 2019) Human - P
-StandWalkJump UCR Archive (Ismail Fawaz et al., 2019) Human - P
-WormsTwoClass UCR Archive (Ismail Fawaz et al., 2019) Human - P
-Lightning7 UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-UWaveGestureLibraryY UCR Archive (Ismail Fawaz et al., 2019) Human - P
-GunPointAgeSpan UCR Archive (Ismail Fawaz et al., 2019) Human - P
-DistalPhalanxOutlineAgeGroup UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-SwedishLeaf UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-LSST UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-CBF UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-BeetleFly UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-Libras UCR Archive (Ismail Fawaz et al., 2019) Human - P
-HouseTwenty UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-ScreenType UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-InsectEPGSmallTrain UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-AllGestureWiimoteZ UCR Archive (Ismail Fawaz et al., 2019) Sensor - P
-DodgerLoopDay UCR Archive (Ismail Fawaz et al., 2019) Sensor - P
-NonInZSasiZSeFetalHealthcareThorax2 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
-BasicHuman UCR Archive (Ismail Fawaz et al., 2019) Bodys - P
-GunPointOldversusYoung UCR Archive (Ismail Fawaz et al., 2019) Human - P
-FordA UCR Archive (Ismail Fawaz et al., 2019) Sensor - P
-InsectWingbeat UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-ItalyPowerDemand UCR Archive (Ismail Fawaz et al., 2019) Power - P
-ProximalPhalanxOutlineAgeGroup UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-ACSF1 UCR Archive (Ismail Fawaz et al., 2019) Power - P
-GunPoint UCR Archive (Ismail Fawaz et al., 2019) Human - P
-RacketSports UCR Archive (Ismail Fawaz et al., 2019) Human - P
-SmallKitchenAppliances UCR Archive (Ismail Fawaz et al., 2019) Power - P
-ProximalPhalanxTW UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-DuckDuckGeese UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-PickupGestureWiimoteZ UCR Archive (Ismail Fawaz et al., 2019) Human - P
-EthanolLevel UCR Archive (Ismail Fawaz et al., 2019) Power - P
-SpokenArabicDigits UCR Archive (Ismail Fawaz et al., 2019) Audio - P
-SonyAIBORobotSurface1 UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
-HandOutlines UCR Archive (Ismail Fawaz et al., 2019) Human - P
-PowerCons UCR Archive (Ismail Fawaz et al., 2019) Power - P
-PhalangesOutlinesCorrect UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-BirdChicken UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-ToeSegmentation2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
-PigCVP UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
-CricketY UCR Archive (Ismail Fawaz et al., 2019) Human - P
-FingerMovements UCR Archive (Ismail Fawaz et al., 2019) Human - P
-ElectricDevices UCR Archive (Ismail Fawaz et al., 2019) Power - P
-InsectEPGRegularTrain UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-DodgerLoopGame UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
-Fungi UCR Archive (Ismail Fawaz et al., 2019) Nature - P
-Symbols UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-MixedShapesRegularTrain UCR Archive (Ismail Fawaz et al., 2019) Generated - P
-ArticularyWordRecognition UCR Archive (Ismail Fawaz et al., 2019) Human - P
-UWaveGestureLibraryZ UCR Archive (Ismail Fawaz et al., 2019) Human - P
-Epilepsy UCR Archive (Ismail Fawaz et al., 2019) Human - P
-Healthcare200 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
-Weatherbench (Daily) Weather Bench Weather D P P possible P P P P P
-Weatherbench (Hourly) Weather Bench Weather H P P possible P P P P P
-Weatherbench (Weekly) Weather Bench Weather W P P possible P P P P P
-Wiki hourly Wikipedia Metrics Web H P P
-Wiki daily Wikipedia Metrics Web D P P P possible P P P P P
-Wiki weekly Wikipedia Metrics Web W P P
-Wiki monthly Wikipedia Metrics Web M P P
-Global Temp Corrformer (Wu et al., 2023) Weather 10T T/T
-TSMixup 10M Chronos (Ansari et al., 2024) Synthetic - P
-KernelSynth 1M Chronos (Ansari et al., 2024) Synthetic - P
-TaxiBJ Zhang et al., 2017 Traffic and weather 30T ZS
-ERA5 Munoz-Sabater et al., 2021 Weather 3h P P
-ERA5 Munoz-Sabater et al., 2021 Weather 12h P
-ERA5 Munoz-Sabater et al., 2021 Weather D P
-ERA5 Munoz-Sabater et al., 2021 Weather W P
-ERA5 Munoz-Sabater et al., 2021 Weather M P
-ERA5 Munoz-Sabater et al., 2021 Weather Q P
-Finance (not published) Liu et al., 2025 (Sundial) Finance ? P
-IoT (not published) Liu et al., 2025 (Sundial) IoT ? P
-BOOM Cohen et al. (2025) (TOTO) Observational 10S ZS
-BOOM Cohen et al. (2025) (TOTO) Observational T ZS
-BOOM Cohen et al. (2025) (TOTO) Observational 5T ZS
-BOOM Cohen et al. (2025) (TOTO) Observational 10T ZS
-BOOM Cohen et al. (2025) (TOTO) Observational 30T ZS
-BOOM Cohen et al. (2025) (TOTO) Observational H ZS
-BOOM Cohen et al. (2025) (TOTO) Observational D
+Seed Learning rate Batch size Epochs Dropout FC Dropout Number of layers Number of heads $d_{\text{model}}$ $d_{\text{ff}}$
+410 9.687e-04 64 11 0.272 0.572 6 16 128 64
+1425 5.889e-04 64 19 0.201 0.593 5 32 256 256
+1680 5.773e-04 64 18 0.162 0.251 7 16 128 32
+1825 4.392e-03 128 17 0.009 0.439 6 4 64 512
+2287 4.828e-04 32 17 0.447 0.644 2 32 256 64
+3658 3.502e-04 64 6 0.001 0.211 2 16 128 128
+4013 5.348e-03 128 11 0.096 0.439 3 16 64 128
+4507 3.656e-04 64 14 0.007 0.369 3 8 128 32
+8936 5.428e-04 128 9 0.003 0.595 8 32 64 512
+9675 3.102e-04 64 14 0.006 0.504 2 4 64 256
+Table 6: Best hyperparameters per seed for no-leak models
+Seed Learning rate Batch size Epochs Dropout FC Dropout Number of layers Number of heads $d_{\text{model}}$ $d_{\text{ff}}$
+410 1.851e-04 32 18 0.334 0.095 7 16 32 256
+1425 1.156e-02 64 8 0.067 0.691 2 16 32 256
+1680 3.105e-04 128 16 0.032 0.267 3 4 128 256
+1825 8.714e-04 128 19 0.093 0.696 6 32 32 256
+2287 2.703e-03 128 19 0.183 0.435 2 8 64 128
+3658 1.645e-03 128 19 0.064 0.293 5 16 64 32
+4013 2.590e-03 128 14 0.090 0.547 7 32 64 512
+4507 4.160e-04 64 20 0.395 0.006 2 8 64 256
+8936 1.328e-04 64 5 0.011 0.501 8 16 128 512
+9675 3.373e-04 128 13 0.010 0.428 3 4 64 256
+Table 7: Best hyperparameters per seed for leak models